### v2

# Finite-difference least square methods for solving Hamilton-Jacobi equations using neural networks

Abstract
We present a simple algorithm to approximate the viscosity solution of Hamilton-Jacobi (HJ) equations by means of an artificial deep neural network. The algorithm uses a stochastic gradient descent-based algorithm to minimize the least square principle defined by a monotone, consistent numerical scheme. We analyze the least square principle’s critical points and derive conditions that guarantee that any critical point approximates the sought viscosity solution. The use of a deep artificial neural network on a finite difference scheme lifts the restriction of conventional finite difference methods that rely on computing functions on a fixed grid. This feature makes it possible to solve HJ equations posed in higher dimensions where conventional methods are infeasible. We demonstrate the efficacy of our algorithm through numerical studies on various canonical HJ equations across different dimensions, showcasing its potential and versatility.
Keywords Hamilton-Jacobi equations, deep learning, finite-difference methods, least squares principle, optimal control and differential games.
AMS subject classifications (MSC2020) 49L25, 35A15, 68T07, 49K20, 65N06
1 Introduction
In recent years, significant research has focused on computing solutions for partial differential equations in high dimensions. These endeavours are driven by advancements in deep learning and new numerical optimization algorithms. In particular, there is great interest in numerical algorithms for Hamilton-Jacobi equations arising in optimal control problems and differential games.
We consider the boundary value problem associated with a Hamilton-Jacobi equation of the form
$\begin{cases}H(x,\nabla u(x))=0&x\in\Omega\\ u(x)=g(x)&x\in\partial\Omega,\end{cases}$ (1.1)
where $\Omega\subset\mathbb{R}^{d}$, with $d\geq 1$, is an open bounded domain with Lipschitz boundary, and
$H:\Omega\times\mathbb{R}^{d}\to\mathbb{R}\quad\text{and}\quad g:\partial\Omega\to\mathbb{R}$
are the given functions, which we shall refer to as the Hamiltonian and the boundary condition, respectively. Throughout the paper, we shall assume that
$x\mapsto g(x)\ \text{is Lipschitz on $\partial\Omega$}$ (1.2)
and
$(x,p)\mapsto H(x,p)\ \text{is locally Lipschitz in $\Omega\times\mathbb{R}^{d}$ and differentiable with respect to $p$.}$ (1.3)
Certain types of Hamilton-Jacobi equations describe the optimality condition for an optimal control problem. By applying Bellman’s dynamic programming principle, one can show that the value function of the associated optimal control problem formally satisfies the so-called Hamilton-Jacobi-Bellman equation (see [1]). However, it is well-known that even if the boundary condition $g$ is very smooth, the boundary value problem (1.1) is not guaranteed a classical solution. This happens when the characteristics of (1.1), representing the optimal trajectories of the optimal control problem, collide with each other. On the other hand, continuous functions that satisfy the PDE for almost every $x\in\Omega$ along with the boundary condition are not necessarily unique. Crandall and Lions introduced the notion of viscosity solution to single out a unique weak solution for the problem [5]. Furthermore, in many situations, one can prove that the viscosity solution coincides with the value function, and one may use it to synthesize the optimal feedback for the control problem.
A characterisation of the viscosity solution can be obtained through the so-called vanishing viscosity method, which consists of approximating the viscosity solution by the unique classical solution to the problem
$\begin{cases}H(x,\nabla u^{\varepsilon}(x))-\varepsilon\Delta u^{\varepsilon}(x)=0,&x\in\Omega,\\ u^{\varepsilon}(x)=g(x),&x\in\partial\Omega,\end{cases}$ (1.4)
for $\varepsilon>0$. The viscosity solution to (1.1) is obtained as the point-wise limit of $u^{\varepsilon}(x)$ as $\ to0^{+}$. A different characterisation of the viscosity solution exists, using smooth test functions, which is based on the comparison principle (see [4]). Depending on the structure of the Hamiltonian and the boundary data, the viscosity solution may be discontinuous (see [1, 9]). This is typically true for Hamilton–Jacobi equations arising in game theory or when the boundary condition is prescribed along a characteristic curve. However, in this paper, we make the following assumption:
Assumption 1.
The boundary value problem (1.1) has a Lipschitz continuous viscosity solution.
The design and implementation of numerical algorithms to approximate the viscosity solution to (1.1) is a classical problem widely considered throughout the years. In two or three dimensions, Hamilton-Jacobi equations are typically solved by special finite-difference methods operating on uniform Cartesian grids. The Hamiltonian $H$ is approximated by a so-called numerical Hamiltonian $\widehat{H}$ defined on the grid and the grid function. If $\widehat{H}$ is consistent to $H$ and monotone (see Definition 1), the solution to the system of discrete equations defined by $\widehat{H}$ on the grid (properly extended to the domain $\overline{\Omega}$) converges uniformly to the viscosity solution as the grid spacing tends to 00 (see e.g. [6, 2]). Similar to the development of numerical schemes for hyperbolic conservation laws, higher-order approximations to the partial derivatives are proposed to be used with a monotone and consistent numerical Hamiltonian. In parallel, fast algorithms such as the fast marching methods [19, 15, 16, 8] and fast sweeping methods [20, 18] are developed for certain class of Hamilton-Jacobi equations to solve the resulting nonlinear system with computational complexities that are linear in $N_{d}$, the total number of unknowns (for the fast marching methods, the formal complexity is $N_{d}\log N_{d}$).
The main drawback of finite-difference methods implemented in the classical fashion is that they involve solving a system of nonlinear equations with as many unknowns as the number of grid points. This means that the complexity of these methods scales at least exponentially with the dimension of the domain. Therefore, the classical method of formulating and solving for finite difference approximations on grid functions cannot be carried out in the high-dimensional setup. In many control theory applications, the state’s dimension is large, and thence, the associated Hamilton-Jacobi-Bellman equation has to be stated in a high-dimensional domain. Recently, there has been substantial development on algorithms for solving high dimensional problems utilizing ideas from the Hopf-Lax formula and its generalization; see, e.g. [7],[3].
In recent years, Deep Learning approaches have proven very effective at solving some non-linear issues in high dimensions. These methods consist on approximating a target function (in our case, the viscosity solution) by an artificial neural network, whose parameters are optimised by minimising a loss function (typically) through a gradient-based method. This can be formulated as a variational problem of the form
${u}^{*}\in\arg\min_{u\in\mathcal{F}}\mathcal{J}(u),$ (1.5)
where
:=$\mathcal{F}:=\{u(\cdot,\theta)\,:\ \theta\in\mathbb{R}^{P}\}\subset C(\overline{\Omega})$
is a parameterized class of functions given by a Neural Network architecture, and $\mathcal{J}:C(\overline{\Omega})\to\mathbb{R}^{+}$ is the loss functional. Of course, the success of this method heavily relies on the choice of $\mathcal{F}$, $\mathcal{J}(\cdot)$ and the optimization algorithm.
Concerning the choice of the functional, it is necessary that $\mathcal{J}(u)$ yields smaller values for functions $u$ close to the target function. A common practice in Deep Learning is to seek a minimizer of $\mathcal{J}(\cdot)$ in $\mathcal{F}$ using information derived from $\nabla_{\theta}\mathcal{J}(\Phi(\cdot,\theta)).$ Let us consider the following continuous version of the gradient descent method to find a minimizer of $\mathcal{J}$ in $\mathcal{F}$:
$\frac{d}{dt}\theta_{t}=-\nabla_{\theta}\mathcal{J}(u(\cdot,\theta_{t}))=-\mathcal{J}^{\prime}(u(\cdot,\theta_{t}))\,\nabla_{\theta}u(\cdot,\theta_{t}),$ (1.6)
where, for any continuous function $u\in C(\overline{\Omega})$, $\mathcal{J}^{\prime}(u)\in C(\overline{\Omega})^{\ast}$ denotes the Fréchet derivative of $\mathcal{J}(\cdot)$ at $u$, and $\nabla_{\theta}u(\cdot,\theta)$ denotes the vector–valued function with the partial derivatives of $u(x,\theta)$ with respect to the parameter $\theta$. We see that a parameter $\theta\in\mathbb{R}^{P}$ is a stationary point of (1.6) if and only if all the elements of the vector $\nabla_{\theta}u(\cdot,\theta)\in[C(\overline{\Omega})]^{P}$ lie in the kernel of the linear functional $\mathcal{J}^{\prime}(u(\cdot,\theta))$. The precise form of the vector-valued function $\nabla_{\theta}u(\cdot,\theta)$ depends on the architecture of the Neural Network and will not be discussed in this paper. On the other hand, any parameter $\theta\in\mathbb{R}^{P}$ such that the function $u=u(\cdot,\theta)$ is a solution to the Euler-Lagrange equation $\mathcal{J}^{\prime}(u)=0$ is a critical point of $\theta\mapsto\mathcal{J}(u(\cdot,\theta))$, and therefore, a stationary point for the gradient descent method.
In the current paper, we investigate the choice of the functional $\mathcal{J}(\cdot)$. Namely, we are interested in constructing a functional with the two following properties:
  1. i.
    Any global minimiser of $\mathcal{J}(u)$ approximates the viscosity solution of (1.1).
  2. ii.
    Any critical point of $\mathcal{J}(u)$ is a global minimizer.
Note that the second property above is crucial in Deep Learning since the optimisation algorithm used to minimise $\mathcal{J}(u)$ is typically a variant of gradient descent, similar to (1.6). In particular, we shall look at functionals with the structure
$\mathcal{J}(u)=\mathcal{L}(u;g)+\mathcal{R}(u),$
where $\mathcal{L}(u;g)$ and $\mathcal{R}(u)$ are, respectively, fitting terms related to the boundary data $g$, and the PDE in (1.1). The main contribution of the current paper is to prove that if we construct $\mathcal{R}(u)$ based on a suitable Lax-Friedrichs numerical Hamiltonian, then one can ensure that the two properties stated above hold.
In the next section, we illustrate our findings with a simple example. In particular, we consider the Eikonal equation in a one-dimensional interval. First, we compute the optimality condition for a functional based on the PDE residual. Then, we carry out the same computation for a functional based on the residual of a finite-difference approximation of the Hamiltonian. In Section 3, we discuss existing approaches based on Deep Learning for the numerical approximation of PDEs, such as the well-known framework known as Physics informed Neural Networks, or PINNs. In Section 4, we study the critical points of loss functionals based on numerical Hamiltonians of Lax-Friedrichs type and present our main theoretical contributions. Section 5 is devoted to the numerical aspects of our approach. First, we describe our algorithm to approximate the viscosity solution employing a Neural Network, and then, we test our method in several examples of Hamilton-Jacobi equations. In Section 6, we present the proofs of the theoretical results presented in Section 4. Finally, we conclude the paper in Section 7, summarising our contributions and some perspectives for future research.
The code to reproduce the numerical experiments from Section 5 is publicly available in the following link:
2 Example
Consider the model viscous Eikonal equation on the unit interval, with Dirichlet boundary conditions:
$\begin{cases}(\partial_{x}u)^{2}=1+\varepsilon\partial_{xx}u&x\in(0,1),\\ u(0)=u_{0},\quad u(1)=u_{1}.\end{cases}$ (2.1)
It is well-known that, for any $\varepsilon\neq 0$ and $u_{0},u_{1}\in\mathbb{R}$, the problem (2.1) has a unique $C^{\infty}$-solution.
Let us consider the minimisation problem
[…]
among functions in $C^{\infty}(0,1)$ satisfying the boundary conditions $u(0)=u_{0}$ and $u(1)=u_{1}$.
For any test function $v\in C_{0}^{2}(0,1)$, the Gâteaux derivative of $\mathcal{R}_{\varepsilon}(\cdot)$ at $u$ in the direction $v$ can be written as
[…]
Now, suppose that $u^{*}\in C^{\infty}(0,1)$ is a critical point of $\mathcal{R}_{\varepsilon}(\cdot)$. Then, $w$ is a weak solution of
$\begin{cases}\partial_{x}(w\partial_{x}u^{*})+\varepsilon\partial_{xx}w=0&\text{in}\ (0,1),\\ w(0)=w(1)=0.\end{cases}$
It is clear that, for any $\varepsilon\neq 0$, the unique weak solution to this boundary value problem is $w=0$. Therefore any critical point $u^{*}\in C^{\infty}(0,1)$ of $\mathcal{R}_{\varepsilon}(\cdot)$, satisfying the boundary conditions is a classical solution of
$(\partial_{x}u)^{2}-\varepsilon\partial_{xx}u-1=0.$
Since this PDE, coupled with boundary conditions $u(0)=u_{0}$ and $u(1)=u_{1}$, has a unique classical solution, we deduce that this solution is the unique critical point of $\mathcal{R}_{\varepsilon}(\cdot)$ in $C^{\infty}(0,1)$.
The sign of $\varepsilon$ determines the convexity of the solution. The commonly sought-after viscosity solution, in this case, corresponds to the limit of the solution as $\varepsilon\rightarrow 0^{+},$ and can be proven to be concave. If we consider $\varepsilon=0$, the boundary value problem (2.1) has no classical solution, so there are no critical points of $\mathcal{R}_{0}(\cdot)$ in $C^{\infty}(0,1)$. This implies, in particular, that the functional $\mathcal{R}_{0}(\cdot)$ is not minimised by any smooth function. One can, however, construct minimisers of $\mathcal{R}_{0}(u)$, which are Lipschitz continuous. However, these are not unique, as illustrated in Figure 1.
Figure 1: Four weak solutions of the Eikonal equation $(\partial_{x}u)^{2}-1=0$ in $\Omega:=(-1,1)$ with boundary condition $u(-1)=u(1)=0$, correspodning to four global minimizers to the functional (3.2).
Despite having proved the uniqueness of a critical point in $C^{\infty}(0,1)$ for any $\varepsilon>0$, it is worth mentioning that there exist other continuous global minimizers of $\mathcal{R}_{\varepsilon}(\cdot)$ which satisfy the boundary condition and are not $C^{\infty}$. Consider, for instance, the boundary value problems
$\begin{cases}(\partial_{x}\phi_{1})^{2}=1+\varepsilon\partial_{xx}\phi_{1}\qquad x\in(0,\,1/2),\\ \phi_{1}(0)=u_{0},\quad\phi_{1}(1/2)=u_{1/2}.\end{cases}\qquad\text{and}\qquad\begin{cases}(\partial_{x}\phi_{2})^{2}=1+\varepsilon\partial_{xx}\phi_{2}\qquad x\in(1/2,\,1),\\ \phi_{2}(1/2)=u_{1/2},\quad\phi_{2}(1)=u_{1},\end{cases}$ (2.3)
[…]
is continuous and minimises the functional $\mathcal{R}_{\varepsilon}(\cdot)$, since it satisfies the boundary condition and the PDE for a.e. $x\in(0,1)$. This proves that the minimiser of $\mathcal{R}_{\varepsilon}(\cdot)$ is not unique in the space of continuous functions. Although we could constraint the optimization to a class of $C^{\infty}$ functions, this would not solve the problem due to the density of $C^{\infty}(0,1)$ in $C(0,1)$. Indeed, there exist minimizing sequences of $C^{\infty}$ functions converging to global minimisers of $\mathcal{R}_{\varepsilon}(\cdot)$ which are not the classical solution to (2.1).
To overcome the above issue, we consider minimising the $L^{2}$ residual of a finite-difference numerical scheme instead. For any $N\in\mathbb{N}$ and $\delta:=1/N$, set the uniform grid $\overline{\Omega}_{\delta}:=\delta\mathbb{Z}\cap[0,1]$, and consider the following discrete version of the PDE-residual functional $\mathcal{R}_{\varepsilon}(\cdot)$ defined in (2.2), with $\varepsilon=0$:
$\min_{u}\widehat{\mathcal{R}}(u):=F(U)=\sum_{i=1}^{N-1}\delta\left[\widehat{H}_{\alpha}\left(\frac{u_{i+1}-u_{i}}{\delta},\frac{u_{i}-u_{i-1}}{\delta}\right)\right]^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to}\ u_{0}=u_{N}=0,$ (2.4)
where, for any $u\in C([0,1])$, the vector $U:=(u_{0},\ldots,u_{N})\in\mathbb{R}^{N+1}$ is the grid function on $\overline{\Omega}_{\delta}$ associated to $u$, and
$\widehat{H}_{\alpha}(p^{+},p^{-})=\left(\dfrac{p^{+}+p^{-}}{2}\right)^{2}-\alpha\dfrac{p^{+}-p^{-}}{2}-1$
for some $\alpha>0$, is the Lax-Friedrichs numeical Hamiltonian associated to $H(p)=p^{2}-1$.
Note that the functional $F(\cdot)$ is differentiable in $\mathbb{R}^{N+1}$, and then, the grid function $U\in\mathbb{R}^{N+1}$ associated to any global minimizer $u\in C([0,1])$ must satisfy the first-order optimality condition $\nabla F(U)=0.$
Denoting $p^{\pm}_{i}=\pm(u_{i\pm 1}-u_{i})/\delta$, we can write the partial derivative of $F(U)$ with respect to $u_{i}$ as
$\partial_{u_{i}}F(U)$ $=$ $\widehat{H}_{\alpha}(p_{i-1}^{+},p_{i-1}^{-})\left(p_{i-1}^{+}+p_{i-1}^{-}-\alpha\right)-\widehat{H}_{\alpha}(p_{i+1}^{+},p_{i+1}^{-})\left(p_{i+1}^{+}+p_{i+1}^{-}+\alpha\right)+2\alpha\widehat{H}_{\alpha}(p_{i}^{+},p_{i}^{-}),$
for any $i=2,\ldots,N-2$.
Now, let us denote
$w_{i}=\widehat{H}_{\alpha}(p_{i}^{+},p_{i}^{-})\quad\text{and}\quad v_{i}=p_{i}^{+}+p_{i}^{-},\qquad\text{for}\ i=1,\ldots,N-1.$
Using this notation, we can write the optimality conditions for $u_{i}$ with $i=2,\ldots,N-2$ as
$-(\alpha-v_{i-1})w_{i-1}-(\alpha+v_{i+1})w_{i+1}+2\alpha w_{i}=0,\qquad i=2,\ldots,N-2.$
Further, since
$\partial_{u_{1}}F(U)$ $=$ $2\alpha\widehat{H}_{\alpha}(p_{1}^{+},p_{1}^{-})-\widehat{H}_{\alpha}(p_{2}^{+},p_{2}^{-})(p_{2}^{+}+p_{2}^{-}+\alpha)$
$=$ $-(\alpha+v_{2})w_{2}+2\alpha w_{1},$
$\partial_{u_{N-1}}F(U)$ $=$ $\widehat{H}_{\alpha}(p_{N-2}^{+},p_{N-2}^{-})(p_{N-2}^{+}+p_{N-2}^{-}-\alpha)+2\alpha\widehat{H}_{\alpha}(p_{N-1}^{+},p_{N-1}^{-})$
$=$ $-(\alpha-v_{N-2})w_{N-2}+2\alpha w_{N-1},$
we obtain the first-order optimality condition for the minimization problem (2.4), which reads as
$\begin{cases}-v_{2}w_{2}-\alpha(w_{2}-2w_{1})=0\\ v_{i-1}w_{i-1}-v_{i+1}w_{i+1}-\alpha(w_{i+1}+w_{i-1}-2w_{i})=0&\text{for}\ i=2,\ldots,N-2\\ v_{N-2}w_{N-2}-\alpha(w_{N-2}-2w_{N-1})=0.\end{cases}$ (2.5)
Denoting $W=(w_{1},\ldots,w_{N-1})$ and $V=(v_{1},\ldots,v_{N-1})$, this system of equations can be written as
$-(A_{N}(V)+\alpha\Delta_{N})W=0,$ (2.6)
where $A_{N}(V)\in\mathbb{R}^{(N-1)\times(N-1)}$ is a matrix depending on $V$ and $\Delta_{N}\in\mathbb{R}^{(N-1)\times(N-1)}$ is the Toeplitz matrix associated to the discrete Laplace operator on a uniform grid with $N-1$ points. It turns out that, for $\alpha$ and $\delta=1/N$ sufficiently big, the matrix $A_{N}(V)+\alpha\Delta_{N}$ is negative definite, which implies that the optimality system has only the trivial solution $W=0$. This implies that, if $u$ is a critical point of $\widehat{\mathcal{R}}(\cdot)$, then the associated grid function $U=(u_{0},\ldots,u_{N})$ satisfies
$w_{i}=\left(\dfrac{u_{i+1}-u_{i-1}}{2\delta}\right)^{2}-\alpha\dfrac{u_{i+1}+u_{i-1}-2u_{i}}{2\delta}-1=0,\qquad\text{for all $i=1,\ldots,N-1$.}$ (2.7)
In other words, $U$ satisfies the finite-difference equation associated with the Lax-Friedrichs numerical Hamiltonian $\widehat{H}_{\alpha}(p^{+},p^{-})$. See Section 4 for the precise estimates on $\alpha$ and $\delta$, and Section 6 for the precise computations in the multi-dimensional setup. For $\alpha$ sufficiently large, one can prove that the numerical Hamiltonian $\widehat{H}_{\alpha}(p^{+},p^{-})$ is consistent with $H(p)=p^{2}$ and monotone, as per Definition 1. Moreover, it is known (see [6, 2]) that finite-difference solutions for consistent and monotone numerical schemes converge to the viscosity solution as $\delta\to 0^{+}$. In Section 5, we propose a strategy to train a Neural Network, first with a large value of $\delta$, ensuring that any critical point solves the equation 2.7, and then progressively reducing $\delta$ to ensure convergence to the viscosity solution.
Of course, in the multi-dimensional case, a functional such as ^$\widehat{\mathcal{R}}(u)$ in (2.4) would involve a summation over a multidimensional grid. Computing the full gradient of ^$\widehat{\mathcal{R}}(u)$ would be computationally infeasible. However, in practice, one can use a version of stochastic gradient descent, in which at every gradient step, the gradient of ^$\widehat{\mathcal{R}}(u)$ is approximated by the gradient of a randomized selection of the terms in the sum.
3 Deep Learning and Partial Differential Equations
Due to its tremendous success in fields such as computer graphics and computer vision, Deep Learning has recently become a popular candidate for tackling various problems where classical methods face limitations. These include the numerical approximation of PDEs with multiple settings. Under the Deep Learning framework, one typically minimizes a functional $\mathcal{J}:C(\Omega)\to\mathbb{R}^{+}$,
${u}\in\arg\min_{u\in\mathcal{F}}\mathcal{J}(u),$ (3.1)
where $\mathcal{F}\subset C(\overline{\Omega})$ is a class of functions defined by a neural network architecture, and $D$ is a given dataset. If $\mathcal{J}$ is defined without data, the setup is often called “unsupervised” in the machine learning community. One of the reasons why Deep Learning techniques are so powerful is the fact that, for a sufficiently over-parameterized Neural Network architecture, one can prove that $\mathcal{F}$ is arbitrarily close to any function in $C(\overline{\Omega})$. Indeed, one can prove that by letting the number of parameters to infinity in a suitable way, the family of functions $\mathcal{F}$ is dense in $C(\overline{\Omega})$
Another reason for the success of Deep Learning techniques is the development of optimisation methods, which are used to address the problem (3.1). These are generally gradient-based iterative methods, such as stochastic gradient descent or variants. Hence, the choice of the functional $\mathcal{J}(u;D)$ becomes of utmost importance. For finding the viscosity solution to the boundary value problem (1.1), we consider loss functionals of the form
$\mathcal{J}(u)=\mathcal{L}(u;D)+\mathcal{R}(u).$
In [13], the popular Deep Learning framework known as Physics Informed Neural Networks (PINNs) is introduced. In a PINNs approach, $\mathcal{R}$ can be interpreted as a Monte-Carlo approximation of the squared $L^{2}$–norm of the PDE’s residual, and $\mathcal{L}$ may involve both boundary data and possibly additional “collocation” points (pointwise values of the solution in $\Omega$). When the data fitting term $\mathcal{L}$ is present, $\mathcal{R}$ is regarded as a regularisation to some practitioners.
Following the PINNs strategy, a least-square principle for our Hamilton-Jacobi equation would involve
[…]
Assuming that we do not have, in principle, access to the viscosity solution in the interior of the domain, the functional $\mathcal{L}$ can only involve the boundary data:
[…]
In practice, both terms $\mathcal{R}(u)$ and $\mathcal{L}(u;g)$ are approximated through Monte Carlo integration.
We see that any Lipschitz continuous function $u$ satisfying $H(x,\nabla u(x))=0$ for almost every $x\in\Omega$, together with the boundary condition $u(x)=g(x)$ on $\partial\Omega$, is a global minimizer of $\mathcal{J}(u)$. It is well-known that solutions of this type are not unique. This is true even for the simple Eikonal equation on a one-dimensional interval shown in Example 1. This is, of course, very inconvenient since one cannot guarantee by any means that minimizing the PINNs functional $\mathcal{J}(u)$ with (3.2) and (3.3) would produce an approximation of the viscosity solution. See Figure 1 for an illustration of four global minimizers of the PINNs functional associated with the one-dimensional Eikonal equation in $(0,1)$ with zero-boundary condition. To overcome the issue of non-uniqueness of minimisers, and promote convergence to the viscosity solution, we need to modify the functional in a way that functions with certain regularity properties are enhanced. In Deep Learning, a typical technique to achieve this is to add a regularisation term to the functional in the form of some Sobolev norm. However, this is not the only way to enhance regularity on the minimisers. In this paper, we replace the $L^{2}$-norm of the PDE-residual by the least-square principle associated to a finite-difference numerical scheme. We prove that the numerical diffusion of the finite-difference scheme enhances regularity on the minimisers. Moreover, by choosing a suitable consistent and monotone numerical scheme, we can prove that the unique minimiser approximates the viscosity solution.
3.1 Why finite-difference approximations?
Larger domain of dependence/influence.
Finite difference discretizations lead to a larger domain of dependence/influence per collocation point, compared to PDEs. Which feature can lead to a stronger coupling of the “pointwise losses” and to more efficient capturing of the problem’s underlying causality. This is essential as typical solutions of the HJ are not classical, and the singular sets of the solutions are lower dimensional; i.e. the probability of a finite set of collocation points “sampling” the singular set is proportional to the size of the domain of dependence of the residual term to be minimized.
Simpler loss gradient evaluations.
While automatic differentiation can be a very convenient and efficient tool for evaluating first-order derivatives, the evaluation of second-order derivatives can be costly with a large memory footprint. See e.g. [17] and [11], Tab. V.
Convergence to the viscosity solutions.
By using a convergent monotone numerical scheme, e.g. [6], our proposed algorithm is capable of computing the viscosity solution of the given boundary value problem, without explicitly determining the convexity of the solution or where the characteristics of problems should collide.
4 Finite difference residual functionals
Let us consider a numerical Hamiltonian
$\widehat{H}(x,p^{+},p^{-}):\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}.$
For a fixed $\delta>0$, we use the notation
$D^{+}_{\delta}u(x):=\left(\dfrac{u(x+\delta e_{1})-u(x)}{\delta},\ldots,\dfrac{u(x+\delta e_{d})-u(x)}{\delta}\right)\in\mathbb{R}^{d}$
and
$D^{-}_{\delta}u(x):=\left(\dfrac{u(x)-u(x-\delta e_{1})}{\delta},\ldots,\dfrac{u(x)-u(x-\delta e_{d})}{\delta}\right)\in\mathbb{R}^{d},$
respectively for the upwind and downwind finite-difference approximation of the gradient $\nabla u(x)$. Here, the vectors $\{e_{i}\}_{i=1}^{d}$ represent the canonical basis of $\mathbb{R}^{d}$. An important definition is in order:
Definition 1.
(Consistency and monotonicty)
  1. i.
    We say that the numerical Hamiltonian $\widehat{H}(x,p^{+},p^{-})$ is consistent if $\widehat{H}(x,p,p)\equiv H(x,p)$,   $p\in\mathbb{R}^{d}.$
  2. ii.
    We say that the numerical Hamiltonian $\widehat{H}(x,p^{+},p^{-})$ is monotone in $\Omega$ if the function
    $\left(u(x),\{u(x+\delta e_{i})\}_{i=1}^{d},\{u(x-\delta e_{i})\}_{i=1}^{d}\right)\longmapsto\widehat{H}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))$
    is non-decreasing with respect to $u(x)$ and non-increasing with respect to $u(x+\delta e_{i})$ and $u(x-\delta e_{i})$.
The Lax-Friedrichs scheme
In this paper, we work with the Lax-Friedrichs scheme, given by
$\widehat{H}_{\alpha}(x,p^{+},p^{-}):=H\left(x,\dfrac{p^{+}+p^{-}}{2}\right)-\alpha\sum_{i=1}^{d}\dfrac{p_{i}^{+}-p_{i}^{-}}{2},\qquad\text{for}\ \alpha>0.$ (4.1)
This numerical scheme is consistent with $H(u,p)$, and for any $L>0$, if one takes $\alpha$ satisfying
$\alpha\geq C_{H}(L):=\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|,$ (4.2)
then $\widehat{H}_{\alpha}(x,p^{+},p^{-})$ is monotone at any function $u$ with Lipschitz constant $L$.
For any $\delta>0$, let us define the uniform grid
$\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega,$
and we consider the loss functional
^$\widehat{\mathcal{R}}(u):=\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}.$ (4.3)
Global minimisers
We note that a function $u\in C(\overline{\Omega})$ is a global minimiser of ^$\widehat{\mathcal{R}}(u)$ if and only if
$\widehat{H}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0\qquad x\in\Omega_{\delta}.$ (4.4)
In other words, $u$ is a global minimiser of ^$\widehat{\mathcal{R}}(u)$ if and only if the restriction of $u$ to $\Omega_{\delta}$ solves the discrete equation associated with the numerical Hamiltonian $\widehat{H}$.
If $\widehat{H}$ is consistent with $H$ and monotone, one can ensure (see [6, 2]) that any minimiser of ^$\widehat{\mathcal{R}}(u)$ approximates a viscosity solution of the Hamilton-Jacobi equation
$H(x,\nabla u)=0,\qquad\text{in}\ \Omega,$ (4.5)
in the following sense: for any decreasing sequence $\{\delta_{n}\}_{n\geq 1}$ with $\delta_{n}\to 0^{+}$, and an equicontinuous sequence of functions $u_{n}\in C(\overline{\Omega})$, each of them minimising ^$\widehat{\mathcal{R}}(u)$ defined on $\Omega_{\delta_{n}}$. If $u_{n}$ converge point-wise to some function $u^{\ast}\in C(\overline{\Omega})$, then $u^{\ast}$ is a viscosity solution of (4.5).
With the above observation, we accomplished the first requirement for the functional $\mathcal{R}(\cdot)$, i.e. minimisers approximate viscosity solutions to Hamilton-Jacobi equation in (1.1). The second requirement is more subtle and is the main contribution of this paper. Namely, we provide sufficient conditions on the numerical Hamiltonian $\widehat{H}(x,p^{+},p^{-})$, ensuring that any critical point of the functional ^$\widehat{\mathcal{R}}(u)$ defined in (4.3) is indeed a global minimiser. Hence, it approximates a viscosity solution of (4.5).
Critical points
Let $N_{\delta}$ denote the number of grid nodes in $\Omega_{\delta}$. For any $u\in C(\overline{\Omega})$, let $U:=u|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}$ denote the vector corresponding to the values of $u$ on $\Omega_{\delta}$ following a chosen ordering of the grid nodes in $\Omega_{\delta}$. We define the functional $F:\mathbb{R}^{N_{\delta}}\mapsto\mathbb{R},$ given by
$F(U):=\widehat{\mathcal{R}}(u)\quad\text{for any}\leavevmode\nobreak\ u\in C(\overline{\Omega})\ \text{such that}\ U=u|_{\Omega_{\delta}}.$
In view of (1.3), the function $F(\cdot)$ is differentiable in $\mathbb{R}^{N_{\delta}}$, and we denote its gradient by $\nabla F(U)\in\mathbb{R}^{N_{\delta}}$.
Then, the Fréchet differential of $\widehat{\mathcal{R}}(\cdot)$ at any $u\in C(\overline{\Omega})$ can be written as the linear functional on $C(\overline{\Omega})$ given by
$\phi\in C(\overline{\Omega})\longmapsto\Phi\cdot\nabla F(U),\quad\text{where}\ \Phi:=\phi|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}.$
Let us define the residual function $w\in C(\overline{\Omega})$ associated to $u$ as
$w(x)=\widehat{H}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x)),\qquad\forall u\in C(\overline{\Omega}).$
From the form of $\widehat{R}(u)$ in (4.3), which is a least-squares like functional, the gradient of $F$ has the following form
$\nabla F(U)=A_{\delta}(U)W,$ (4.6)
where $W:=w|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}$, and $A_{\delta}(U)$ is a linear bounded operator $\mathbb{R}^{N_{\delta}}\to\mathbb{R}^{N_{\delta}}$, depending on $\widehat{H}$ and $U$ (see Section 6 for more details). Then, for any $u\in C(\overline{\Omega})$ critical point of $\widehat{\mathcal{R}}(\cdot)$, the vector of residuals $W$ satisfies the equation
$A_{\delta}(U)W=0.$ (4.7)
This linear system can be seen as a finite-difference equation for the residual function $w(x)$ on the grid $\Omega_{\delta}$. It can be interpreted as the adjoint equation associated with the finite difference system (4.4). Our next goal is to prove that by choosing a suitable numerical Hamiltonian $\widehat{H}$, one can ensure that the adjoint equation (4.7) has a unique solution, which is the trivial solution $W=0$. This would imply that any critical point of ^$\widehat{\mathcal{R}}(u)$ solves the finite-difference equation (4.4).
4.1 The main results
In the following theorem, for any constant $L>0$, we give a sufficient condition on $\alpha$ and $\delta$ to ensure that any critical point $u$ of $\widehat{R}(u)$ with Lipschitz constant $L$ solves the finite-difference equation (4.4). We consider the domain $\Omega$ to be a $d$-dimensional cube for simplicity. The result can be adapted to more general domains, satisfying suitable regularity properties, albeit with some additional technicalities.
Theorem 1.
Let $H$ be a Hamiltonian satisfying (1.3), and let $\Omega=(0,1)^{d},$ $d\in\mathbb{N}$. For any fixed $\alpha>0$ and $N\in\mathbb{N}$, set $\delta=1/N$ and consider the uniform Cartesian grid $\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega$ and the functional $\widehat{\mathcal{R}}(\cdot)$ defined in (4.3).
Let $u\in C(\overline{\Omega})$ be a critical point of $\widehat{\mathcal{R}}(\cdot)$ with a Lipschitz constant $L$. If
$2\alpha\sin^{2}\left(\frac{\pi}{2}\delta\right)>\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|$ (4.8)
then
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega_{\delta}.$ (4.9)
The proof is given in Section 6; essentially, it shows that if $\alpha,\delta$ and $L$ satisfy (4.8), then the linear operator $A_{\delta}(U)$ in (4.6) associated to $\widehat{H}_{\alpha}$ is invertible for any function $u$ with Lipschitz constant equal or smaller than $L$. This implies that if the first-order optimality condition (4.7) holds, then the residuals $W$ associated with the function $u$ on $\Omega_{\delta}$ satisfy $W=0$, and hence, $u$ solves the system of equations (4.9).
Remark 1.
Condition (4.8) can be written as
$\alpha\dfrac{\lambda_{1}(\Omega_{\delta})}{2d}>\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|,$
where $\lambda_{1}(\Omega_{\delta})$ is the smallest eigenvalue of the matrix associated with the discrete Dirichlet Laplacian on the computational domain $\Omega_{\delta}=\delta\mathbb{Z}\cap\Omega$. A similar result to Theorem 1 can be proved for a more general class of domains $\Omega$ satisfying suitable regularity conditions. However, we have restricted ourselves to the case of a $d$-dimensional cube for the sake of clarity of the arguments.
Since our goal is to use an artificial neural network to approximate the solution of a given Hamilton-Jacobi problem, we do not need to commit ourselves to one fixed grid as in a standard finite difference method. Therefore, we consider the functional $\mathcal{R}(\cdot):C(\overline{\Omega})\to\mathbb{R}^{+}$ given by
[…]
Note that
[…]
where ^$\widehat{\mathcal{R}}(u;z)$ is a discrete functional similar to (4.3), but on the shifted grid $\Omega_{\delta}(z):=(\delta\mathbb{Z}^{d}+z)\cap\Omega$ for $z\in(0,\delta)^{d}$.
By minimising $\mathcal{R}(\cdot)$, one solves the associated finite-difference equation on a family of shifted Cartesian grids, covering the entire domain $\Omega$ (instead of on a fixed grid). This fact is then exploited in Algorithm 2 that minimizes $\mathcal{R}(u)$ for a decreasing sequence of $\delta$.
For the functional $\mathcal{R}(u)$, we have the following result.
Theorem 2.
Let $H$ be a Hamiltonian satisfying (1.3), and let $\Omega=(0,1)^{d}$, $d\in\mathbb{N}$. For any fixed $\alpha>0$ and $N\in\mathbb{N}$, set $\delta=1/(N-1)$ and consider the functional $\mathcal{R}(\cdot)$ defined in (4.10).
Let $u\in C(\overline{\Omega})$ be any critical point of $\mathcal{R}(\cdot)$ with a Lipschitz constant $L$. If (4.8) holds, then
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega.$ (4.11)
The proof is given in Section 6.3.
We make the following remarks concerning the above results:
  1. i.
    It is worth noting that Theorems 1 and 2 are local results in the sense that they only ensure that a critical point $u\in C(\overline{\Omega})$ of $\widehat{\mathcal{R}}(\cdot)$ solves the finite-difference equation (4.9), provided that the Lipschitz constant of $u$ is smaller or equal than $L$. It does not rule out the existence of critical points with larger Lipschitz constants that do not satisfy (4.9). However, one can take $L>0$ as large as needed, at the expense of increasing $\alpha$ and $\delta$ in such a way that (4.8) holds.
  2. ii.
    Choosing large values for $\alpha$ and $\delta$ has the drawback of introducing a lot of numerical diffusion, which implies that minimisers of $\widehat{\mathcal{R}}(\cdot)$ (resp. $\mathcal{R}(\cdot))$) are over-regularised approximations of the actual viscosity solution. Since we want to use an iterative gradient-based method to minimise the functional (as it is typically done in Deep Learning), we must choose large values of $\alpha$ and $\delta$ to ensure global convergence to the minimiser. This, of course, is undesirable as the minimiser will be over-smoothed. As we explain in Section 5, and will be observed in the numerical experiments, a possible strategy to attain a global minimiser for $\alpha$ and $\delta$ small is to sequentially decrease the values of $\alpha$ and $\delta$ throughout the gradient iterations.
  3. iii.
    In Theorem 2, $\delta=1/(N-1)$, as opposed to $\delta=1/N$ in Theorem 1. This difference is only technical. If $\delta=1/N$, then the grid $\Omega_{\delta}=\delta\mathbb{Z}^{d}\cap\Omega$ contains $(N-1)^{d}$ points, whereas for any $z\in(0,\delta)^{d}$, the shifted grid $\Omega_{\delta}(z)=(\delta\mathbb{Z}^{d}+z)\cap\Omega$ has $N^{d}$ points. If $\delta=1/(N-1)$, then the grid $\Omega_{\delta}(z)$ has $(N-1)^{d}$ points, and thus one can use the same condition (4.8) from Theorem 1.
4.2 The least square principle involving numerical Hamiltonian and supervised data
The above section discussed functionals involving a numerical Hamiltonian. We show that under appropriate conditions, any global minimiser approximates a viscosity solution of the Hamilton Jacobi equation
$H(x,\nabla u)=0,\qquad\text{in}\ \Omega.$ (4.12)
The uniqueness of the viscosity solution to the partial differential equation (4.12) typically involves additional (suitable) boundary conditions. Likewise, our ultimate least square formulation has to incorporate boundary data.
For simplicity, consider $\Omega=(0,1)^{d}$ and functionals defined on the uniform Cartesian grid $\overline{\Omega}_{\delta}:=\delta\mathbb{Z}^{d}\cap\overline{\Omega}$, where $\delta=1/N$ for some $N\in\mathbb{N}$. Recall that we denote the interior points of the numerical domain by $\Omega_{\delta}:=\delta\mathbb{Z}\cap\Omega$. Similarly, we define the numerical boundary as
$\partial\Omega_{\delta}:=\delta\mathbb{Z}\cap\partial\Omega.$
Note that $\partial\Omega_{\delta}$ has to be redefined appropriately for more general domains.
For the parameters $\alpha,\gamma>0$, we define the functional
$\widehat{\mathcal{J}}(u):=\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}+\gamma\delta^{d-1}\sum_{x\in\partial\Omega_{\delta}}\left(u(x)-g(x)\right)^{2},$ (4.13)
where $\widehat{H}_{\alpha}$ is the Lax-Friedrichs numerical Hamiltonian defined in (4.1) and $g$ is the boundary condition in (1.1). The following result is a consequence of Theorem 1.
Corollary 1.
Let $H$ be a Hamiltonian satisfying (1.3), $\Omega=(0,1)^{d}$, $d\in\mathbb{N}$, and $g\in C(\partial\Omega)$. For any $N\in\mathbb{N}$, set $\delta=1/N$ and consider the uniform Cartesian grid $\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega$ with the associated numerical boundary $\partial\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\partial\Omega$. For the parameters $\alpha,\gamma>0$, consider the functional $\widehat{\mathcal{J}}(\cdot)$ defined in (4.13). Let $u\in C(\overline{\Omega})$ be any critical point of $\widehat{\mathcal{J}}(\cdot)$ with Lipschitz constant $L$. If (4.8) holds, then
$\begin{cases}\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,&x\in\Omega_{\delta},\\ u(x)=g(x),&x\in\partial\Omega_{\delta}.\end{cases}$ (4.14)
The proof of this corollary can be found in Section 6.2. Under the Assumption 1, and choosing $\alpha>0$ such that the numerical Hamiltonian $\widehat{H}_{\alpha}$ is monotone, Corollary 1 can be used in the following way: if one is able to construct an equicontinuous sequence $u_{n}\in C(\overline{\Omega})$ minimising the functional $\widehat{\mathcal{J}}(\cdot)$, with $\delta_{n}\to 0^{+}$, then $u_{n}(x)\to u^{\ast}(x)$ uniformly in $\Omega$, where $u^{\ast}$ is the unique viscosity solution to (1.1). Of course, it is important to note that condition (4.8) only holds if $\delta$ is large enough. As we will see in Section 5, one can start by training a neural network with a large value of $\delta$, ensuring that any critical point satisfies (4.14), and then re-train the neural network with decreasing values of $\delta$.
A similar result holds for more general domains $\Omega\subset\mathbb{R}^{d}$, and considering the functional
[…]
where $\mu$ is a measure on $\partial\Omega$.
In this paper, we propose to minimize such least square principles via some stochastic gradient descent algorithm. This means that the integrals will be approximated by the Monte Carlo method. The involvement of the Monte-Carlo method would require deciding on a probability distribution over $\Omega$, and the choice of this distribution can depend on the error metric of choice. Finally, in many applications, such as finding the shortest path to the target region, one also has data on $\Gamma$ that rests in the interior of $\Omega$. Therefore, we propose a more general form of the least square principle as follows:
$\mathcal{J}(u):=\int_{\Omega}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}\rho(x)dx+\gamma_{1}\int_{\partial\Omega}\left[u(x)-g(x)\right]^{2}d\mu(x)+\gamma_{2}\int_{\Gamma}\left[u(x)-h(x)\right]^{2}d\nu(x),$ (4.15)
where $\rho$ is a density on $\Omega$, $\mu$ and $\nu$ are measures on $\partial\Omega$ and $\Gamma$.
In Section 5.1, we propose two simple stochastic gradient descent algorithms for minimizing $\mathcal{J}(u)$. Section 5.2.1 presents a study on the choice of $\rho$ using two different metrics.
Hyper-parameters $\alpha$ and $\delta$ influenced by supervised data
Using our results in Theorems 1 and 2 and Corollary 1, we can deduce that adding supervised data in the functional allows the use of a numerical Hamiltonian with less numerical diffusion, while still ensuring that any critical point of $\mathcal{J}(\cdot)$ satisfies (4.14). Indeed, the set $\Gamma$ can be seen as part of the boundary of an auxiliary domain $\Omega^{\ast}:=\Omega\setminus\Gamma$. For a fixed $\delta>0$, the computational grid $\Omega^{\ast}_{\delta}:=\delta\mathbb{Z}\cap\Omega^{\ast}$ is then a subset of the original grid $\Omega_{\delta}:=\delta\mathbb{Z}\cap\Omega$, which implies that the smallest eigenvalue of the discrete Dirichlet Laplacian on $\Omega^{\ast}_{\delta}$ is larger than the one on $\Omega_{\delta}$. Then, in view of Remark 1, it follows that one is allowed to take smaller values of $\alpha$, or $\delta$, while the conclusion of Theorem 1 (resp. Theorem 2 or Corollary 1) still holds. Being able to choose smaller values of $\alpha$ and $\delta$ reduces the numerical diffusion of the solution of (4.14), which in turn implies that the minimiser of $\mathcal{J}(u)$ provides a more accurate approximation of the viscosity solution.
4.3 Uniqueness of a critical point by regularization of the Hamiltonian
In Theorem 1 (resp. Theorem 2), we provide a sufficient condition on the choice of the parameters $\alpha$ and $\delta$, ensuring that any critical point of the functional ^$\widehat{\mathcal{R}}(u)$ (resp. $\mathcal{R}(u)$) satisfies
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\qquad\forall x\in\Omega_{\delta}\ \text{(resp. $\Omega$).}$. (4.16)
This condition (see (4.8)) involves taking $\alpha$ and $\delta$ sufficiently large so that the system of equations associated with the first-order optimality condition, which is of the form $A_{\alpha,\delta}(U)W=0,$ admits a unique solution, i.e. the trivial solution $W=0$. See Section 6 for further details.
An alternative way to ensure that any critical point of the functional solves the finite-difference equation (4.16) is by adding an increasing term to the numerical Hamiltonian. Let us consider the functional
$u\in C(\overline{\Omega})\longmapsto\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))+\tau u(x)\right]^{2},$
for some $\tau>0$. In this case, the first-order optimality condition is of the form
$(A_{\alpha,\delta}(U)+\tau I_{\Omega_{\delta}})W=0,$
where $I_{\Omega_{\delta}}$ is the identity operator in the space of grid functions on $\Omega_{\delta}$, and $W$ is the grid function associated to the function $x\mapsto w(x):=\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))+\tau u(x)$.
Similar arguments to the ones in the proofs of Theorems 1 and 2 can be used to obtain a sufficient condition on the parameters $\alpha,\delta$ and $\tau$, ensuring that the matrix associated to the linear operator
$A_{\alpha,\delta}(U)+\tau I_{\Omega_{\delta}}$
is positive definite, for any grid function $U$ on $\Omega_{\delta}=\delta\mathbb{Z}\cap\Omega$ with Lipschitz constant $L$. This in turn implies that, for any function $u$ with Lipschitz constant $L$, if $u$ is a critical point of ^$\widehat{\mathcal{R}}(u)+\tau\widehat{\mathcal{I}}$ then $u$ satisfies the finite-difference equation (4.16).
4.4 Time dependent problems
There is no difficulty in extending our analysis to the time-dependent case:
$\begin{cases}\partial_{t}u+H(x,\nabla u)=0&\Omega\times(0,T),\\ u(x,t)=g(x)&\partial\Omega\times(0,T),\\ u(x,0)=u_{0}(x)&\Omega,\end{cases}$ (4.17)
with $T\in(0,\infty]$, initial data $u_{0}:\Omega\to\mathbb{R}$ and boundary data $g:\Omega\to\mathbb{R}$.
In this case, a suitable choice for the least squares approach is to use the following finite-difference scheme
$u(x,t+\delta_{t})=u(x,t)-\delta_{t}\widehat{H}_{\alpha}(x,D_{\delta_{x}}^{+}u(x,t),D_{\delta_{x}}^{-}u(x,t)),$ (4.18)
for some $\delta_{t},\delta_{x},\alpha>0$, which combines an explicit Euler scheme for the time derivative and a Lax-Friedrichs numerical Hamiltonian.
The least-squares finite-difference functional associated to the above numerical scheme is
[…]
and can be coupled with the initial and boundary data as
[…]
A similar analysis of the critical points can be carried out for the above functional. Again, the conclusion is that by taking $\alpha$ and $\delta$ sufficiently large, one can ensure that any critical point satisfies the finite-difference equation associated with the numerical scheme (4.18). In Section 5.2.2, we show some numerical experiments, where we apply our least squares approach to time evolution problems of the form (4.17).
5 Numerics
As discussed in the introduction, the purpose of formulating the numerical method for (1.1) as a minimisation problem is that one can approximate the solution using a Neural Network trained through stochastic gradient descent (SGD). To enhance converge to the viscosity solution, we use the residual of a consistent and monotone Lax-Friedrichs numerical Hamiltonian in the loss functional, which ensures that any global minimiser approximates a viscosity solution. Due to the use of stochastic gradient descent as optimisation method, it is of utmost importance that any critical point of the functional is a global minimiser. However, in view of our main results Theorem 1 and Theorem 2, we can only ensure this by considering a numerical Hamiltonian that incorporates enough numerical viscosity ($\alpha$ and $\delta$ big enough in (4.3) or (4.10)). The drawback is that the global minimiser of the resulting minimisation problem is an over-smoothed approximation of the actual viscosity solution. Next, we described an algorithm to approximate the global minimiser of a Lax-Friedrichs finite-difference loss with small parameters $\alpha$ and $\delta$. The idea is to first apply SGD with large $\alpha$ and $\delta$, ensuring convergence to the global minimiser, then decrease these parameters to carry on with more SGD iterations and refine the approximation of the solution.
5.1 The proposed algorithms
Let us consider any NN function $\Phi:\overline{\Omega}\times\Theta\to\mathbb{R}$, where $\Theta\subset\mathbb{R}^{P}$ is the parameter space and $P$ is the number of parameters. We define the parameterized family of functions
:=$\mathcal{F}:=\{x\mapsto\Phi(x;\theta)\ :\quad\theta\in\Theta\}\subset C(\overline{\Omega}).$ (5.1)
Our numerical solution of (1.1) will be a function of the form $u(x)=\Phi(x,\theta^{\ast})$, where $\theta^{\ast}\in\Theta$ is obtained by minimising the function
$\theta\longmapsto\mathcal{J}(\Phi(\cdot;\theta)),$ (5.2)
and $\mathcal{J}$ is defined in (4.15) or in (4.20). The minimisation algorithm to obtain $\theta^{\ast}$ is a version of SGD in which, at every iteration, the functional (5.3) is approximated by a Monte Carlo approximation of the integrals in both terms.
Let $\mathcal{D}_{0}$, $\mathcal{D}_{b_{1}}$, and $\mathcal{D}_{b_{2}}$ be the chosen probability distributions over $\Omega$, $\partial\Omega$, and $\Gamma$ respectively. Our algorithms are built on the typical stochastic gradient descent algorithm:
$\theta_{j+1}=\theta_{j}-\eta_{j}\nabla_{\theta}\widetilde{\mathcal{J}}(\Phi(\cdot,\theta_{j})),$ (5.3)
with
$\widetilde{\mathcal{J}}$ ::=:= $\widetilde{\mathcal{L}}+\widetilde{\mathcal{R}},$
~$\widetilde{\mathcal{L}}(\Phi(\cdot;\theta);\gamma_{1},\gamma_{2})$ ::=:= $\dfrac{\gamma_{1}}{N_{b_{1}}}\sum_{x\in X_{b_{1}}}\left(\Phi(x;\theta)-g(x)\right)^{2}+\dfrac{\gamma_{2}}{N_{b_{2}}}\sum_{x\in X_{b_{2}}}\left(\Phi(x;\theta)-h(x)\right)^{2},$
~$\widetilde{\mathcal{R}}(\Phi(\cdot;\theta))$ ::=:= $\dfrac{1}{N_{0}}\sum_{x\in X}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}.$
and $X$ is a set consisting of $N$ i.i.d. samplings from $\mathcal{D}_{0}$, $X_{b_{1}}$ contains $M_{1}$ i.i.d. samples from both $\mathcal{D}_{b_{1}}$, and $X_{b_{2}}$ contains $M_{2}$ i.i.d. samples from both $\mathcal{D}_{b_{2}}$.
A possible choice for the probability distributions $\mathcal{D}$, $\mathcal{D}_{b_{1}}$, and $\mathcal{D}_{b_{2}}$ are the uniform distributions over $\Omega$, $\partial\Omega$, and $\Gamma$ respectively. However, as we observe in our numerical experiments below, other distributions can be more efficient, especially in high-dimensional domains.
Algorithm 1 SGD for Lax-Friedrichs finite-difference loss
$\gamma_{1},\gamma_{2}>0$, $\{\eta_{j}\}_{j=0}^{J}\in(0,\infty)^{J}$, $N_{0},N_{b_{1}},N_{b_{2}}\in\mathbb{N}$, $\mathcal{D}_{0}\in\mathcal{P}(\Omega)$, $\mathcal{D}_{b_{1}}\in\mathcal{P}(\partial\Omega)$, and $\mathcal{D}_{b_{1}}\in\mathcal{P}(\Gamma)$.
Input: $\theta_{0}\in\Theta$ and $\alpha,\delta>0$.
procedure SGDLxF($\theta_{0},\alpha,\delta$)
     $j\leftarrow 0$
     while $j<J$ do
         $X\leftarrow$ $N_{0}$ i.i.d. samples from $\mathcal{D}_{0}$
         $X_{b_{1}}\leftarrow$ $N_{b_{1}}$ i.i.d. from $\mathcal{D}_{b_{1}}$
         $X_{b_{2}}\leftarrow$ $N_{b_{2}}$ i.i.d. from $\mathcal{D}_{b_{2}}$
         $\theta^{j+1}:=\theta^{j}-\eta_{j}\left(\nabla_{\theta}\widetilde{\mathcal{L}}(\Phi(\cdot;\theta^{j});\gamma_{1},\gamma_{2})+\nabla_{\theta}\widetilde{\mathcal{R}}(\Phi(\cdot;\theta^{j}))\right)$
         $j\leftarrow j+1$
     end whilereturn $\theta^{J}$
end procedure
If the NN is initialised to be close to a global minimiser, then the gradient descent iterations will converge to a global minimiser. By restricting the parameters to a compact set, we can have an estimate of the Lipschitz constant $L$. We can choose $\alpha:=\alpha_{0}>0$ and $\delta:=\delta_{0}>0$ accordingly to meet the condition set in (4.8) for the Lax-Friedrichs numerical Hamiltonian $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))$ from (4.1).
In view of Theorem 2, we can ensure that any critical point of
[…]
with Lipschitz constant $L>0$ satisfies
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0\qquad\text{forall}\quad x\in\Omega.$ (5.4)
Hence, the gradient descent iterations will converge to a function that solves (or approximately solves) the above equation. Then, we can reduce the hyper-parameters $\alpha,\delta$ by choosing
$0<\alpha:=\alpha_{1}\lesssim\alpha_{0}\quad\text{and}\quad 0<\delta:=\delta_{1}\lesssim\delta_{0}$
and continue with the gradient descent iterations applied to the new functional $\mathcal{R}_{\alpha_{1},\delta_{1}}(u)$. Due to the continuity of the critical points of $\mathcal{R}(\cdot)$, and of the solutions of (5.4) with respect to $\alpha$ and $\delta$, we can ensure that, by initializing the NN close to a global minimiser of $\mathcal{R}_{\alpha_{0},\delta_{0}}(u)$, the iterations will converge to a global minimiser of $\mathcal{R}_{\alpha_{1},\delta_{1}}(u)$. This procedure can be iterated with a reduction schedule for $\alpha$ and $\delta$. Every time we reduce the parameters $\alpha$ and $\delta$, we obtain a function which is closer to the viscosity solution of (1.1). This procedure can be interpreted as an approximation of the vanishing viscosity method, in which the viscosity solution to the boundary value problem (1.1) is obtained as the limit $\varepsilon\to 0^{+}$ of the viscous approximation (1.4).
Algorithm 2 Optimisation for a Lax-Friedrichs finite-difference loss with a progressive refinement schedule
$K\in\mathbb{N}$, $L>0$, $\mathcal{F}$ as in (5.1) and $\{(\alpha_{k},\delta_{k})\}_{k=0}^{K}\in(0,\infty)^{2K}$.
$(L,\alpha_{0},\delta_{0})$ satisfy (4.8) and $\alpha_{k+1}<\alpha_{k}$ and $\delta_{k+1}<\delta_{k}$ for all $1\leq k\leq K$.
Initialise $\theta_{0}\in\Theta$ in a way that $x\mapsto\Phi(x;\theta)$ has Lipschitz constant $L$.
$k\leftarrow 0$
while $k<K$ do
     $\theta_{k+1}:=\operatorname{SGDLxF}(\theta_{k},\alpha_{k},\delta_{k})$ as defined in Algorithm 1
     $k\leftarrow k+1$
end whilereturn $\theta_{K}$
5.2 Computational experiments
Next, we test the efficacy of our algorithm on various canonical HJ equations across different dimensions. We start with the eikonal equation posed on different domains (Section 5.2.1). Then, we consider a time-evolution HJ equation of the form (4.17) (Section 5.2.2). Finally, we consider HJ equations arising in control problems and differential games associated to curvature constrained dynamics. In particular, we consider two problems associated to the Reeds-Shepp’s car model: the shortest path to a given target (Section 5.2.3) and the pursuit evasion game (Section 5.2.4).
In all the experiments, we use a fully connected multi-layer perceptron as NN architecture. For a chosen number of hidden layers $\ell$ (NN depth) and number of neurons per layer $n$ (NN width), this can be written as
$\Phi(x;\theta):=W_{\ell}\sigma\left(W_{\ell-1}x\sigma\left(\cdots\sigma(W_{0}x+b_{0})\right)+b_{\ell-1}\right)+b_{\ell},$ (5.5)
where $\theta:=\{(W_{i},b_{i})\}_{i=0}^{\ell}$ is the parameter of the NN. The matrices $W_{0}\in\mathbb{R}^{d\times n}$, $\{W_{i}\}_{i=1}^{\ell-1}\in\left(\mathbb{R}^{n\times n}\right)^{\ell-1}$ and $W_{\ell}\in\mathbb{R}^{n\times 1}$ are typically called weights and the vectors $\{b_{i}\}_{i=0}^{\ell-1}\in\left(\mathbb{R}^{n}\right)^{\ell}$ and the scalar $b_{\ell}\in\mathbb{R}$ are called biases. The function $\sigma(\cdot)\in C(\mathbb{R};\mathbb{R}^{+})$ defined as $\sigma(s):=\max\{0,s\}$ is the activation function, which acts component wise on vectors.
The NN function can also be written as
$\Phi(x;\theta)=W_{\ell}y_{\ell}+b_{\ell},\quad\text{where}\quad\begin{cases}y_{i}=\sigma\left(W_{i-1}y_{i-1}+b_{i-1}\right),&\text{for}\ i=1,\ldots,\ell\\ y_{0}=x.\end{cases}$
In the experiments, we vary the number of hidden layers and the number of neurons per layer depending on the complexity of the problem. For each case, this will be indicated in the experiment description. For the problems in Sections (5.2.3) and (5.2.4), we consider a slightly different architecture, due to the fact that the problem’s domain requires the solution to be $2\pi$-periodic with respect to some of the variables. This is achieved by considering a linear combination of trigonometric functions, for which the coefficients are given by NN as defined above.
We have not explored other NN architectures. Depending on the specific problem under consideration, the use of other NN architectures such as ResNet might improve the performance of the method, however, this is left for future research.
5.2.1 Eikonal equations
To illustrate the algorithm presented above, we carry out some numerical experiments for the Eikonal equation in a $d$-dimensional cube
$\begin{cases}|\nabla u(x)|^{2}=f(x)&x\in\Omega\\ u(x)=g(x)&x\in\partial\Omega.\end{cases}$ (5.6)
Let us start with a simple example, taking $f\equiv 1$ and $g\equiv 0$, and the domain $\Omega:=(-3,3)^{d}$. The solution, in this case, corresponds to the distance function to the boundary of the $d$-dimensional cube and can be written as
$u(x)=3-\|x\|_{\infty},\qquad\forall x\in(-3,3)^{d}.$ (5.7)
We will use the explicit solution as a ground truth to evaluate the accuracy of the numerical approximation in each experiment.
We first consider the two-dimensional case. We use Algorithm 2, with $K=5$, to train a two-layer fully connected Neural Network with 20 neurons in the hidden layer. The parameters $\alpha$ and $\delta$ in each round/iteration are taken as
$\alpha=[2.5,2,1.5,1,0.5]\quad\text{and}\quad\delta=[0.75,0.5,0.3,0.1,0.05].$ (5.8)
Observe that, after each round, both parameters $\alpha$ and $\delta$ are decreased. This is designed to progressively reduce the effect of the numerical viscosity to obtain higher fidelity in the approximate solution in the final round. In each round, we applied Algorithm 1 with $J=1000$ steps. The number of supervised data on the boundary $\partial\Omega$ in (5.3) is taken as $N_{b_{1}}=30$, which are uniformly sampled from the domain boundary for each iteration. Concerning the number of collocation points, we increase the batch size by decreasing the discretisation step’s size in the numerical Hamiltonian. Namely, we consider the values
$N_{0}=[30,40,80,100,1000].$ (5.9)
The collocation points are sampled from a uniform distribution on $\Omega$ for each iteration of Algorithm 1. The need to increase the number of collocation points is because when decreasing $\delta$, each collocation point covers a smaller portion of the domain, and hence, more collocation points are needed. This may be another argument supporting the use of a finite-difference-based loss function over the PDE residual, which would correspond to the limit $\delta\to 0^{+}$.
In Figure 2, we can see the numerical approximation provided by the NN after each of the five rounds in Algorithm 2. In Figure 3, we see the evolution throughout the rounds of the Mean Square Error and the $L^{\infty}$-Error (see (5.10)) of the numerical solution with respect to the ground truth solution (5.7). Each round of the algorithm took one second on the CPU, except for the last round, which, due to the large number of collocation points, took around two seconds. Thus, the NN’s training took a total of six seconds.
Figure 2: Neural network solutions of the Eikonal equation in $\Omega=(-3,3)^{2}$ constructed by Algorithm 1. Each plot corresponds to the approximation after every iteration. The values of $\alpha$ and $\delta$ are taken as in 5.8.
Figure 3: Evolution of the Mean Square Error and the $L^{\infty}$-Error after each round of Algorithm 2, for the Eikonal equation in $(-3,3)^{2}$. The parameters $\alpha$ and $\delta$ in each of the five iterations are taken as in (5.8).
Next, we consider the same problem in dimensions $d=5$ and $d=10$. In each experiment, we conduct five rounds in the Algorithm 2, again with decreasing parameters $\alpha$ and $\delta$ as in (5.8). As we increase the dimension, we need to increase the complexity of the NN and the sampling size for the iterations in Algorithm 1. Consequently, we also need to increase the number of iterations per round (we chose $J=3000$ for the $5$-dimensional case and $J=5000$ for the $10$-dimensional case). We used a three-layer, fully connected NN. The complexity of the NN is increased by increasing the number of neurons in the hidden layers, i.e. the width of the NN.
See Tables 1 and 2 for a summary of the numerical experiments (each row corresponds to an independent experiment). In all the experiments, we used a uniform distribution to sample the points in $\Omega$ and $\partial\Omega$ during the iterations of Algorithm 1. In Table 1, we report the Mean Square Error of the numerical solution after each round, and in Table 2, we can see the $L^{\infty}$-Error. These errors are defined respectively as
[…]
where $u\in C(\overline{\Omega})$ is the viscosity solution, which, in this case, is given by (5.7).
Table 1: Mean Square Error of the numerical solution for the Eikonal equation in $(-3,3)^{d}$ with $d=5,10$ after each round of Algorithm 2. In each experiment we carried out $5$ rounds with parameters $\alpha$ and $\delta$ as in (5.8). We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
$d$ $n$ $N_{0}$ $N_{b_{1}}$ $MSE\left(\theta_{1}\right)$ $MSE\left(\theta_{2}\right)$ $MSE\left(\theta_{3}\right)$ $MSE\left(\theta_{4}\right)$ $MSE\left(\theta_{5}\right)$ Runtime(s)
5 20 60 20 0.074888 0.032949 0.017251 0.002257 0.002136 28.834377
5 20 100 50 0.074242 0.031021 0.008205 0.001884 0.001798 30.431862
5 20 300 100 0.079349 0.028334 0.008436 0.004028 0.003564 40.061352
5 40 60 20 0.080704 0.034279 0.009987 0.003943 0.007453 32.218208
5 40 100 50 0.073138 0.031448 0.007935 0.002752 0.003224 34.528339
5 40 300 100 0.074896 0.025999 0.007483 0.002128 0.006383 46.840443
10 30 80 30 0.433648 0.297972 0.312044 0.273188 0.194381 65.347862
10 30 200 60 0.426775 0.372508 0.188064 0.115416 0.046867 91.703835
10 30 400 80 0.364606 0.206969 0.072394 0.046537 0.032831 107.685104
10 60 80 30 0.401053 0.250574 0.069570 0.056949 0.029952 76.941849
10 60 200 60 0.351684 0.223897 0.063929 0.047159 0.035578 115.046974
10 60 400 80 0.367076 0.219113 0.045828 0.031053 0.028625 187.062515
Table 2: $L^{\infty}$-Error of the numerical solution for the Eikonal equation in $(-3,3)^{d}$ with $d=5,10$ after each round of Algorithm 2. In each experiment we carried out $5$ rounds with parameters $\alpha$ and $\delta$ as in (5.8). We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
$d$ $n$ $N_{0}$ $N_{b_{1}}$ $E_{\infty}(\theta_{1})$ $E_{\infty}(\theta_{2})$ $E_{\infty}(\theta_{3})$ $E_{\infty}(\theta_{4})$ $E_{\infty}(\theta_{5})$ Runtime(s)
5 20 60 20 1.417063 0.899929 0.563600 0.295139 0.196418 28.834377
5 20 100 50 1.404727 0.917647 0.468370 0.208034 0.145593 30.431862
5 20 300 100 1.424190 0.915070 0.477517 0.338038 0.298342 40.061352
5 40 60 20 1.419317 0.937307 0.467741 0.284296 0.306632 32.218208
5 40 100 50 1.429622 0.914904 0.450887 0.235697 0.229808 34.528339
5 40 300 100 1.403689 0.924961 0.453269 0.234614 0.265901 46.840443
10 30 80 30 1.484837 1.284743 1.702758 1.825397 1.558153 65.347862
10 30 200 60 1.447259 1.162474 1.003959 0.971083 0.494156 91.703835
10 30 400 80 1.424744 1.096113 0.710787 0.598849 0.509197 107.685104
10 60 80 30 1.435504 1.137931 0.865097 0.855500 0.611055 76.941849
10 60 200 60 1.403080 1.038334 0.877222 0.787736 0.603456 115.046974
10 60 400 80 1.405046 1.006007 0.744690 0.655944 0.520338 187.062515
As it turns out, especially when the dimension is high, the sampling strategy used during the iterations of Algorithm 1 dramatically impacts the accuracy of the numerical solution. Let the domain $\Omega$ be a $d$-dimensional ball. If we choose a uniform distribution over $\Omega$, as we did in the previous experiments, it is well-known that the sampled points will be closer to the boundary with higher probability. This is simply because the outer rings of the ball occupy a larger volume (of the order $\sim r^{d-1}$). This causes that the accuracy near the centre of the ball is poor as compared to the accuracy of the solution near the boundary. This issue can be overcome by using a different distribution for the sampling. In the case when $\Omega$ is a $d$-dimensional ball of radius $R$, one possibility is to use a radially uniform sampling, i.e.
$X=rv,\quad\text{where}\ r\sim\operatorname{Unif}(0,R)\ \text{and}\ v\sim\operatorname{Unif}(\mathbb{S}^{d-1}),$ (5.11)
where $\mathbb{S}^{d-1}$ denotes the $(d-1)-$dimensional unit sphere.
In Tables 3 and 4, we see a summary of the numerical experiments for the Eikonal equation in a $10$-dimensional ball of radius $4$. Again, for each experiment we carried out $5$ rounds of Algorithm 2 with parameters $\alpha$ and $\delta$ as in 5.8. In each round, the number of iterations of Algorithm 1 is $J=5000$. We observe that when a uniform distribution is used to sample the points in $\Omega$, the accuracy of the numerical solution is slightly better in terms of Mean Square Error (see Table 3). However, if we look at the accuracy in terms of $L^{\infty}$-Error (see Table 4), the accuracy is remarkably better when using the radially uniform sampling defined in (5.11). This is because, when using uniform sampling, most of the points are close to the boundary, and the accuracy near the centre of the ball is poorer. See Figure 4 for a representation of the numerical solution using different sampling distributions for the Eikonal equation in the $10$-dimensional ball $\Omega=B_{10}(0,4)$. In the plot at the left, we used a uniform sampling for the collocation points during the optimisation algorithm, whereas in the plot at the right, we used the radially uniform sampling described in (5.11). In both cases, the plots represent the NN evaluated on a uniform grid of the two-dimensional cross-section $B_{2}(0,4)\times\{0\}^{8}$.
Table 3: Mean Square Error of the numerical solution for the Eikonal equation in the $10$-dimensional ball of radius $4$. We compare the use of a uniform sampling (Unif.) in $\Omega=B_{10}(0,4)$ and a radially uniform sampling (Rad.) during Algorithm 1. We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
Sampl. $n$ $N_{0}$ $N_{b_{1}}$ $MSE\left(\theta_{1}\right)$ $MSE\left(\theta_{2}\right)$ $MSE\left(\theta_{3}\right)$ $MSE\left(\theta_{4}\right)$ $MSE\left(\theta_{5}\right)$ Runtime(s)
Unif. 30 200 60 0.110328 0.061621 0.027269 0.011736 0.004897 73.800888
Unif. 30 400 80 0.118788 0.070565 0.031719 0.007372 0.004355 98.934196
Unif. 60 200 60 0.109524 0.057619 0.022191 0.006970 0.001976 106.323158
Unif. 60 400 80 0.114257 0.063988 0.021899 0.008254 0.002056 182.277634
Rad. 30 200 60 0.116404 0.067086 0.033722 0.009239 0.005982 78.508209
Rad. 30 400 80 0.128343 0.080259 0.036268 0.025176 0.006872 101.379951
Rad. 60 200 60 0.121584 0.067197 0.036193 0.012041 0.011173 99.182808
Rad. 60 400 80 0.114929 0.064326 0.026376 0.011320 0.003026 181.791611
Table 4: $L^{\infty}$-Errors of the numerical solutions for the Eikonal equation in $\Omega=B_{10}(0,4)$, the $10$-dimensional ball of radius $4$. The numerical approximations are computed using Algorithm 1 using uniform sampling (Unif.) in $\Omega$ and a radially uniform sampling (Rad.) during Algorithm 1. We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
Samp. $n$ $N_{0}$ $N_{b_{1}}$ $E_{\infty}(\theta_{1})$ $E_{\infty}(\theta_{2})$ $E_{\infty}(\theta_{3})$ $E_{\infty}(\theta_{4})$ $E_{\infty}(\theta_{5})$ Runtime(s)
Unif. 30 200 60 3.234767 2.765964 2.106851 1.739295 1.332692 73.800888
Unif. 30 400 80 3.216162 2.713652 2.118615 1.699433 1.266500 98.934196
Unif. 60 200 60 3.215516 2.758886 2.189581 1.706580 1.348982 106.323158
Unif. 60 400 80 3.227854 2.743320 2.109617 1.654431 1.295758 182.277634
Rad. 30 200 60 3.232331 2.719774 1.956412 1.243235 0.330742 78.508209
Rad. 30 400 80 3.247878 2.749383 1.975342 1.289011 0.316167 101.379951
Rad. 60 200 60 3.238025 2.730297 1.964284 1.233331 0.326179 99.182808
Rad. 60 400 80 3.231674 2.714149 1.933717 1.226952 0.247510 181.791611
Figure 4: The numerical solution for the Eikonal equation in $\Omega=B_{10}(0,4)\subset\mathbb{R}^{10}$, evaluated on a uniform grid of the two-dimensional cross-section $B_{2}(0,4)\times\{0\}^{8}$. In the figure at the left, a uniform distribution over $\Omega$ has been used in Algorithm 1, whereas in the figure at the right, we used the radially uniform sampling defined in 5.11.
In Tables 5 and 6, we can see the results of some numerical experiments in terms of MSE and $L^{\infty}$-Error, for the solution of the Eikonal equation in an annulus of dimension $d=15$ and $20$, with inner radius $1$ and outer radius $4$, i.e.
$\Omega:=\{x\in\mathbb{R}^{d}\ :\quad 1<|x|<4\}.$
We applied $5$ rounds of Algorithm 2 with $\alpha$ and $\delta$ as in (5.8). We carried out $5000$ iterations of Algorithm 1 in each round. See Figure 5 for a representation of the numerical solution in a central cross-section of the domain.
Table 5: Mean Square Error of the numerical solution for the Eikonal equation in a multi-dimensional annulus. We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
d $n$ $N_{0}$ $N_{b_{1}}$ $MSE\left(\theta_{1}\right)$ $MSE\left(\theta_{2}\right)$ $MSE\left(\theta_{3}\right)$ $MSE\left(\theta_{4}\right)$ $MSE\left(\theta_{5}\right)$ Runtime(s)
15 30 200 60 0.072798 0.057440 0.057066 0.023764 0.014624 92.204529
15 30 400 100 0.100825 0.058277 0.050973 0.035672 0.018985 148.414368
15 60 200 60 0.071785 0.047736 0.019479 0.011221 0.009534 160.298050
15 60 400 100 0.069589 0.044236 0.025286 0.008536 0.006582 299.889526
20 40 400 100 0.071917 0.041953 0.041093 0.019704 0.012875 235.390549
20 40 600 150 0.063643 0.035669 0.032586 0.017674 0.012418 340.445892
20 80 400 100 0.060304 0.039937 0.030478 0.012338 0.008163 496.888367
20 80 600 150 0.066939 0.039615 0.021738 0.013110 0.008335 746.963623
Table 6: $L^{\infty}$-error of the numerical solution for the Eikonal equation in a multi-dimensional annulus. We used the NN architecture defined in (5.5), with different number of neurons per layer $n$.
d $n$ $N_{0}$ $N_{b_{1}}$ $E_{\infty}(\theta_{1})$ $E_{\infty}(\theta_{2})$ $E_{\infty}(\theta_{3})$ $E_{\infty}(\theta_{4})$ $E_{\infty}(\theta_{5})$ Runtime(s)
15 30 200 60 1.217133 1.057912 0.913309 0.673560 0.659246 92.204529
15 30 400 100 1.301586 1.067243 0.833179 0.749360 0.620970 148.414368
15 60 200 60 1.212686 0.990640 0.713578 0.558775 0.402660 160.298050
15 60 400 100 1.205524 0.974108 0.744321 0.568417 0.374233 299.889526
20 40 400 100 1.356932 1.118282 0.945220 0.649021 0.624245 235.390549
20 40 600 150 1.312017 1.084736 0.934323 0.631151 0.564639 340.445892
20 80 400 100 1.297654 1.101432 0.823067 0.526710 0.366261 496.888367
20 80 600 150 1.306950 1.101824 0.807403 0.509917 1.310886 746.963623
Figure 5: Cross-section of the numerical solution provided by a trained NN for the Eikonal equation in a $20$-dimensional annulus. The NN approximates the viscosity solution of the Eikonal equation in the annulus.
5.2.2 Time-dependent problems
As mentioned in Section 4.4, our method can be applied to time-evolution Hamilton-Jacobi of the form (4.17). Here, we consider the initial value problem (4.17), in the domain $\Omega:=\mathbb{R}^{d}\times(0,\infty)$ with Hamiltonian
$H(x,p):=\dfrac{|p|^{2}}{2}+\dfrac{|x|^{2}}{2},\qquad(x,p)\in\mathbb{R}^{d}\times\mathbb{R}^{d},$
and initial condition
$u_{0}(x)=\dfrac{1}{2}\left(\langle x,\,Ax\rangle-1\right),$
[…]
We consider the bounded domain $\Omega:=(-3,3)^{d}\times(0,1/2)$, and minimize the functional
[…]
where $\mathcal{R}(u)$ is defined as in (4.19). Note that we do not use supervised data for the entire boundary of $\Omega$. We only prescribe the initial condition on $(-3,3)^{d}\times\{0\}$. However, due to the convexity of Hamiltonian with respect to $p$ and of the initial condition $u_{0}(\cdot)$, the characteristic curves starting outside $(-3,3)^{d}$ never enter in $(-3,3)^{d}$. This implies that the viscosity solution in $(-3,3)^{d}\times(0,T)$ is uniquely determined by the initial condition in $(-3,3)^{d}$, i.e. the initial value problem
$\begin{cases}\partial_{t}u+H(x,\nabla u)=0&(x,t)\in(-3,3)^{d}\times(0,1/2)\\ u(x,0)=u_{0}(x)&x\in(-3,3)^{d}\end{cases}$ (5.12)
has a unique solution. Then, by choosing a finite-difference functional $\mathcal{R}(u)$ such that any critical point approximates a viscosity solution, one can deduce that any critical point of $\mathcal{J}(u)$ approximates the unique viscosity solution to the initial value problem (5.12).
We run two experiments with $d=2$ and with $d=5$, in which the viscosity solution of (5.12) is a approximated by a three-layer NN with $50$ and $150$ neurons per layer respectively. We run $4$ iterations of Algorithm 2 with decreasing values of $\alpha$ and $\delta$ for each experiment. The results are reported in Table 7. Figure 6 shows the evolution of the 00-level set of the approximated solution for the two-dimensional case (left) and for the five-dimensional case (right).
Table 7: Summary of the experiments for time-dependent problems in dimension $d=2$ and $d=5$. We used the NN architecture defined in (5.5) with $\ell=3$, and $n=50$ and $n=200$ respectively.
d $n$ $\alpha$ $\delta_{x}$ $\delta_{t}$ $N_{0}$ $N_{b_{1}}$ SGD iter. $\mathcal{R}(u)$ $\mathcal{L}(u)$ Runtime
2 50 2.5 0.5 0.05 40 40 5000 0.0205 0.0025 10 sec
2 0.3 0.03 80 40 5000 0.0247 0.0017 10 sec
1.5 0.2 0.05 120 40 5000 0.0258 0.0019 11 sec
1 0.1 0.01 200 40 8000 0.0262 0.0013 19 sec
5 200 2.5 0.5 0.05 250 200 20000 0.0484 0.0111 3 min 58 sec
2 0.3 0.03 250 200 20000 0.0303 0.0033 4 min 2sec
1.5 0.2 0.02 250 200 20000 0.0362 0.0037 4 min 4sec
1 0.1 0.01 250 200 50000 0.0160 0.0025 10 min 9 sec
Figure 6: Two numerical solutions to the time-dependent problem, in dimension $d=2$ (left) and $d=5$ (right), defined in Section 5.2.2. In the plots, the zero-level sets of the approximate solutions, $\{x\in\mathbb{R}^{d}\,:\,\Phi(x,t;\theta)=0\}$, are depicted for $t=0,0.1,0.2,0.3,0.4$ and $0.5$. The label for the curves reveal the corresponding values of $t$. In the 5-dimensional case (right), the plot represents the central cross-section of the domain $(-3,3)^{2}\times\{0\}^{3}$. See Table 7 for a summary of the experiments.
5.2.3 Shortest path to a target for a Reeds-Shepp’s car
Here, we consider an optimal control problem for a dynamical system with curvature constraints. In particular, we consider the Reeds and Shepp’s car model, which can move backwards and forward [14] with a limited turning radius. The pose of the car is represented by $(x,y,\omega)\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$, where $(x,y)$ represents the spatial position of the centre of mass of the car, and $\omega$ represents its orientation angle. Here, $\mathbb{T}_{[0,2\pi)}$ denotes the one-dimensional flat torus of length $2\pi$. The optimal control problem involves steering the car to the origin $(x,y)=0$ from the given initial position $(x_{0},y_{0},\omega_{0})$ in the shortest possible time.
Denoting by $t\mapsto(x(t),y(t),\omega(t))\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$ the trajectory of the car, the dynamics are given by the following ODE:
$\begin{cases}\dot{x}(t)=\sigma a(t)\cos\omega(t)\\ \dot{y}(t)=\sigma a(t)\sin\omega(t)\\ \dot{\omega}(t)=\dfrac{b(t)}{\rho}\\ x(0)=x_{0},\ y(0)=y_{0},\ \omega(0)=\omega_{0},\end{cases}$ (5.13)
where $\sigma>0$ is the maximum speed of the car, $\rho>0$ is the inverse of the angular velocity of the car (the turning radius is therefore $\sigma\rho$) and $(x_{0},y_{0},\omega_{0})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$ is the initial position of the car. The time dependent functions $a(\cdot),b(\cdot):[0,\infty)\to[-1,1]$ are the controls of the car.
Given the dynamics of the car, the Hamiltonian associated with the shortest path problem is given by
$H(x,y,\omega,\nabla u)=\sigma|\partial_{x}u\cos\omega+\partial_{y}u\sin\omega|+\dfrac{1}{\rho}|\partial_{\omega}u|-1,$
where $\nabla u=(\partial_{x}u,\partial_{y}u,\partial_{\omega}u)$ represents the gradient of $u:\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}\to\mathbb{R}$ with repect to the position of the game. We consider the target to be the ball of radius $r=0.2$ centred at the origin. Since our framework applies to bounded domains, we set $R=5$, and define the domain
$\Omega:=\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)},$
where $\mathbb{A}_{r,R}$ denotes the two-dimensional annulus with inner radius $r$ and outer radius $R$
$\mathbb{A}_{r,R}:=\{(x,y)\in\mathbb{R}^{2}\,:\ r<\sqrt{x^{2}+y^{2}}<R\}.$ (5.14)
As for the boundary conditions, we set 00 on the inner boundary of the annulus and $R$ on the outer boundary. Hence, the resulting boundary value problem reads as
$\begin{cases}H(x,y,\omega,\nabla u)=0&(x,y,\omega)\in\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=0&(x,y,\omega)\in\{\sqrt{x^{2}+y^{2}}=r\}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=R&(x,y,\omega)\in\{\sqrt{x^{2}+y^{2}}=R\}\times\mathbb{T}_{[0,2\pi)}.\end{cases}$ (5.15)
We used our approach to approximate the viscosity solution of the above boundary value problem using a NN. Since the solution $u(x,y,\omega)$ is $2\pi$-periodic with respect to $\omega$, we considered a parameterised family of functions of the form
$\Phi(x,y,\omega;\theta)=\sum_{n=0}^{10}\phi(x,y;\theta_{n}^{(1)})\cos(n\omega)+\sum_{m=1}^{10}\phi(x,y;\theta_{m}^{(2)})\sin(m\omega),$
where the function $\phi(\cdot,\cdot;\theta):\mathbb{R}^{2}\to\mathbb{R}$ is a three-layer fully connected NN with 60 neurons in each hidden layer. We trained the NN for three rounds of Algorithm 2, using the following values for the parameters $\alpha$ and $\delta$:
$\alpha=[2.5,\,2.5,\,2.5]\quad\text{and}\quad\delta=[0.75,0.6,0.3].$
In each application of Algorithm 1, we run $3000$ iterations of stochastic gradient descent, using $N_{b_{1}}=100$ boundary points for the supervised data, and in each of the three rounds, we took $N=[800,1000,1200]$ as number of collocation points to approximate the finite-difference loss. The time it took to train the NN was $90$ seconds on a CPU. Figure 7 shows some trajectories obtained by the trained Neural Network.
Figure 7: Five sample trajectories for the Reeds-Shepp’s car, obtained using Neural Network approximating the viscosity solution of (5.15). The underlying optimal control problem is the shortest path problem to the origin (represented by the blue dot), subject to the curvature-constrained dynamics in (5.13). The red star represents the initial position of the car, and the blue circle represents the turning circumference from the initial position and orientation of the car. The bottom left figure is the fifth trajectory, zoomed-in around the initial position. The car maneuvers to find a shorter path to the target.
5.2.4 Pursuit-evasion game with Reeds-Shepp’s cars
As an example of applying our method to a Hamilton-Jacobi equation arising from differential game theory, we consider the pursuit-evasion game for two players that move according to Reeds and Shepp’s car model (5.13). Given the initial poses for the Evader and the Pursuer, represented by
$(x_{e},y_{e},\omega_{e})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)},\qquad\text{and}\qquad(x_{p},y_{p},\omega_{p})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$
We consider the game in which the Evader tries to intercept the Pursuer in the shortest possible time and delay the interception for as long as possible.
Since the position of the game only depends on the relative position between the players, we introduce the variables $X=x_{e}-x_{p}$ and $Y=y_{e}-y_{p}$. The vector $(X,Y)$ represents the vector joining the Evader with the Pursuer. The position of the game is therefore represented by $(X,Y,\omega_{e},\omega_{p})\in\mathbb{R}^{2}\times\mathbb{T}_{[1,2\pi)}^{2}$. In view of the dynamics of the two cars in (5.13), we can write down the dynamics for the evolution of the game as
$\begin{cases}\dot{X}(t)=\sigma_{e}a_{e}(t)\cos\omega_{e}(t)-\sigma_{p}a_{p}(t)\cos\omega_{p}(t)\\ \dot{Y}(t)=\sigma_{e}a_{e}(t)\sin\omega_{e}(t)-\sigma_{p}a_{p}(t)\sin\omega_{p}(t)\\ \dot{\omega}_{e}(t)=\dfrac{b_{e}(t)}{\rho_{e}},\qquad\dot{\omega}_{p}(t)=\dfrac{b_{p}(t)}{\rho_{p}}\\ X(0)=x_{e}-x_{p},\ Y(0)=y_{e}-y_{p},\ \omega_{e}(0)=\omega_{e},\ \omega_{p}(0)=\omega_{p},\end{cases}$ (5.16)
where $\sigma_{e}$ and $\sigma_{p}$ represent the maximum velocities for the Evader and the Pursuer, respectively, and $\rho_{e}$ and $\rho_{p}$ represent the inverse of their angular velocity (the turning radius for the evader and the pursuer is $\sigma_{e}\rho_{e}$ and $\sigma_{p}\rho_{p}$ respectively). The functions $a_{e}(\cdot),a_{p}(\cdot):[0,\infty)\to[-1,1]$ are the controls for the velocities of the players, and $b_{e}(\cdot),b_{p}(\cdot):[0,\infty)\to[-1,1]$ are the controls for the direction.
In this case, we define the domain of the game as
$\Omega:=\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}^{2},$
with $r=0.2$, $R=4$, where $\mathbb{A}_{r,R}$ is the two-dimensional annulus defined in (5.14), and $\mathbb{T}_{[0,2\pi)}^{2}$ is the two-dimensional flat torus with side-length $2\pi$. Recall that the position of the game is represented by the vector joining the players $(X,Y)=(x_{e}-x_{p},\,y_{e}-y_{p})\in\mathbb{R}^{2}$ and the orientation angle of the players $(\omega_{e},\omega_{p})\in\mathbb{T}_{[0,2\pi)}^{2}$. The target for the Pursuer is the inner boundary of the annulus $\mathbb{A}_{r,R}$, where we set zero boundary condition. The target for the Evader is the outer boundary, where we set the boundary condition equal to $10$.
In view of the dynamics of the game (5.16), and since the Pursuer is minimizing the output of the game and the Evader is maximizing it, we can derive the Hamiltonian associated with the game, which reads as
$H(X,Y,\omega_{e},\omega_{p},\nabla u):=\sigma_{p}\left|\partial_{X}u\cos\omega_{p}+\partial_{Y}u\sin\omega_{p}\right|+\dfrac{1}{\rho_{p}}\left|\partial_{\omega_{p}}u\right|-\sigma_{e}\left|\partial_{X}u\cos\omega_{e}+\partial_{Y}u\sin\omega_{e}\right|-\dfrac{1}{\rho_{e}}\left|\partial_{\omega_{e}}u\right|,$
where, for a function $u:\Omega\to\mathbb{R}$ defined on the game domain, $\nabla u=(\partial_{X}u,\partial_{Y}u,\partial_{\omega_{e}}u,\partial_{\omega_{p}}u)$ denotes the gradient of $u$ with respect to the position of the game. The boundary value problem associated with the game reads as follows.
$\begin{cases}H(X,Y,\omega_{e},\omega_{p},\nabla u)=0&(X,Y,\omega_{e},\omega_{p})\in\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}^{2}\\ u(x,y,\omega)=0&(x,y,\omega)\in\{\sqrt{X^{2}+Y^{2}}=r\}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=R&(x,y,\omega)\in\{\sqrt{X^{2}+Y^{2}}=R\}\times\mathbb{T}_{[0,2\pi)}.\end{cases}$ (5.17)
The viscosity solution of this boundary value problem characterises the value of the game and can be used to characterise the optimal feedback strategies for the players as
$a_{p}(X,Y,\omega_{e},\omega_{p})=-\operatorname{sgn}\left(\partial_{X}u\cos\omega_{p}+\partial_{Y}u\sin\omega_{p}\right)\qquad b_{p}=-\operatorname{sgn}\left(\partial_{\omega_{p}}u\right)$ (5.18)
for the Pursuer, and
$a_{e}(X,Y,\omega_{e},\omega_{p})=\operatorname{sgn}\left(\partial_{X}u\cos\omega_{e}+\partial_{Y}u\sin\omega_{e}\right)\qquad b_{e}=\operatorname{sgn}\left(\partial_{\omega_{e}}u\right)$ (5.19)
for the Evader.
For different values of the players’ velocities $(\sigma_{e},\rho_{e})$ and $(\sigma_{p},\rho_{p})$, we used our approach to approximate the viscosity solution of (5.17) using a NN. Similarly to the previous example, since the solution $u(X,X,\omega_{e},\omega_{p})$ is $2\pi$-periodic with respect to $\omega_{e}$ and $\omega_{p}$, we considered a parameterised family of functions of the form
$\Phi(X,Y,\omega_{e},\omega_{p};\theta)$ $=$ $\sum_{n=0}^{4}\phi(X,Y;\theta_{n}^{(1)})\cos(n\omega_{e})+\sum_{m=1}^{4}\phi(X,Y;\theta_{m}^{(2)})\sin(m\omega_{e})$
$+\sum_{n=0}^{4}\phi(X,Y;\theta_{n}^{(3)})\cos(n\omega_{p})+\sum_{m=1}^{4}\phi(X,Y;\theta_{m}^{(4)})\sin(m\omega_{p})$
where the function $\phi(\cdot,\cdot;\theta):\mathbb{R}^{2}\to\mathbb{R}$ is a three-layer fully connected NN with 80 neurons in each hidden layer. We trained the NN for three rounds of Algorithm 2, using the following values for the parameters $\alpha$ and $\delta$:
$\alpha=[2.5,2,1.5]\quad\text{and}\quad\delta=[0.7,0.5,0.3].$
In each application of Algorithm 1, we run $3000$ iterations of stochastic gradient descent, using $N_{b_{1}}=100$ boundary points for the supervised data, and $N=800$ collocation points to approximate the least-squares finite-difference loss. The running time to train each NN was 120 seconds on a CPU. In Figure 8, each sample trajectory of the game is computed according to (5.18) and (5.19), with a trained NN approximating the value function $u$. In the first example, the Pursuer moves faster than the Evader ($\sigma_{p}>\sigma_{e}$), but the Evader has a higher angular velocity ($1/\rho_{e}>1/\rho_{p}$). The plot at the right represents the evolution of the distance between the players throughout the game. We see oscillatory behaviour due to the turns that the Evader performs for not being intercepted. Near the maxima of the curve, we see that the distance increases abruptly before it starts decreasing. This is due to the small maneuver of the Pursuer (similar to last example in Figure 7) allowing to change the direction faster. In the second example, we see a trajectory in which both players have the same angular velocity, but the pursuer moves faster. Hence, the turning radius of the Evader is smaller ($\sigma_{e}\rho_{e}<\sigma_{p}\rho_{p}$). In this case, the game enters in a loop. Finally, in the third example, we see a trajectory in which the pursuer moves faster, and both players have the same turning radius ($\sigma_{e}\rho_{e}=\sigma_{p}\rho_{p}$). In this case, the Pursuer has a clear advantage over the Evader, and interception occurs in a short time interval.
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=1.2$ Refer to caption Refer to caption
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=1$ Refer to caption Refer to caption
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=0.8$ Refer to caption Refer to caption
Players’ velocities Game trajectory Time versus the dist. between players
Figure 8: Three trajectories of the Pursuit Evasion game with Reeds-Shepp’s cars discussed in Section 5.2.4. We consider different velocity parameters for the players (left) and different initial positions shown by the stars.
6 Proofs of the main results
6.1 Notation and preliminaries
In view of Theorems 1 and 2 and Corollary 1, in this section, we only consider the $d$-dimensional cube $\Omega:=(0,1)^{d}$. For any $N\in\mathbb{N}$, fix $\delta=1/N$, and consider the uniform Cartesian grids of the domains $\overline{\Omega}$ and $\Omega$, given by
$\overline{\Omega}_{\delta}:=\delta\mathbb{Z}^{d}\cap\overline{\Omega}=\left\{x_{\beta}:=\delta\beta\,:\ \beta\in\{0,1,\ldots,N\}^{d}\right\}$ (6.1)
$\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega=\left\{x_{\beta}:=\delta\beta\,:\ \beta\in\{1,\ldots,N-1\}^{d}\right\}$ (6.2)
From now on, we will use the multi-index notation $\beta:=(\beta_{1},\ldots,\beta_{d})\in\{0,1,\ldots,N\}^{d}$. Furthermore, we denote by $\{e_{i}\}_{i=1}^{d}$ the canonical basis of $\mathbb{R}^{d}$, so that $x_{\beta+e_{i}}=x_{\beta}+\delta e_{i}$ and $x_{\beta-e_{i}}=x_{\beta}-\delta e_{i}$. We also introduce the index domains
$\overline{\mathfrak{B}}:=\{0,\ldots,N\}^{d}\quad\text{and}\quad\mathfrak{B}:=\{1,\ldots,N-1\}^{d}.$ (6.3)
Then, we can then write $\overline{\Omega}_{\delta}=\{x_{\beta}\,:\ \beta\in\overline{\mathfrak{B}}\}$ and $\Omega_{\delta}=\{x_{\beta}\,:\ \beta\in\mathfrak{B}\}$.
For a given a function $u\in C(\overline{\Omega})$, let us denote by $U$ the grid function associated to $u$ on the grid $\overline{\Omega}_{\delta}$, i.e.
$U:=\left\{U_{\beta}=u(x_{\beta})\,:\ \beta\in\overline{\mathfrak{B}}\right\}.$
Using the above notation, we can write the functional ^$\widehat{\mathcal{R}}(u)$ as
^$\widehat{\mathcal{R}}(u)$ $=$ $\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}$ (6.4)
$=$ $\delta^{d}\sum_{x\in\Omega_{\delta}}\left[H\left(x,\dfrac{D_{\delta}^{+}u(x)+D_{\delta}^{-}u(x)}{2}\right)-\alpha\sum_{i=1}^{d}\left(\dfrac{D_{\delta}^{+}u(x)-D_{\delta}^{-}u(x)}{2}\right)_{i}\right]^{2}$
$=$ $\delta^{d}\sum_{\beta\in\mathfrak{B}}\left[H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}\right]^{2}=:F(U),$
where, for any $\beta\in\mathfrak{B}$, $D_{\delta}U_{\beta}$ denotes the centred finite-difference approximation of the gradient, i.e.
$D_{\delta}U_{\beta}=\left(\dfrac{U_{\beta+e_{1}}-U_{\beta-e_{1}}}{2\delta},\ldots,\dfrac{U_{\beta+e_{d}}-U_{\beta-e_{d}}}{2\delta}\right)\in\mathbb{R}^{d},$ (6.5)
and, for any $i\in\{1,\ldots,d\},$ $D_{\delta,i}^{2}U_{\beta}$ is given by
$D_{\delta,i}^{2}U_{\beta}=\dfrac{U_{\beta+e_{i}}+U_{\beta-e_{i}}-2U_{\beta}}{2\delta}.$ (6.6)
In view of the assumption (1.3), the functional $U\mapsto F(U)$ defined in (6.4) is differentiable in the space of grid functions on $\overline{\Omega}_{\delta}$. The main step in the proof of Theorem 1 is to verify that if $\nabla F(U)=0$, then
$H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}=0,\qquad\forall\beta\in\mathfrak{B}.$
In the next lemma, we compute the partial derivatives of $F(U)$ with respect to $U_{\beta}$, for all $\beta\in\mathfrak{B}$.
Lemma 1.
Let $H$ satisfy (1.3), $\alpha>0$ and, for any $N\in\mathbb{N}$ and $\delta=1/N$, consider the grid $\overline{\Omega}_{\delta}$ as in (6.1), and the index domains $\overline{\mathfrak{B}}$ and $\mathfrak{B}$ as in (6.3). Let $F(U)$ be defined as in (6.4) for any grid function $U$ on $\overline{\Omega}_{\delta}$. Then we have
$\partial_{U_{\beta}}F(U)=-\delta^{d-1}\left(\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right)+\alpha\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right)\right),\qquad\forall\beta\in\mathfrak{B},$
where
$W_{\beta}:=\begin{cases}H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta},&\forall\beta\in\mathfrak{B},\\ 0&\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B},\end{cases}$ (6.7)
and for all $i\in\{1,\ldots,d\}$,
$V_{\beta}^{(i)}:=\begin{cases}\partial_{p_{i}}H\left(x_{\beta},D_{\delta}U_{\beta}\right),&\forall\beta\in\mathfrak{B}\\ 0&\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}.\end{cases}$ (6.8)
Proof.
Note that, in view of the definition of $F(U)$ in (6.4) and the definition of $W=\{W_{\beta}\}_{\beta\in\overline{\mathfrak{B}}}$ in (6.7), we have
$F(U)=\sum_{\beta\in\overline{\mathfrak{B}}}W_{\beta}^{2}.$
For any $\beta\in\mathfrak{B}$, the term $U_{\beta}$ appears only in the terms associated to the index $\beta$ and the indexes $\beta\pm e_{i}$ for all $i\in\{1,\ldots,d\}$. Moreover, we note that, in view of (6.5), the term $H(x_{\beta},D_{\delta}U_{\beta})$ is independent of $U_{\beta}$, and for any $i$, the terms $H(x_{\beta\pm e_{i}},D_{\delta}U_{\beta\pm e_{i}})$ depend on $U_{\beta}$ only through the $i$-th component of the vector $D_{\delta}U_{\beta\pm e_{i}}$.
Hence, we can write $\partial_{U_{\beta}}F(U)$ as
$\partial_{U_{\beta}}F(U)$ $=$ $-2\delta^{d}\alpha W_{\beta}\sum_{i=1}^{d}\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta}\right)$
$+2\delta^{d}\sum_{i=1}^{d}W_{\beta+e_{i}}\Big{(}\underbrace{\partial_{p_{i}}H\left(x_{\beta+e_{i}},D_{\delta}U_{\beta+e_{i}}\right)}_{V_{\beta+e_{i}}^{(i)}}\partial_{U_{\beta}}\left(D_{\delta}U_{\beta+e_{i}}\right)_{i}-\alpha\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta+e_{i}}\right)\Big{)}$
$+2\delta^{d}\sum_{i=1}^{d}W_{\beta-e_{i}}\Big{(}\underbrace{\partial_{p_{i}}H\left(x_{\beta-e_{i}},D_{\delta}U_{\beta-e_{i}}\right)}_{V_{\beta-e_{i}}^{(i)}}\partial_{U_{\beta}}\left(D_{\delta}U_{\beta-e_{i}}\right)_{i}-\alpha\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta-e_{i}}\right)\Big{)}.$
Using (6.6), we can compute
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+e_{i}}+U_{\beta-e_{i}}-2U_{\beta}}{2\delta}\right)=-\dfrac{1}{\delta}.$
Next, for any $i\in\{1,\ldots,d\}$ such that $\beta+e_{i}\in\mathfrak{B}$ (note that the terms such that $\beta+e_{i}\not\in\mathfrak{B}$ do not matter since $W_{\beta+e_{i}}$ would be zero), we can use (6.5) and (6.6) to compute
$\partial_{U_{\beta}}\left(D_{\delta}U_{\beta+e_{i}}\right)_{i}=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+2e_{i}}-U_{\beta}}{2\delta}\right)=-\dfrac{1}{2\delta},$
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta+e_{i}}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+2e_{i}}+U_{\beta}-2U_{\beta+e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta}.$
Similarly, for any $i\in\{1,\ldots,d\}$ such that $\beta-e_{i}\in\mathfrak{B}$, we can compute
$\partial_{U_{\beta}}\left(D_{\delta}U_{\beta-e_{i}}\right)_{i}=\partial_{U_{\beta}}\left(\dfrac{U_{\beta}-U_{\beta-2e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta},$
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta-e_{i}}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta}+U_{\beta-2e_{i}}-2U_{\beta-e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta}.$
By plugging all these values in (6.1), we obtain
$\partial_{U_{\beta}}F(U)$ $=$ $\delta^{d-1}\sum_{i=1}^{d}2\alpha W_{\beta}+\delta^{d-1}\sum_{i=1}^{d}\left(-V_{\beta+e_{1}}^{(i)}-\alpha\right)W_{\beta+e_{i}}+\delta^{d-1}\sum_{i=1}^{d}\left(V_{\beta-e_{i}}^{(i)}-\alpha\right)W_{\beta-e_{i}},$
and the conclusion follows by regrouping the terms in the above sum. ∎
6.2 Proof of Theorem 1 and Corollary 1
We are now in position to carry out the proof of Theorem 1, which consists of two steps:
  1. i.
    First we prove that if $u\in C(\overline{\Omega})$ is a critical point of $\widehat{R}(\cdot)$, then the associated grid function $U$ on $\overline{\Omega}_{\delta}$ satisfies $\nabla F(U)=0$, where $F(U)$ is given by (6.4).
  2. ii.
    Then, we use Lemma 1 to prove that if $u$ is $L$-Lipschitz, and $\alpha$ and $\delta$ satisfy (4.8), then $\nabla F(U)=0$ implies that $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0$ for all $x\in\Omega_{\delta}$.
Proof.
Step 1: Let $u\in C(\overline{\Omega})$. If $u$ is a critical point of $\widehat{\mathcal{R}}(\cdot)$, then it must hold that
$\lim_{\varepsilon\to 0}\dfrac{\widehat{\mathcal{R}}(u+\varepsilon\phi)-\widehat{\mathcal{R}}(u)}{\varepsilon}=0,\qquad\text{for any function $\phi\in C(\overline{\Omega})$.}$ (6.10)
Now, for any grid function $\Phi$ on $\overline{\Omega}_{\delta}$, let $\phi\in C(\overline{\Omega})$ be a function such that $\phi(x_{\beta})=\Phi_{\beta}$ for all $\beta\in\overline{\mathfrak{B}}$. In other words, consider a continuous function $\phi$ such that the associated grid function is $\Phi$. In view of (6.10), and using the function $F(\cdot)$ as defined in (6.4), we deduce that
$0=\lim_{\varepsilon\to 0}\dfrac{\widehat{\mathcal{R}}(u+\varepsilon\phi)-\widehat{\mathcal{R}}(u)}{\varepsilon}=\lim_{\varepsilon\to 0}\dfrac{F(U+\varepsilon\Phi)-F(U)}{\varepsilon}=\nabla F(U)\cdot\Phi.$
Since this holds for any grid function $\Phi$, we deduce that $\nabla F(U)=0$.
Step 2: In view of Lemma 1, the equation $\nabla F(U)=0$ can be written as
$-\delta^{d-1}(A_{N}(V)+\alpha\Delta_{N})W=0,$
where $W$ and $V=(V^{(1)},\ldots,V^{(d)})$ are the grid functions on $\Omega_{\delta}$ defined in the statement of Lemma 1, and $A_{N}(V)$ and $\Delta_{N}$ are linear operators on the space of grid functions on $\Omega_{\delta}$, defined in Lemmas 2 and 3 respectively. The conclusion of the proof follows after proving that the linear operator $A_{N}(V)+\alpha\Delta_{N}$ is invertible, so the optimality condition $\nabla F(U)=0$ implies that $W=0$, and hence,
$\widehat{H}_{\alpha}(x_{\beta},D_{\delta}^{+}u(x_{\beta}),D_{\delta}^{-}u(x_{\beta}))=H(x_{\beta},D_{\delta}U_{\beta})-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}=W_{\beta}=0,\quad\forall\beta\in\mathfrak{B}.$
Recall that, by the Definition (6.2), $\Omega_{\delta}=\{x_{\beta}\,:\ \beta\in\mathfrak{B}\}$. To prove that the operator $A_{N}(V)+\alpha\Delta_{N}$ is invertible, it is enough to prove that
$\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)=\min_{W}\dfrac{\|\left(A_{N}(V)+\alpha\Delta_{N}\right)W\|_{2}}{\|W\|_{2}}>0.$
Note that $\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)$ is a lower bound for the modulus of the eigenvalues of $A_{N}(V)+\alpha\Delta_{N}$. Using Lemmas 2 and 3 below, we have that
$\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)$$\geq$ $\min_{W}\dfrac{\alpha\|\Delta_{N}W\|_{2}-\|A_{N}(V)W\|_{2}}{\|W\|_{2}}$
$\geq$ $\min_{W}\dfrac{\alpha\|\Delta_{N}W\|_{2}}{\|W\|_{2}}-\max_{W}\dfrac{\|A_{N}(V)W\|_{2}}{\|W\|_{2}}$
$\geq$ $4\alpha d\sin^{2}\left(\dfrac{\pi}{2}\delta\right)-2dC_{H}(L).$
Hence, we deduce that $\underline{\sigma}(A_{N}(V)+\alpha\Delta_{N})>0$ provided $\alpha>\dfrac{C_{H}(L)}{2\sin^{2}\left(\frac{\pi}{2}\delta\right)}$, where $C_{H}(L)$ is defined in the statement of Lemma 2. ∎
Next, we prove the two following lemmas, which are used, at the end of the proof of Theorem 1, to bound the spectrum of the operators $A_{N}(V)$ and $\Delta_{N}$. Although these results might be known to the expert reader, we include them for completeness.
Lemma 2.
Let $H$ satisfy (1.3), consider the domain $\Omega=(0,1)^{d}$ and, for any $N\in\mathbb{N}$ and $\delta=1/N$, let $\overline{\Omega}_{\delta}$ be the grid defined in (6.1). For any function $u\in C(\overline{\Omega})$ with Lipschitz constant $L>0$, let $U$ be the grid function associated to $u$ on the grid $\overline{\Omega}_{\delta}$, and set $V=(V^{(1)},\ldots,V^{(d)})$, where for each $i\in\{1,\ldots,d\}$, $V^{(i)}$ is the grid function on $\overline{\Omega}_{\delta}$ defined by (6.8).
Then, the linear operator defined, for any grid function $W$ on $\overline{\Omega}_{\delta}$, by
$A_{N}(V)W=\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right),$
satisfies
$\overline{\sigma}\left(A_{N}(V)\right):=\max_{W}\dfrac{\|A_{N}(V)W\|_{2}}{\|W\|_{2}}\leq 2dC_{H}(L),$
where $C_{H}(L)=\displaystyle\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|$.
Proof.
Using the definition of $V^{(i)}$ in (6.8) and the fact that $u$ has Lipschitz constant $L$, it holds that
$V_{\beta}^{(i)}\leq C_{H}(L),\quad\forall\beta\in\mathfrak{B}\ \text{and}\ i\in\{1,\ldots,d\}.$
Then, in view of the definition of $A_{N}(V)$, it follows that
$\overline{\sigma}(A_{N}(V))\leq C_{H}(L)\overline{\sigma}(\overline{A}),$ (6.11)
where $\overline{A}$ is the linear operator, on the space of grid functions on $\Omega_{\delta}$, given by
$\overline{A}=\sum_{i=1}^{d}A^{(i)},$
with
$A^{(1)}$ $=$ $B\otimes I\otimes\cdots\otimes I$
$A^{(2)}$ $=$ $I\otimes B\otimes\cdots\otimes I$
$\vdots$
$A^{(d)}$ $=$ $I\otimes I\otimes\cdots\otimes B,$
where $I$ is the $(N-1)\times(N-1)$ identity matrix, and $B\in\mathbb{R}^{(N-1)\times(N-1)}$ is given by
$B=\begin{bmatrix}0&1&\\ 1&0&1\\ &1&0&1\\ &&\ddots&\ddots&\ddots\\ &&&1&0&1\\ &&&&1&0\end{bmatrix}.$
Finally, by Gershgorin Theorem, the largest eigenvalue of $B$ smaller or equal than $2$, so we conclude that
$\overline{\sigma}(\overline{A})\leq\sum_{i=1}^{d}\overline{\sigma}\left(A^{(i)}\right)\leq 2d,$
and the conclusion of the lemma follows from (6.11). ∎
Lemma 3.
Let $N\in\mathbb{N}$, $\delta=1/N$, and consider the grid $\Omega_{\delta}$ as in (6.1) and the index domain $\mathfrak{B}$ as in (6.3). The linear operator $\Delta_{N}$, defined for any grid function $W$ on $\Omega_{\delta}$ as
$\Delta_{N}W=\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right),$
with $W_{\beta\pm e_{i}}=0$ whenever $\beta\pm e_{i}\not\in\mathfrak{B}$, satisfies
$\underline{\sigma}(\Delta_{N}):=\min_{W}\dfrac{\|\Delta_{N}W\|_{2}}{\|W\|_{2}}=4d\sin^{2}\left(\dfrac{\pi}{2}\delta\right).$
Proof.
The operator $\Delta_{N}$ is the Dirichlet Laplacian associated with the $d$-dimensional grid $\Omega_{\delta}$, and is a symmetric negative definite operator. Hence, $\underline{\sigma}(\Delta_{N})$ coincides with the smallest eigenvalue of $-\Delta_{N}$. We can write it as
$-\Delta_{N}=-\sum_{i=1}^{d}\Delta_{N}^{(i)},$
with
$\Delta_{N}^{(1)}$ $=$ $L_{N}\otimes I\otimes\cdots\otimes I$
$\Delta_{N}^{(2)}$ $=$ $I\otimes L_{N}\otimes\cdots\otimes I$
$\vdots$
$\Delta_{N}^{(d)}$ $=$ $I\otimes I\otimes\cdots\otimes L_{N},$
and where
$L_{N}=\begin{bmatrix}-2&1&\\ 1&-2&1\\ &1&-2&1\\ &&\ddots&\ddots&\ddots\\ &&&1&-2&1\\ &&&&1&-2\end{bmatrix}\in\mathbb{R}^{(N-1)\times(N-1)}.$
Any eigenvector of $-\Delta_{N}$ can be written as $\Phi=\bigotimes_{i=1}^{d}\phi^{(i)},$ where $\phi^{(i)}\in\mathbb{R}^{N-1}$ is an eigenvector of $-L_{N}$ with associated eigenvalue $\Lambda=\sum_{i=1}^{d}\lambda^{(i)}$, where each $\lambda^{(i)}$ is the eigenvalue associated to $\phi^{(i)}$. Hence, the smallest eigenvalue of $-\Delta_{N}$ is $d\lambda_{1}$, where $\lambda_{1}$ is the smallest eigenvalue of $-L_{N}$. Since the latter is well-known to be $\lambda_{1}=4\sin^{2}\left(\frac{\pi}{2N}\right)$, we conclude that
$\underline{\sigma}(\Delta_{N})=d\lambda_{1}=4d\sin^{2}\left(\dfrac{\pi}{2N}\right)=4d\sin^{2}\left(\dfrac{\pi}{2}\delta\right).$
Next, we give the proof of Corollary 1. The proof is almost identical to that of Theorem 1, with the only difference being that one has to consider the functional’s partial derivatives with respect to the boundary points.
Proof of Corollary 1.
First, for any function $u\in C(\overline{\Omega})$, let us write the functional $\widehat{\mathcal{J}}(u)$ as
$\widehat{\mathcal{J}}(u)=F(U)+\gamma G(U),$
where $F(U)$ is defined as in (6.4), and $G(U)$ is given by
$G(U):=\delta^{d-1}\sum_{\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}}\left(U_{\beta}-g(x_{\beta})\right)^{2}.$
Here, we use the multi-index notation introduced in (6.3).
Using precisely the same argument as in step 1 in the proof of Theorem 1, we have that if $u$ is a critical point of $\widehat{\mathcal{J}}(u)$, then it must hold that $\nabla F(U)+\gamma\nabla G(U)=0$.
Since $G(U)$ does not depend on $U_{\beta}$ for any $\beta\in\mathfrak{B}$, we have that $u$ being a critical point of $\widehat{\mathcal{J}}_{\alpha,\delta,\gamma}(u)$ implies that
$\partial_{U_{\beta}}F(U)=-\delta^{d-1}\left(\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right)+\alpha\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right)\right)=0,\qquad\forall\beta\in\mathfrak{B},$
where $W_{\beta}$ and $V_{\beta}^{(i)}$ are given as in (6.7) and (6.8). Now, using the step 2 in the proof of Theorem 1, we obtain that, if (4.8) holds, then $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$.
Since $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$, we have that
$\partial_{U_{\beta}}F(U)+\gamma\partial_{U_{\beta}}G(U)=\gamma\partial_{U_{\beta}}G(U)=2\gamma\delta^{d-1}\left(U_{\beta}-g(x_{\beta})\right),\qquad\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}.$
Hence, $\partial_{U_{\beta}}F(U)+\gamma\partial_{U_{\beta}}G(U)=0$ implies that $U_{\beta}=g(x_{\beta})$ for all $\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}$. This, combined with $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$ concludes the proof. ∎
6.3 Proof of Theorem 2
The main idea of the proof is to re-write the functional $\mathcal{R}(\cdot)$ in such a way that we can apply the conclusion of Theorem 1 for the discrete functional $\widehat{\mathcal{R}}(\cdot)$.
Proof.
Let $N\in\mathbb{N}$ and fix $\delta=1/(N-1)$. The interval $[0,1)$ can be partitioned in $N-1$ disjoint intervals as
$[0,1)=\left[0,\dfrac{1}{N-1}\right)\sqcup\left[\dfrac{1}{N-1},\dfrac{2}{N-1}\right)\sqcup\ldots\sqcup\left[\dfrac{N-2}{N-1},1\right)=\bigsqcup_{\beta\in\{1,\ldots,N-1\}}\left(\delta\beta+[-\delta,0)\right).$
Hence, for any dimension $d$, we can write $[0,1)^{d}$ similarly as a disjoint union of $(N-1)^{d}$ disjoint $d$-dimensional cubes as
$[0,1)^{d}=\bigsqcup_{\beta\in\mathfrak{B}}\left(\delta\beta+[-\delta,0)^{d}\right),\quad\text{where $\mathfrak{B}:=\{1,\ldots,N-1\}^{d}$.}$
Now, recalling that $\Omega=(0,1)^{d}$, we can use this partition to write $\mathcal{R}(u)$ defined in (4.10) as
[…]
$=$ $\int_{(-\delta,0)^{d}}\sum_{\beta\in\mathfrak{B}}\left[\widehat{H}_{\alpha}(z-\delta\beta,D_{\delta}^{+}u(z-\delta\beta),D_{\delta}^{-}u(z-\delta\beta))\right]^{2}dz$
[…]
where
^$\widehat{\mathcal{R}}(u;z):=\sum_{x\in\Omega_{\delta}(z)}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2},$
with $\Omega_{\delta}(z):=\{x_{\beta}=z+\delta\beta\,:\ \beta\in\mathfrak{B}\}$ for all $z\in(0,1)^{d}$.
Note that $\Omega_{\delta}(z)$ is simply a shift of the grid $\Omega_{\delta}$ defined in (6.2). Hence, by virtue of Theorem 1, if $u\in C(\overline{\Omega})$ is a critical point of ^$\widehat{\mathcal{R}}(u;z)$, then $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0$ for all $x\in\Omega_{\delta}(z)$.
For any functions $u,v\in C(\overline{\Omega})$, and any $z\in(0,1)^{d}$, let $U(z)$ and $V(z)$ be the grid functions of $u$ and $v$ associated to the grid $\Omega_{\delta}(z)$, i.e.
$U(z):=\{u(x)\,:x\in\Omega_{\delta}(z)\},\quad\text{and}\quad V(z):=\{v(x)\,:x\in\Omega_{\delta}(z)\}.$
We also define, for any $z\in(0,\delta)^{d}$, the functional on the space of grid functions on $\Omega_{\delta}(z)$ given by
$F_{z}(U(z)):=\sum_{x\in\Omega_{\delta}(z)}\left[\widehat{H}_{\alpha}\left(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x)\right)\right]^{2}=\widehat{R}_{\alpha,\delta}(u;z).$
If $u\in C(\overline{\Omega})$ is a critical point of $\mathcal{R}(\cdot)$, then for any function $v\in C(\overline{\Omega})$ it must hold that
$\lim_{\varepsilon\to 0}\dfrac{\mathcal{R}(u+\varepsilon v)-\mathcal{R}(u)}{\varepsilon}=0,$
or equivalently, we can use (6.12) to write
[…]
Since this holds for any function $v\in C(\overline{\Omega})$, we deduce that $\nabla F_{z}\left(U(z)\right)$ is identically 00 for all $z\in(-\delta,0)^{d}$. Whence, $u$ is a critical point of ^$\widehat{\mathcal{R}}(u;z)$ for all $z$, and the conclusion follows. ∎
7 Summary and perspectives
In this paper, we propose to solve the boundary value problems for a class of Hamilton-Jacobi equations using Deep Learning. Artificial neural networks are trained by typical gradient descent-based algorithms, optimizing for the least square problem defined by the residual of a consistent and monotone finite difference scheme for the equation. We analyze the critical points of the least square problem in the space of continuous functions and reveal conditions that guarantee the critical points being the unique viscosity solution of the problem. These conditions may involve the parameters of the chosen finite difference scheme or those coming from additional “supervised” data points. We demonstrated numerically that the proposed approach effectively solves Hamilton-Jacobi equations in higher dimensions under a favourable setup.
An important point is that artificial neural networks are used instead of grid functions. Thus, one may take advantage of neural networks’ approximation power for the smooth parts of a solution and the scaling property of Monte-Carlo sampling involved in the training procedure. Thus, under favourable conditions, it is reasonable to expect that the proposed approach would suffer less the so-called “curse of dimensionality” than the classical algorithms.
But during the training iterations, one must ensure that the (high dimensional) computational domain is sufficiently “covered” — enough to effectively follow the characteristics from the boundaries to everywhere in the domain and “monitor” how the monotone scheme resolves the collision of the characteristics. This is crucial for computing accurate approximation of the viscosity solutions.
In PINNs, it is common knowledge that the choice of collocation points can determine the success of the methods. This is also true for the proposed approach, given a limited training budget. In higher dimensions, due to the “concentration of measure” errors, the errors “deeper” in the interior of $\Omega$ are significantly “discounted.” However, reducing the errors in the interior may be essential to constructing approximate solutions with the proper global structure (such as convexity). Therefore, uniform sampling (relative to the Lebesque measure) in training or assigning additional “supervised” data may be sub-optimal. We plan to develop adaptive algorithms for the proposed approach in the near future.
We point out that it is possible to use high-resolution schemes. However, suppose a gradient descent-based optimization scheme that utilizes automatic differentiation is used. In that case, the chosen numerical scheme must be a differentiable function to the variables defined by the scheme’s stencil. This formerly precludes the use of higher order essentially non-oscillatory (ENO) schemes [12]. However, the weighted essentially oscillatory (WENO) schemes [10] can be considered.
Finally, we remark that the idea of using a convergent numerical method to form least square principles is not limited to only finite difference methods. It is possible to develop such an algorithm using the finite element methodology. In that case, we envision that during training, proper mesh refinement strategies may be involved to enhance error control.
Acknoledgement
Tsai is partially supported by NSF grant DMS-2110895 and by Army Research Office Grant W911NF2320240.