### v4

# Finite-difference least square methods for solving Hamilton-Jacobi equations using neural networks

Abstract
We present a simple algorithm to approximate the viscosity solution of Hamilton-Jacobi (HJ) equations using an artificial deep neural network. The algorithm uses a stochastic gradient descent-based method to minimize the least square principle defined by a monotone, consistent numerical scheme. We analyze the least square principle’s critical points and derive conditions that guarantee that any critical point approximates the sought viscosity solution. Using a deep artificial neural network lifts the restriction of conventional finite difference methods that rely on computing functions on a fixed grid. This feature makes it possible to solve HJ equations posed in higher dimensions where conventional methods are infeasible. We demonstrate the efficacy of our algorithm through numerical studies on various canonical HJ equations across different dimensions, showcasing its potential and versatility.
Keywords Hamilton-Jacobi equations, deep learning, finite-difference methods, least squares principle, optimal control, and differential games.
AMS subject classifications (MSC2020) 49L25, 35A15, 68T07, 49K20, 65N06
1 Introduction
In recent years, significant research has focused on computing solutions for partial differential equations in high dimensions. These endeavours are driven by advancements in deep learning and new numerical optimisation algorithms. In particular, there is great interest in numerical algorithms for Hamilton-Jacobi equations arising in optimal control problems and differential games.
We consider the boundary value problem associated with a Hamilton-Jacobi equation of the form
$\begin{cases}H(x,\nabla u(x))=0&x\in\Omega\\ u(x)=g(x)&x\in\partial\Omega,\end{cases}$ (1.1)
where $\Omega\subset\mathbb{R}^{d}$, with $d\geq 1$, is an open bounded domain with Lipschitz boundary, and
$H:\Omega\times\mathbb{R}^{d}\to\mathbb{R}\quad\text{and}\quad g:\partial\Omega\to\mathbb{R}$
are the given functions, which we shall refer to as the Hamiltonian and the boundary condition, respectively. Throughout the paper, we shall assume that
$x\mapsto g(x)\ \text{is Lipschitz on $\partial\Omega$}$ (1.2)
and
$(x,p)\mapsto H(x,p)\ \text{is locally Lipschitz in $\Omega\times\mathbb{R}^{d}$ and differentiable with respect to $p$.}$ (1.3)
Certain types of Hamilton-Jacobi equations describe the optimality condition for an optimal control problem. By applying Bellman’s dynamic programming principle, one can show that the value function of the associated optimal control problem formally satisfies the so-called Hamilton-Jacobi-Bellman equation (see [2]). However, it is well-known that even if the boundary condition $g$ is very smooth, the boundary value problem (1.1) is not guaranteed a classical solution. This happens when the characteristics of (1.1), representing the optimal trajectories of the optimal control problem, collide with each other. On the other hand, continuous functions that satisfy the PDE for almost every $x\in\Omega$ along with the boundary condition are not necessarily unique. In [8], Crandall and Lions introduced the notion of viscosity solution to single out a unique weak solution for the problem (1.1). Furthermore, in many situations, one can prove that the viscosity solution coincides with the value function, and one may use it to synthesize the optimal feedback for the control problem.
A characterisation of the viscosity solution can be obtained through the so-called vanishing viscosity method, which consists of approximating the viscosity solution by the unique classical solution to the problem
$\begin{cases}H(x,\nabla u^{\varepsilon}(x))-\varepsilon\Delta u^{\varepsilon}(x)=0,&x\in\Omega,\\ u^{\varepsilon}(x)=g(x),&x\in\partial\Omega,\end{cases}$ (1.4)
for $\varepsilon>0$. The viscosity solution to (1.1) is obtained as the point-wise limit of $u^{\varepsilon}(x)$ as $\varepsilon\to 0^{+}$. A different characterisation of the viscosity solution exists, using smooth test functions based on the comparison principle (see [7]). Depending on the structure of the Hamiltonian and the boundary data, the viscosity solution may be discontinuous (see [2, 16]). This is typically true for Hamilton–Jacobi equations arising in differential game theory or when the boundary condition is prescribed along a characteristic curve. However, in this paper, we make the following assumption:
Assumption 1.
The boundary value problem (1.1) has a Lipschitz continuous viscosity solution.
The design and implementation of numerical algorithms to approximate the viscosity solution of (1.1) is a classical problem widely considered throughout the years. In two or three dimensions, Hamilton-Jacobi equations are typically solved by special finite-difference methods operating on uniform Cartesian grids. The Hamiltonian $H$ is approximated by a so-called numerical Hamiltonian $\widehat{H}$ defined on the grid and the grid function. If $\widehat{H}$ is consistent with $H$ and monotone (see Definition 1), the solution to the system of discrete equations defined by $\widehat{H}$ on the grid (properly extended to the domain $\overline{\Omega}$) converges uniformly to the viscosity solution as the grid spacing tends to 00 (see e.g. [9, 3]). Similarly to the development of numerical schemes for hyperbolic conservation laws, higher-order approximations to the partial derivatives are proposed to be used with a monotone and consistent numerical Hamiltonian. In parallel, fast algorithms such as the fast marching methods [34, 27, 28, 15] and fast sweeping methods [38, 33] are developed for a certain class of Hamilton-Jacobi equations to solve the resulting nonlinear system with computational complexities that are linear in $N_{d}$, the total number of unknowns (for the fast marching methods, the formal complexity is $N_{d}\log N_{d}$).
The main drawback of finite-difference methods implemented in the classical fashion is that they involve solving a system of nonlinear equations with as many unknowns as the number of grid points. This means that the complexity of these methods scales at least exponentially with the dimension of the domain. Therefore, the classical method of formulating and solving for finite difference approximations on grid functions cannot be carried out in the high-dimensional setup. However, many Hamilton-Jacobi-Bellman equations for optimal control applications are posed in high dimensions. Recently, there has been substantial development in algorithms for solving high dimensional problems utilising ideas from the Hopf-Lax formula and its generalisation; see, e.g. [11],[6]. Other algorithms [12, 18, 13] use tensor decomposition to address special classes of high-dimensional Hamilton-Jacobi equations arising in optimal control theory.
In recent years, Deep Learning approaches have proven very effective at solving some non-linear problems in high dimensions. These methods consist of approximating a target function (in our case, the viscosity solution) by an artificial neural network, whose parameters are optimised by minimising a loss function (typically) through a gradient-based method. This can be formulated as a variational problem of the form
${u}^{*}\in\arg\min_{u\in\mathcal{F}}\mathcal{J}(u),$ (1.5)
where
:=$\mathcal{F}:=\{u(\cdot,\theta)\,:\ \theta\in\mathbb{R}^{P}\}\subset C(\overline{\Omega})$
is a parameterized class of functions given by a Neural Network architecture, and $\mathcal{J}:C(\overline{\Omega})\to\mathbb{R}^{+}$ is the loss functional. Of course, the success of this method heavily relies on the choice of $\mathcal{F}$, $\mathcal{J}(\cdot)$ and the optimisation algorithm.
Ideally, $\mathcal{J}(u)$ should yield smaller values for functions $u$ close to the target function. A common practice in Deep Learning is to seek a minimizer of $\mathcal{J}(\cdot)$ in $\mathcal{F}$ using information derived from $\nabla_{\theta}\mathcal{J}(\Phi(\cdot,\theta)).$ Let us consider the following continuous version of the gradient descent method to find a minimizer of $\mathcal{J}$ in $\mathcal{F}$:
$\frac{d}{dt}\theta_{t}=-\nabla_{\theta}\mathcal{J}(u(\cdot,\theta_{t}))=-\mathcal{J}^{\prime}(u(\cdot,\theta_{t}))\,\nabla_{\theta}u(\cdot,\theta_{t}),$ (1.6)
where, for any continuous function $u\in C(\overline{\Omega})$, $\mathcal{J}^{\prime}(u)\in C(\overline{\Omega})^{\ast}$ denotes the Fréchet derivative of $\mathcal{J}(\cdot)$ at $u$, and $\nabla_{\theta}u(\cdot,\theta)$ denotes the vector–valued function with the partial derivatives of $u(x,\theta)$ with respect to the parameter $\theta$. We see that a parameter $\theta\in\mathbb{R}^{P}$ is a stationary point of (1.6) if and only if all the elements of the vector $\nabla_{\theta}u(\cdot,\theta)\in[C(\overline{\Omega})]^{P}$ lie in the kernel of the linear functional $\mathcal{J}^{\prime}(u(\cdot,\theta))$. The precise form of the vector-valued function $\nabla_{\theta}u(\cdot,\theta)$ depends on the architecture of the Neural Network and will not be discussed in this paper. On the other hand, any parameter $\theta\in\mathbb{R}^{P}$ such that the function $u=u(\cdot,\theta)$ is a solution to the Euler-Lagrange equation $\mathcal{J}^{\prime}(u)=0$ is a critical point of $\theta\mapsto\mathcal{J}(u(\cdot,\theta))$, and therefore, a stationary point for the gradient descent method.
In this paper, we investigate the choice of the functional $\mathcal{J}(\cdot)$. Namely, we are interested in constructing a functional with the two following properties:
  1. i.
    Any global minimiser of $\mathcal{J}(u)$ approximates the viscosity solution of (1.1).
  2. ii.
    Any critical point of $\mathcal{J}(u)$ is a global minimizer.
Note that the second property above is crucial in Deep Learning since the optimisation algorithm used to minimise $\mathcal{J}(u)$ is typically a variant of gradient descent, similar to (1.6). In particular, we shall look at functionals with the structure
$\mathcal{J}(u)=\mathcal{L}(u;g)+\mathcal{R}(u),$ (1.7)
where $\mathcal{L}(u;g)$ and $\mathcal{R}(u)$ are, respectively, fitting terms related to the boundary data $g$ and the PDE in (1.1), with the following specific form:
[…]
where $\widehat{H}$ is a numerical Hamiltonian, and $D^{+}_{\delta}u(x)$ and $D^{-}_{\delta}u(x)$ are the forward and backward finite-difference approximations of the gradient $\nabla u(x)$. The main contribution of the current paper is to prove that if we construct $\mathcal{R}(u)$ based on a suitable Lax-Friedrichs numerical Hamiltonian, then one can ensure that the two properties stated above hold.
We point out that other variational approaches exist for solving Hamilton-Jacobi equations. In [22, 23], the viscosity solution of a Hamilton-Jacobi equation is identified with the saddle-point of a min-max problem, whose solution is addressed through a primal-dual hybrid gradient algorithm. In contrast, in developing our proposed formulation, we aim to benefit from the optimisation techniques commonly used in Deep Learning.
In practice, the optimisation problem (1.5) is addressed by considering a discrete version of the gradient descent method (1.6), applied to a discrete approximation of the functional $\mathcal{J}(u)$ in (1.7). In this paper, we propose an algorithm that leverages the stochastic gradient descent algorithm to minimize $\mathcal{J}$ in the form
$\theta_{n+1}:=\theta_{n}-\eta_{n}\nabla_{\theta}J_{n}(u(\cdot,\theta_{k})),$ (1.8)
where $J_{n}$ is a Monte-Carlo approximation of the integral terms in $\mathcal{J}$, involving a batch of randomly positioned “collocation points” $B_{n}\subset\overline{\Omega}$. We refer to the strategy of changing $B_{n}$ at every iteration as resampling of the domain. This is a crucial step to the success of the proposed algorithm.
In the next section, we illustrate our findings with a simple example. In particular, we consider the Eikonal equation in a one-dimensional interval. First, we compute the optimality condition for a functional based on the PDE residual. Then, we carry out the same computation for a functional based on the residual of a finite-difference approximation of the Hamiltonian. In Section 3, we discuss existing approaches based on Deep Learning for the numerical approximation of PDEs, such as the well-known framework known as Physics informed Neural Networks, or PINNs. In Section 4, we study the critical points of loss functionals based on numerical Hamiltonians of Lax-Friedrichs type and present our main theoretical contributions. Sections 5 and 6 are devoted to the numerical aspects of our approach. In Section 5, we describe our algorithm to approximate the viscosity solution employing a Neural Network, and then, we test our method in several examples of Hamilton-Jacobi equations. In Section 6 we analyse, through numerical experiments, various aspects of our algorithm, such as data efficiency and accuracy. We also show how the resampling strategy dramatically outperforms other approximation strategies, in which the collocation points are reused throughout the gradient descent iterations. In Section 7, we present the proofs of the theoretical results presented in Section 4. Finally, we conclude the paper in Section 8, summarising our contributions and some perspectives for future research.
The code to reproduce the numerical experiments from Sections 5 and 6 is publicly available in the following link:
2 Example
Consider the model viscous Eikonal equation in the unit interval, with Dirichlet boundary conditions:
$\begin{cases}(\partial_{x}u)^{2}=1+\varepsilon\partial_{xx}u&x\in(0,1),\\ u(0)=u_{0},\quad u(1)=u_{1}.\end{cases}$ (2.1)
It is well-known that, for any $\varepsilon\neq 0$ and $u_{0},u_{1}\in\mathbb{R}$, the problem (2.1) has a unique $C^{\infty}$-solution.
Let us consider the minimisation problem
[…]
among functions in $C^{\infty}(0,1)$ satisfying the boundary conditions $u(0)=u_{0}$ and $u(1)=u_{1}$.
For any $u\in C^{\infty}(0,1)$, let us define the PDE-residual function $w(x)=(\partial_{x}u(x))^{2}-\varepsilon\partial_{xx}u(x)-1$. For any test function $v\in C_{0}^{2}(0,1)$, the Gâteaux derivative of $\mathcal{R}_{\varepsilon}(\cdot)$ at $u$ in the direction $v$ can be written as
[…]
Now, suppose that $u^{*}\in C^{\infty}(0,1)$ is a critical point of $\mathcal{R}_{\varepsilon}(\cdot)$. Then, $w$ is a weak solution of
$\begin{cases}\partial_{x}(w\partial_{x}u^{*})+{\dfrac{\varepsilon}{2}}\partial_{xx}w=0&\text{in}\ (0,1),\\ w(0)=w(1)=0.\end{cases}$
It is clear that, for any $\varepsilon\neq 0$, the unique weak solution to this boundary value problem is $w=0$. Therefore any critical point $u^{*}\in C^{\infty}(0,1)$ of $\mathcal{R}_{\varepsilon}(\cdot)$, satisfying the boundary conditions is a classical solution of
$(\partial_{x}u)^{2}-\varepsilon\partial_{xx}u-1=0.$
Since this PDE, coupled with boundary conditions $u(0)=u_{0}$ and $u(1)=u_{1}$, has a unique classical solution, we deduce that this solution is the unique critical point of $\mathcal{R}_{\varepsilon}(\cdot)$ in $C^{\infty}(0,1)$.
The sign of $\varepsilon$ determines the convexity of the solution. In this case, the commonly sought-after viscosity solution corresponds to the solution’s limit as $\varepsilon\rightarrow 0^{+},$ and can be proven to be concave. If we consider $\varepsilon=0$, the boundary value problem (2.1) has no classical solution, so there are no critical points of $\mathcal{R}_{0}(\cdot)$ in $C^{\infty}(0,1)$. This implies, in particular, that the functional $\mathcal{R}_{0}(\cdot)$ is not minimised by any smooth function. However, one can construct minimisers of $\mathcal{R}_{0}(u)$, which are continuous Lipschitz. However, these are not unique, as illustrated in Figure 1.
Figure 1: Four weak solutions of the Eikonal equation $(\partial_{x}u)^{2}-1=0$ in $\Omega:=(0,1)$ with boundary condition $u(0)=u(1)=0$, correspodning to four global minimizers to the functional (3.2).
Despite having proved the uniqueness of a critical point in $C^{\infty}(0,1)$ for any $\varepsilon>0$, it is worth mentioning that there exist other continuous global minimizers of $\mathcal{R}_{\varepsilon}(\cdot)$ which satisfy the boundary condition and are not $C^{\infty}$. Consider, for instance, the boundary value problems
$\begin{cases}(\partial_{x}\phi_{1})^{2}=1+\varepsilon\partial_{xx}\phi_{1}\qquad x\in(0,\,1/2),\\ \phi_{1}(0)=u_{0},\quad\phi_{1}(1/2)=u_{1/2}.\end{cases}\qquad\text{and}\qquad\begin{cases}(\partial_{x}\phi_{2})^{2}=1+\varepsilon\partial_{xx}\phi_{2}\qquad x\in(1/2,\,1),\\ \phi_{2}(1/2)=u_{1/2},\quad\phi_{2}(1)=u_{1},\end{cases}$ (2.3)
[…]
is continuous and minimises the functional $\mathcal{R}_{\varepsilon}(\cdot)$ since it satisfies the boundary condition and the PDE for a.e. $x\in(0,1)$. This proves that the minimiser of $\mathcal{R}_{\varepsilon}(\cdot)$ is not unique in the space of continuous functions. Although we could constrain the optimisation to a class of $C^{\infty}$ functions, this would not solve the problem. Indeed, there exist minimising sequences of $C^{\infty}$ functions converging, pointwise in $(0,1)$, to global minimisers of $\mathcal{R}_{\varepsilon}(\cdot)$, in $C(0,1)$, which are not the classical solution to (2.1).
One could think of introducing regularization terms to convexify the functional and single out the correct viscosity solution. This idea was exploited in [19]. There, the convexity of the functional is achieved by considering the viscous Hamilton-Jacobi equation with a Carleman weight function and a Sobolev norm, which allows for a Carleman-type inequality. We do not pursue in that direction.
Instead of considering functionals based on the PDE residual, we consider minimising the $L^{2}$–residual of a finite-difference numerical scheme. The needed regularisation will come implicitly from the numerical scheme. For any $N\in\mathbb{N}$, set $\delta:=1/N$, the uniform grid $\overline{\Omega}_{\delta}:=\delta\mathbb{Z}\cap[0,1]$, and consider the following discrete version of the PDE-residual functional $\mathcal{R}_{\varepsilon}(\cdot)$ defined in (2.2), with $\varepsilon=0$:
$\min_{u}\widehat{\mathcal{R}}(u):=F(U)=\sum_{i=1}^{N-1}\delta\left[\widehat{H}_{\alpha}\left(\frac{u_{i+1}-u_{i}}{\delta},\frac{u_{i}-u_{i-1}}{\delta}\right)\right]^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to}\ u_{0}=u_{N}=0,$ (2.4)
where, for any $u\in C([0,1])$, the vector $U:=(u_{0},\ldots,u_{N})\in\mathbb{R}^{N+1}$ is the grid function on $\overline{\Omega}_{\delta}$ associated to $u$, and
$\widehat{H}_{\alpha}(p^{+},p^{-})=\left(\dfrac{p^{+}+p^{-}}{2}\right)^{2}-\alpha\dfrac{p^{+}-p^{-}}{2}-1$
for some $\alpha>0$, is the Lax-Friedrichs numerical Hamiltonian associated to $H(p)=p^{2}-1$.
Note that the functional $F(\cdot)$ is differentiable in $\mathbb{R}^{N+1}$, and then, the grid function $U\in\mathbb{R}^{N+1}$ associated to any local minimizer $u\in C([0,1])$ must satisfy the first-order optimality condition $\nabla F(U)=0.$
Denoting $p^{\pm}_{i}=\pm(u_{i\pm 1}-u_{i})/\delta$, we can write the partial derivative of $F(U)$ with respect to $u_{i}$ as
$\partial_{u_{i}}F(U)$ $=$ $\widehat{H}_{\alpha}(p_{i-1}^{+},p_{i-1}^{-})\left(p_{i-1}^{+}+p_{i-1}^{-}-\alpha\right)-\widehat{H}_{\alpha}(p_{i+1}^{+},p_{i+1}^{-})\left(p_{i+1}^{+}+p_{i+1}^{-}+\alpha\right)+2\alpha\widehat{H}_{\alpha}(p_{i}^{+},p_{i}^{-}),$
for any $i=2,\ldots,N-2$.
Now, let us denote
$w_{i}=\widehat{H}_{\alpha}(p_{i}^{+},p_{i}^{-})\quad\text{and}\quad v_{i}=p_{i}^{+}+p_{i}^{-},\qquad\text{for}\ i=1,\ldots,N-1.$
Using this notation, we can write the optimality condition for $u_{i}$ with $i=2,\ldots,N-2$ as
$-(\alpha-v_{i-1})w_{i-1}-(\alpha+v_{i+1})w_{i+1}+2\alpha w_{i}=0,\qquad i=2,\ldots,N-2.$
Further, since
$\partial_{u_{1}}F(U)$ $=$ $2\alpha\widehat{H}_{\alpha}(p_{1}^{+},p_{1}^{-})-\widehat{H}_{\alpha}(p_{2}^{+},p_{2}^{-})(p_{2}^{+}+p_{2}^{-}+\alpha)$
$=$ $-(\alpha+v_{2})w_{2}+2\alpha w_{1},$
$\partial_{u_{N-1}}F(U)$ $=$ $\widehat{H}_{\alpha}(p_{N-2}^{+},p_{N-2}^{-})(p_{N-2}^{+}+p_{N-2}^{-}-\alpha)+2\alpha\widehat{H}_{\alpha}(p_{N-1}^{+},p_{N-1}^{-})$
$=$ $-(\alpha-v_{N-2})w_{N-2}+2\alpha w_{N-1},$
we obtain the first-order optimality condition for the minimisation problem (2.4), which reads as
$\begin{cases}-v_{2}w_{2}-\alpha(w_{2}-2w_{1})=0\\ v_{i-1}w_{i-1}-v_{i+1}w_{i+1}-\alpha(w_{i+1}+w_{i-1}-2w_{i})=0&\text{for}\ i=2,\ldots,N-2\\ v_{N-2}w_{N-2}-\alpha(w_{N-2}-2w_{N-1})=0.\end{cases}$ (2.5)
Denoting $W=(w_{1},\ldots,w_{N-1})$ and $V=(v_{1},\ldots,v_{N-1})$, this system of equations can be written as
$-(A_{N}(V)+\alpha\Delta_{N})W=0,$ (2.6)
where $A_{N}(V)\in\mathbb{R}^{(N-1)\times(N-1)}$ is a matrix depending on $V$ and $\Delta_{N}\in\mathbb{R}^{(N-1)\times(N-1)}$ is the Toeplitz matrix associated to the discrete Laplace operator on a uniform grid with $N-1$ points. It turns out that, for $\alpha$ and $\delta=1/N$ sufficiently large [^1] One can ensure that $A_{N}(V)+\alpha\Delta_{N}$ is invertible by choosing $N$ sufficiently small in $\delta=1/N$. This choice depends on the maximum value in the vector $V=(v_{1},\ldots,v_{N-1})$, or equivalently, on the Lipschitz constant of the grid function $U=(u_{0},\ldots,u_{N})$. See section 7 for the rigorous arguments., the matrix $A_{N}(V)+\alpha\Delta_{N}$ is invertible, which implies that the optimality system (2.6) has only the trivial solution $W=0$. This implies that, if $u$ is a critical point of $\widehat{\mathcal{R}}(\cdot)$, then the associated grid function $U=(u_{0},\ldots,u_{N})$ satisfies
$w_{i}=\left(\dfrac{u_{i+1}-u_{i-1}}{2\delta}\right)^{2}-\alpha\dfrac{u_{i+1}+u_{i-1}-2u_{i}}{2\delta}-1=0,\qquad\text{for all $i=1,\ldots,N-1$.}$ (2.7)
In other words, $U$ satisfies the finite-difference equation associated with the Lax-Friedrichs numerical Hamiltonian $\widehat{H}_{\alpha}(p^{+},p^{-})$. See Section 4 for the precise estimates on $\alpha$ and $\delta$, and Section 7 for the precise computations in the multi-dimensional setup. For $\alpha$ sufficiently large, one can prove that the numerical Hamiltonian $\widehat{H}_{\alpha}(p^{+},p^{-})$ is consistent with $H(p)=p^{2}$ and monotone, as per Definition 1. Moreover, it is known (see [9, 3]) that finite-difference solutions for consistent and monotone numerical schemes converge to the viscosity solution as $\delta\to 0^{+}$. In Section 5, we propose a strategy to train a Neural Network, first with a large value of $\delta$, ensuring that any critical point solves the equation 2.7, and then progressively reducing $\delta$ to ensure convergence to the viscosity solution.
Of course, in the multi-dimensional case, a functional such as ^$\widehat{\mathcal{R}}(u)$ in (2.4) would involve a summation over a multidimensional grid. Computing the full gradient of ^$\widehat{\mathcal{R}}(u)$ would be computationally infeasible. In practice, one can use a version of stochastic gradient descent, in which at every gradient step, the gradient of ^$\widehat{\mathcal{R}}(u)$ is approximated by the gradient of a randomized selection of the terms in the sum.
3 Deep Learning and Partial Differential Equations
Due to its tremendous success in fields such as computer graphics and computer vision, Deep Learning has recently become a popular candidate for tackling various problems where classical methods face limitations. These include the numerical approximation of PDEs in multiple settings. Under the Deep Learning framework, one typically minimizes a functional $\mathcal{J}:C(\Omega)\to\mathbb{R}^{+}$,
${u}^{\ast}\in\arg\min_{u\in\mathcal{F}}\mathcal{J}(u),$ (3.1)
where $\mathcal{F}\subset C(\overline{\Omega})$ is a class of functions defined by a neural network architecture. If $\mathcal{J}(u)\equiv J(u;D)$ is based on some notion of misfit between the minimiser $u$ and a given data set $D$, the setup is often called “supervised learning” in the machine learning community. If $\mathcal{J}$ is defined without labelled data, the setup is often called “unsupervised”.
One of the reasons for the success of Deep Learning techniques is that the commonly used (deep) neural networks can be constructed to approximate continuous functions in high dimensions conveniently [4, 37, 30] and very efficiently for certain classes of smooth functions [1, 31, 32, 36, 35].
Another reason for the success of Deep Learning techniques is the development of optimisation methods, which are used to address the problem (3.1). These are generally gradient-based iterative methods, such as stochastic gradient descent or variants. Hence, choosing the functional $\mathcal{J}(u)$ becomes paramount. For finding the viscosity solution to the boundary value problem (1.1), we consider loss functionals of the form
$\mathcal{J}(u)=\mathcal{L}(u;D)+\mathcal{R}(u).$
In [25], the popular Deep Learning framework known as Physics Informed Neural Networks (PINNs) was introduced. In a PINNs approach, $\mathcal{R}$ can be interpreted as a Monte-Carlo approximation of the squared $L^{2}$–norm of the PDE residual, and $\mathcal{L}$ may involve both boundary data and possibly additional “collocation” points (pointwise values of the solution in $\Omega$). When the data fitting term $\mathcal{L}$ is present, $\mathcal{R}$ is regarded as a regularisation term to some practitioners.
Following the PINNs strategy, a least-square principle for the Hamilton-Jacobi equation would involve
[…]
Assuming that we do not have, in principle, access to the viscosity solution in the interior of the domain, the functional $\mathcal{L}$ can only involve the boundary data:
[…]
In practice, both terms $\mathcal{R}(u)$ and $\mathcal{L}(u;g)$ are approximated through Monte Carlo integration.
We see that any Lipschitz continuous function $u$ satisfying $H(x,\nabla u(x))=0$ for almost every $x\in\Omega$, together with the boundary condition $u(x)=g(x)$ on $\partial\Omega$, is a global minimizer of $\mathcal{J}(u)$. It is well-known that solutions of this type are not unique. This is true even for the simple Eikonal equation on a one-dimensional interval shown in Example 1. This is, of course, very inconvenient since one cannot guarantee by any means that minimising the PINNs functional $\mathcal{J}(u)$ with (3.2) and (3.3) would produce an approximation of the viscosity solution. See Figure 1 for an illustration of four global minimizers of the PINNs functional associated with the one-dimensional Eikonal equation in $(0,1)$ with zero-boundary condition. To overcome the issue of non-uniqueness of minimisers, and promote convergence to the viscosity solution, we replace the $L^{2}$-norm of the PDE-residual with the least-square principle associated with a finite-difference numerical scheme. We prove that the numerical diffusion of the finite-difference scheme enhances regularity on the minimisers. Moreover, by choosing a suitable consistent and monotone numerical scheme, we can prove that the unique minimiser approximates the viscosity solution.
There are also supervised Deep Learning approaches to address Hamilton-Jacobi equations. Different Deep Learning approaches for learning, through data, a weak solution of Hamilton-Jacobi equations are proposed in [10] and in [5]. In [10], data is generated by numerically evolving the Hamiltonian system (the characteristic equations of the Hamilton-Jacobi equation) in the phase space, and a least square principle is used to assign the solution gradient. However, it is unclear if the minimiser is the viscosity solution. In [5], the issue of non-uniqueness of minimisers of the PDE-residual arising in Hamilton-Jacobi equations is alleviated by pre-training the NN with supervised data. In contrast, our approach requires no supervised data for convergence to the viscosity solution.
3.1 Why finite-difference approximations?
Larger domain of dependence/influence.
Finite difference discretization leads to a larger domain of dependence/influence per collocation point compared to the pointwise evaluation of the PDE residual. This feature can lead to a stronger coupling of the “pointwise losses” and more efficient capturing of the problem’s underlying causality. In a finite-difference approximation of the differential operator, each evaluation involves points in a stencil that covers a region of positive measure, related to the discretization step. Instead, a pointwise evaluation of the differential operator using automatic differentiation only involves a single point, which has measure zero. In other words, a residual-based Deep Learning model using finite differences has better data efficiency compared to a similar approach using automatic differentiation. This property is essential as typical solutions of the HJ equation are not classical, and the singular sets of the solutions are lower dimensional. The probability of a finite set of collocation points “sampling” the singular set is proportional to the size of the domain of dependence of the residual term to be minimized. Therefore, using a pointwise evaluation of the PDE residual, it is impossible to distinguish between solutions that satisfy the PDE in an almost everywhere sense. See Section 5.1 for further details.
A similar reasoning can be done in terms of the complexity of the NN. Consider a real-valued function, $u$, defined by a ReLU-activated feedforward neural network. This function is a continuous piecewise linear function. The collocation of $u$ with the given partial differential equation directly regulates only one of its “linear” pieces. Each linear piece is defined by a subset of the network’s parameters. It is possible that the collocated linear pieces are not connected due to the number of collocation points being far fewer than the number of linear pieces in $u$. In contrast, collocating a finite difference equation regulates the relationship of possibly multiple linear pieces, influencing a larger subset of the parameters in $u$. This means that a finite difference-based residual functional reduces the chance of overfitting.
In Section 6.1, our numerical experiments suggest that using a finer discretization of the PDE in the loss functional requires a larger number of collocation points during the optimisation algorithm. Evaluating the PDE residual at the collocation points would be equivalent to letting the discretization step go to zero. In this case, the number of collocation points needed to compute a reasonable approximation of the solution would be significantly larger.
Simpler loss gradient evaluations.
While automatic differentiation can be a very convenient and efficient tool for evaluating first-order derivatives, the evaluation of second-order derivatives can be costly with a large memory footprint. See e.g. [29] and [20], Tab. V.
Convergence to the viscosity solutions.
By using a convergent monotone numerical scheme, e.g. [9], our proposed algorithm is capable of computing the viscosity solution of the given boundary value problem, without explicitly determining the convexity of the solution or where the characteristics of problems should collide.
4 Finite difference residual functionals
Let us consider a numerical Hamiltonian
$\widehat{H}(x,p^{+},p^{-}):\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}.$
For a fixed $\delta>0$, we use the notation
$D^{+}_{\delta}u(x):=\left(\dfrac{u(x+\delta e_{1})-u(x)}{\delta},\ldots,\dfrac{u(x+\delta e_{d})-u(x)}{\delta}\right)\in\mathbb{R}^{d}$
and
$D^{-}_{\delta}u(x):=\left(\dfrac{u(x)-u(x-\delta e_{1})}{\delta},\ldots,\dfrac{u(x)-u(x-\delta e_{d})}{\delta}\right)\in\mathbb{R}^{d},$
respectively for the upwind and downwind finite-difference approximation of the gradient $\nabla u(x)$. Here, the vectors $\{e_{i}\}_{i=1}^{d}$ represent the canonical basis of $\mathbb{R}^{d}$. An important definition is in order:
Definition 1.
(Consistency and monotonicity)
  1. i.
    We say that the numerical Hamiltonian $\widehat{H}(x,p^{+},p^{-})$ is consistent if $\widehat{H}(x,p,p)\equiv H(x,p)$,   $p\in\mathbb{R}^{d}.$
  2. ii.
    We say that the numerical Hamiltonian $\widehat{H}(x,p^{+},p^{-})$ is monotone in $\Omega$ if the function
    $\left(u(x),\{u(x+\delta e_{i})\}_{i=1}^{d},\{u(x-\delta e_{i})\}_{i=1}^{d}\right)\longmapsto\widehat{H}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))$
    is non-decreasing with respect to $u(x)$ and non-increasing with respect to $u(x+\delta e_{i})$ and $u(x-\delta e_{i})$.
The Lax-Friedrichs scheme
In this paper, we work with the Lax-Friedrichs scheme, given by
$\widehat{H}_{\alpha}(x,p^{+},p^{-}):=H\left(x,\dfrac{p^{+}+p^{-}}{2}\right)-\alpha\sum_{i=1}^{d}\dfrac{p_{i}^{+}-p_{i}^{-}}{2},\qquad\text{for}\ \alpha>0.$ (4.1)
This numerical scheme is consistent with $H(x,p)$, and for any $L>0$, if one takes $\alpha$ satisfying
$\alpha\geq C_{H}(L):=\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|,$ (4.2)
then $\widehat{H}_{\alpha}(x,p^{+},p^{-})$ is monotone at any function $u$ with Lipschitz constant $L$.
For any $\delta>0,$ let us define the uniform grid
$\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega,$
and consider the residual functional
^$\widehat{\mathcal{R}}(u):=\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}.$ (4.3)
Global minimisers
We note that a function $u\in C(\overline{\Omega})$ is a global minimiser of ^$\widehat{\mathcal{R}}(u)$ if and only if
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0\qquad x\in\Omega_{\delta}.$ (4.4)
In other words, $u$ is a global minimiser of ^$\widehat{\mathcal{R}}(u)$ if and only if the restriction of $u$ to $\Omega_{\delta}$ solves the discrete equation associated with the numerical Hamiltonian $\widehat{H}_{\alpha}$.
If $\widehat{H}_{\alpha}$ is consistent with $H$ and monotone, one can ensure (see [9, 3]) that any minimiser of ^$\widehat{\mathcal{R}}(u)$ approximates a viscosity solution of the Hamilton-Jacobi equation
$H(x,\nabla u)=0,\qquad\text{in}\ \Omega,$ (4.5)
in the following sense: for any decreasing sequence $\{\delta_{n}\}_{n\geq 1}$ with $\delta_{n}\to 0^{+}$, and an equicontinuous sequence of functions $u_{n}\in C(\overline{\Omega})$, each of them minimising ^$\widehat{\mathcal{R}}(u)$ defined on $\Omega_{\delta_{n}}$. If $u_{n}$ converge point-wise to some function $u^{\ast}\in C(\overline{\Omega})$, then $u^{\ast}$ is a viscosity solution of (4.5).
With the above observation, we accomplished the first requirement for the functional $\mathcal{R}(\cdot)$, i.e. its global minimisers approximate viscosity solutions to Hamilton-Jacobi equation in (1.1). The second requirement is more subtle and is the main contribution of this paper. Namely, we provide sufficient conditions on $\alpha$ relative to the numerical Hamiltonian $\widehat{H}_{\alpha}(x,p^{+},p^{-})$ and the grid spacing $\delta$ ensuring that any critical point of the functional ^$\widehat{\mathcal{R}}(u)$ defined in (4.3) is indeed a global minimiser. Hence, it approximates a viscosity solution of (4.5).
Critical points
Let $N_{\delta}$ denote the number of grid nodes in $\Omega_{\delta}$. For any $u\in C(\overline{\Omega})$, let $U:=u|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}$ denote the vector corresponding to the values of $u$ on $\Omega_{\delta}$ following a chosen ordering of the grid nodes in $\Omega_{\delta}$. We define the functional $F:\mathbb{R}^{N_{\delta}}\mapsto\mathbb{R},$ given by
$F(U):=\widehat{\mathcal{R}}(u)\quad\text{for any}\leavevmode\nobreak\ u\in C(\overline{\Omega})\ \text{such that}\ U=u|_{\Omega_{\delta}}.$
In view of (1.3), the function $F(\cdot)$ is differentiable in $\mathbb{R}^{N_{\delta}}$, and we denote its gradient by $\nabla F(U)\in\mathbb{R}^{N_{\delta}}$.
Then, the Fréchet differential of $\widehat{\mathcal{R}}(\cdot)$ at any $u\in C(\overline{\Omega})$ can be written as the linear functional on $C(\overline{\Omega})$ given by
$\phi\in C(\overline{\Omega})\longmapsto\Phi\cdot\nabla F(U),\quad\text{where}\ \Phi:=\phi|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}.$
Let us define the residual function $w\in C(\overline{\Omega})$ associated to $u$ and $\widehat{H}_{\alpha}$ as
$w(x)=\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x)),\qquad\forall u\in C(\overline{\Omega}).$
From the form of ^$\widehat{\mathcal{R}}(u)$ in (4.3), which is a least-squares like functional, the gradient of $F$ has the following form
$\nabla F(U)=A_{\alpha,\delta}(U)W,$ (4.6)
where $W:=w|_{\Omega_{\delta}}\in\mathbb{R}^{N_{\delta}}$, and $A_{\alpha,\delta}(U)$ is a linear bounded operator $\mathbb{R}^{N_{\delta}}\to\mathbb{R}^{N_{\delta}}$, depending on $\widehat{H}_{\alpha}$ and $U$ (see Section 7 for the explicit form). Then, for any $u\in C(\overline{\Omega})$ critical point of $\widehat{\mathcal{R}}(\cdot)$, the vector of residuals $W$ satisfies the equation
$A_{\alpha,\delta}(U)W=0.$ (4.7)
This linear system can be seen as a finite-difference equation for the residual function $w(x)$ on the grid $\Omega_{\delta}$. It can be interpreted as the adjoint equation associated with the finite difference system (4.4). Our next goal is to prove that by choosing suitable parameters $\alpha$ and $\delta$ in the numerical Hamiltonian, one can ensure that the adjoint equation (4.7) has a unique solution, which is the trivial solution $W=0$. This implies that any critical point of ^$\widehat{\mathcal{R}}(u)$ solves the finite-difference equation (4.4).
4.1 A sufficient condition for residuals to be zero
In the following theorem, for any constant $L>0$, we give a sufficient condition on $\alpha$ and $\delta$, ensuring that any critical point $u$ of ^$\widehat{\mathcal{R}}(u)$ with Lipschitz constant $L$ solves the finite-difference equation (4.4). We consider the domain $\Omega$ to be a $d$-dimensional cube for simplicity. The result can be adapted to more general domains, satisfying suitable regularity properties, albeit with some additional technicalities.
Theorem 1.
Let $H$ be a Hamiltonian satisfying (1.3), and let $\Omega=(0,1)^{d},$ $d\in\mathbb{N}$. Consider the uniform Cartesian grid $\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega$, for some $\delta=1/N$ and $N\in\mathbb{N}$, and the functional $\widehat{\mathcal{R}}(\cdot)$ defined in (4.3).
Let $u\in C(\overline{\Omega})$ be a critical point of $\widehat{\mathcal{R}}(\cdot)$ with a Lipschitz constant $L$. If
$2\alpha\sin^{2}\left(\frac{\pi}{2}\delta\right)>\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|$ (4.8)
then
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall x\in\Omega_{\delta}.$ (4.9)
The proof is given in Section 7. Essentially, it shows that if $\alpha,\delta$ and $L$ satisfy (4.8), then the linear operator $A_{\alpha,\delta}(U)$ in (4.6) associated to $\widehat{H}_{\alpha}$ is invertible for any function $u$ with Lipschitz constant equal or smaller than $L$. This implies that if the first-order optimality condition (4.7) holds, then the residuals $W$ associated with the function $u$ on $\Omega_{\delta}$ satisfy $W=0$, and hence, $u$ solves the system of equations (4.9).
We further note that condition (4.8) can be written as
$\alpha\dfrac{\lambda_{1}(\Omega_{\delta})}{2d}>\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|,$ (4.10)
where $\lambda_{1}(\Omega_{\delta})$ is the smallest eigenvalue of the matrix associated with the discrete Dirichlet Laplacian on the computational domain $\Omega_{\delta}=\delta\mathbb{Z}\cap\Omega$. A similar result to Theorem 1 can be proved for a more general class of domains $\Omega$ satisfying suitable regularity conditions. However, we have restricted ourselves to the case of a $d$-dimensional cube for the sake of clarity of the arguments.
Since our goal is to use an artificial neural network to approximate the solution of a given Hamilton-Jacobi problem, we do not need to commit ourselves to one fixed grid as in a standard finite difference method. Therefore, we consider the functional $\mathcal{R}(\cdot):C(\overline{\Omega})\to\mathbb{R}^{+}$ given by
[…]
Note that
[…]
where ^$\widehat{\mathcal{R}}(u;z)$ is a discrete functional similar to (4.3), but on the shifted grid $\Omega_{\delta}(z):=(\delta\mathbb{Z}^{d}+z)\cap\Omega$ for $z\in(0,\delta)^{d}$.
By minimising $\mathcal{R}(\cdot)$, one solves the associated finite-difference equation on a family of shifted Cartesian grids, covering the entire domain $\Omega$ (instead of on a fixed grid). This fact is then exploited in Algorithm 2, which minimizes $\mathcal{R}(u)$ for a decreasing sequence of $\delta$.
We have the following result for the functional $\mathcal{R}(u)$.
Theorem 2.
Let $H$ be a Hamiltonian satisfying (1.3), and let $\Omega=(0,1)^{d}$, $d\in\mathbb{N}$. For any fixed $\alpha>0$ and $N\in\mathbb{N}$, set $\delta=1/(N-1)$ and consider the functional $\mathcal{R}(\cdot)$ defined in (4.11).
Let $u\in C(\overline{\Omega})$ be any critical point of $\mathcal{R}(\cdot)$ with a Lipschitz constant $L$. If (4.8) holds, then
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall x\in\Omega.$ (4.12)
The proof is given in Section 7.3.
We make the following remarks concerning the above results:
  1. i.
    Theorems 1 and 2 are local results in the sense that they only ensure that a critical point $u\in C(\overline{\Omega})$ of $\widehat{\mathcal{R}}(\cdot)$ solves the finite-difference equation (4.9), provided that the Lipschitz constant of $u$ is smaller or equal than $L$. They do not rule out the possibility of critical points with larger Lipschitz constants failing to satisfy (4.9). However, one can take $L>0$ as large as needed, at the expense of increasing $\alpha$ and $\delta$ in such a way that (4.8) holds.
  2. ii.
    Choosing large values for $\alpha$ and $\delta$ has the drawback of larger errors in approximating the Hamilton-Jacobi equation due to the numerical diffusion. In a sense, there is no surprise in Theorem 1. If $\alpha\sim\mathcal{O}(\delta^{-2})$, the resulting numerical scheme is consistent with the Laplace equation! This is a linear equation, and thus, the associated least square principle is a strictly convex functional. Therefore, one should only use this condition away from the asymptotic regime. Introducing a lot of numerical diffusion, implies that minimisers of $\widehat{\mathcal{R}}(\cdot)$ (resp. $\mathcal{R}(\cdot)$) are over-regularised approximations of the actual viscosity solution. Since we will use an iterative gradient-based method to minimise the functional (as it is typically done in Deep Learning), we can choose large values of $\alpha$ and $\delta$ in the initial stage of optimization to ensure global convergence to the minimiser and successively reduce $\alpha$ and $\delta$ in later stages. See section 5.1 for a detailed description of the proposed algorithm.
  3. iii.
    A natural question is whether a continuous solution to (4.12) exists at all. There is indeed a constructive way to prove existence of such continuous solutions. Consider the family of grids $\Omega_{\delta}(z):=(\delta\mathbb{Z}^{d}+z)\cap\Omega$, for $z\in[0,\delta)^{d}$. This is simply a shifted version of the original grid $\Omega_{\delta}$. Now, for each $z$, there is a solution to the discrete equation (4.9) on the grid $\Omega_{\delta}(z)$, and this solution depends continuously on the parameter $z$. Then, one can construct the function $u(x)$ which associates, to every $x\in\Omega$, the solution of the discrete equation (4.9) on the unique grid $\Omega_{\delta}(z)$ such that $x\in\Omega_{\delta}(z)$. This function solves (4.12) and is continuous by construction.
4.2 Uniqueness of critical points by regularisation of the Hamiltonian
In Theorem 1 (resp. Theorem 2), we provide a sufficient condition on the choice of the parameters $\alpha$ and $\delta$, ensuring that any critical point of the functional ^$\widehat{\mathcal{R}}(u)$ (resp. $\mathcal{R}(u)$) satisfies
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,\qquad\forall x\in\Omega_{\delta}\ \text{(resp. $\Omega$).}$. (4.13)
This condition (see (4.8)) involves taking $\alpha$ and $\delta$ sufficiently large so that the system of equations associated with the first-order optimality condition, which is of the form $A_{\alpha,\delta}(U)W=0,$ admits a unique solution, i.e. the trivial solution $W=0$. See Section 7 for further details.
An alternative way to ensure that any critical point of the functional solves the finite-difference equation (4.13) is by adding a monotonically increasing term to the numerical Hamiltonian. Let us consider the functional
$u\in C(\overline{\Omega})\longmapsto\widehat{\mathcal{R}}(u):=\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))+\tau u(x)\right]^{2},$
for some $\tau>0$. In this case, the first-order optimality condition is of the form
$(A_{\alpha,\delta}(U)+\tau I_{\Omega_{\delta}})W=0,$
where $I_{\Omega_{\delta}}$ is the identity operator in the space of grid functions on $\Omega_{\delta}$, and $W$ is the grid function associated to the function $x\mapsto w(x):=\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))+\tau u(x)$.
Similar arguments to the ones in the proofs of Theorems 1 and 2 can be used to obtain the sufficient condition on the parameters $\alpha,\delta$ and $\tau$, i.e.
$2\alpha\sin^{2}\left(\frac{\pi}{2}\delta\right)+\frac{\tau}{d}>\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|,$ (4.14)
ensuring that the matrix associated with the linear operator $A_{\alpha,\delta}(U)+\tau I_{\Omega_{\delta}}$ is invertible, for any grid function $U$ on $\Omega_{\delta}=\delta\mathbb{Z}\cap\Omega$ with Lipschitz constant $L$. This in turn implies that, for any function $u$ with Lipschitz constant $L$, if $u$ is a critical point of ^$\widehat{\mathcal{R}}(u)$ then $u$ satisfies the finite-difference equation
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))+\tau u(x)=0,\qquad\forall x\in\Omega_{\delta}.$
However, in this paper, we predominantly consider the case $\tau\equiv 0$.
4.3 The least square principle involving a numerical Hamiltonian and supervised data
In the above section, we discussed functionals involving a numerical Hamiltonian. We show that under appropriate conditions, any global minimiser approximates a viscosity solution of the Hamilton Jacobi equation
$H(x,\nabla u)=0,\qquad\text{in}\ \Omega.$ (4.15)
The uniqueness of the viscosity solution to the partial differential equation (4.15) typically involves additional (suitable) boundary conditions. Likewise, our ultimate least square formulation has to incorporate boundary data.
For a given $\delta>0$, let us consider the uniform Cartesian grid $\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega$, where $\delta=1/N$ for some $N\in\mathbb{N}$. Similarly, we define the numerical boundary as
$\partial\Omega_{\delta}:=\{x\in\delta\mathbb{Z}^{d}\setminus\Omega_{\delta}\,:\ \min_{y\in\partial\Omega}|x-y|<\sqrt{d}\leavevmode\nobreak\ \delta\}.$ (4.16)
In order to set the boundary condition on the numerical boundary, the function $g:\partial\Omega\to\mathbb{R}$ in (4.5) needs to be extended to a $\delta$-neighbourhood of the boundary, defined as
$\Gamma_{\delta}:=\{x\in\mathbb{R}^{d}\setminus\Omega\,:\ \min_{y\in\partial\Omega}|x-y|<\sqrt{d}\leavevmode\nobreak\ \delta\}.$ (4.17)
Since $g$ is a continuous function, the extension of $g$ to $\Gamma_{\delta}$ can be done in a continuous way.
For the parameters $\alpha,\gamma>0$, we define the functional
$\widehat{\mathcal{J}}(u):=\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}+\gamma\delta^{d-1}\sum_{x\in\partial\Omega_{\delta}}\left(u(x)-g(x)\right)^{2},$ (4.18)
where $\widehat{H}_{\alpha}$ is the Lax-Friedrichs numerical Hamiltonian defined in (4.1) and $g$ is the boundary condition in (1.1), continuously extended to $\Gamma_{\delta}$. The following result is a consequence of Theorem 1.
Corollary 1.
Let $H$ be a Hamiltonian satisfying (1.3), $\Omega=(0,1)^{d}$, $d\in\mathbb{N}$, and $g\in C(\partial\Omega)$. For any $N\in\mathbb{N}$, set $\delta=1/N$ and the uniform Cartesian grid $\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega$, with the associated numerical boundary $\partial\Omega_{\delta}$ defined in (4.16), and consider a continuous extension of $g$ to $\Gamma_{\delta}$. For the parameters $\alpha,\gamma>0$, consider the functional $\widehat{\mathcal{J}}(\cdot)$ defined in (4.18). Let $u\in C(\overline{\Omega})$ be any critical point of $\widehat{\mathcal{J}}(\cdot)$ with Lipschitz constant $L$. If (4.8) holds, then
$\begin{cases}\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0,&x\in\Omega_{\delta},\\ u(x)=g(x),&x\in\partial\Omega_{\delta}.\end{cases}$ (4.19)
By the monotonicity of the numerical Hamiltonian, the system (4.19) has at most one solution. Hence, the above corollary proves that the functional $\widehat{\mathcal{J}}(\cdot)$ has at most one critical point with Lipschitz constant $L$. This uniqueness result is up to the continuous extension of $g$ from $\partial\Omega$ to $\Gamma_{\delta}$. That being said, note that in the limit when $\delta\to 0^{+}$, the choice of this extension is irrelevant.
The proof of this corollary can be found in Section 7.2. Under the Assumption 1, and choosing $\alpha>0$ such that the numerical Hamiltonian $\widehat{H}_{\alpha}$ is monotone, Corollary 1 can be used in the following way: if one is able to construct an equicontinuous sequence $u_{n}\in C(\overline{\Omega})$ minimising the functional $\widehat{\mathcal{J}}(\cdot)$, with $\delta_{n}\to 0^{+}$, then $u_{n}(x)\to u^{\ast}(x)$ uniformly in $\Omega$, where $u^{\ast}$ is the unique viscosity solution to (1.1). Of course, the condition (4.8) only holds if $\delta$ is large enough. As we will see in Section 5, one can start by training a neural network with a large value of $\delta$, ensuring that any critical point satisfies (4.19), and then re-train the neural network with decreasing values of $\delta$.
A similar result holds for the functional on $C(\overline{\Omega})$ given by
[…]
where $\rho$ and $\mu$ are probability measures on $\Omega$ and $\Gamma_{\delta}$ respectively, and $\Gamma_{\delta}$ is defined in (4.17).
In this paper, we propose to minimize such least square principle via a stochastic gradient descent algorithm. This means that the integrals will be approximated by the Monte Carlo method. The involvement of the Monte-Carlo method requires deciding on a probability distribution $\rho$ over $\Omega$, and, as we will show in Section 6.2, the choice of this distribution can depend on the error metric of choice.
Ultimately, we shall construct a sequence $u_{k}$ that minimizes $\mathcal{J}$ for a sequence of $\delta_{k}$ that decreases to 00. Also, since we shall consider functions given by a NN, which are continuous by construction, and the continuous extension of the boundary condition to $\Gamma_{\delta}$ is irrelevant in the limit $\delta\to 0^{+}$, for convenience, we will consider
[…]
where $\mu$ is a probability measure on $\partial\Omega$.
Incorporating additional labelled data
Finally, in many applications, labelled data is available in the interior of $\Omega$. Let us denote this interior set by $\Gamma\subset\Omega$, and the associated data by the function $h\in C(\Gamma)$. We can of course introduce an additional data-fitting term to form a least square principle as follows:
$\mathcal{J}(u):=\int_{\Omega}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}d\rho(x)+\gamma_{1}\int_{\partial\Omega}\left[u(x)-g(x)\right]^{2}d\mu(x)+\gamma_{2}\int_{\Gamma}\left[u(x)-h(x)\right]^{2}d\nu(x),$ (4.21)
where $\rho$ is a probability measure on $\Omega$, $\mu$ and $\nu$ are probability measures on $\partial\Omega$ and $\Gamma$.
The addition of additional labelled data, a.k.a. supervised data, changes the optimization landscape and can influence the properties of the critical points. Indeed, if one interprets $\Gamma$ as part of the boundary of $\Omega$, the associated discrete Laplacian has larger eigenvalues, and one may employ smaller values of $\delta$, while (4.10) still holds true. See Section 6.3 for experiments showing this phenomenon.
4.4 Time dependent problems
There is no difficulty in extending our analysis to the time-dependent case:
$\begin{cases}\partial_{t}u+H(x,\nabla u)=0&\Omega\times(0,T),\\ u(x,t)=g(x)&\partial\Omega\times(0,T),\\ u(x,0)=u_{0}(x)&\Omega,\end{cases}$ (4.22)
with $T\in(0,\infty]$, initial data $u_{0}:\Omega\to\mathbb{R}$ and boundary data $g:\Omega\to\mathbb{R}$.
In this case, a suitable choice for the least squares approach is to use the following finite-difference scheme
$u(x,t+\delta_{t})=u(x,t)-\delta_{t}\widehat{H}_{\alpha}(x,D_{\delta_{x}}^{+}u(x,t),D_{\delta_{x}}^{-}u(x,t)),$ (4.23)
for some $\delta_{t},\delta_{x},\alpha>0$, which combines an explicit Euler scheme for the time derivative and a Lax-Friedrichs numerical Hamiltonian.
The least square finite-difference functional associated to the above numerical scheme is
[…]
and can be coupled with the initial and boundary data as
[…]
A similar analysis of the critical points can be carried out for the above functional. Again, the conclusion is that by taking $\alpha$ and $\delta_{x}$ sufficiently large, one can ensure that any critical point satisfies the finite-difference equation associated with the numerical scheme (4.23). In Section 5.2.2, we show some numerical experiments, where we apply our least squares approach to time evolution problems of the form (4.22).
5 Numerics
As discussed in the introduction, the purpose of formulating the numerical method for (1.1) as a minimisation problem is that one can approximate the solution using a Neural Network trained through stochastic gradient descent (SGD). To enhance converge to the viscosity solution, we use the residual of a consistent and monotone Lax-Friedrichs numerical Hamiltonian in the loss functional, which ensures that any global minimiser approximates a viscosity solution. Due to the use of stochastic gradient descent, it is of utmost importance that any critical point of the function is a global minimiser. However, in view of our main results Theorem 1 and Theorem 2, we can only ensure this by considering a numerical Hamiltonian that incorporates enough numerical diffusion ($\alpha$ and $\delta$ big enough in (4.3) or (4.11)). The drawback is that the global minimiser of the resulting minimisation problem is an over-smoothed approximation of the actual viscosity solution. Next, we describe an algorithm to approximate the global minimiser of a Lax-Friedrichs finite-difference loss with small parameters $\alpha$ and $\delta$. The idea is to first apply SGD with large values of $\alpha$ and $\delta$, ensuring convergence to the global minimiser, and then decrease these parameters to carry on with more SGD iterations and refine the approximation of the solution.
5.1 The proposed algorithms
Let us consider any NN function $\Phi:\overline{\Omega}\times\Theta\to\mathbb{R}$, where $\Theta\subset\mathbb{R}^{P}$ is the parameter space and $P$ is the number of parameters. We define the parameterized family of functions
:=$\mathcal{F}:=\{x\mapsto\Phi(x;\theta)\ :\quad\theta\in\Theta\}\subset C(\overline{\Omega}).$ (5.1)
Our numerical solution of (1.1) will be a function of the form $u(x)=\Phi(x;\theta^{\ast})$, where $\theta^{\ast}\in\Theta$ is obtained by minimising the function
$\theta\longmapsto\mathcal{J}(\Phi(\cdot;\theta)),$ (5.2)
and $\mathcal{J}$ is defined in (4.21) (or in (4.25) for time-dependent Hamilton-Jacobi equations). The minimisation algorithm to obtain $\theta^{\ast}$ is a version of SGD in which, at every iteration, the functional $\mathcal{J}$ in (4.21) is approximated by a Monte Carlo approximation of the integrals.
Let $\mathcal{D}_{0}$, $\mathcal{D}_{b_{1}}$, and $\mathcal{D}_{b_{2}}$ be the chosen probability distributions over $\Omega$, $\partial\Omega$, and $\Gamma$ respectively. Our algorithms are built on the typical stochastic gradient descent algorithm:
$\theta^{k+1}=\theta^{k}-\eta_{k}\left(\nabla_{\theta}\widetilde{\mathcal{L}}(\Phi(\cdot;\theta);\gamma_{1},\gamma_{2})+\widetilde{\mathcal{R}}(\Phi(\cdot;\theta))\right),$ (5.3)
with
~$\widetilde{\mathcal{L}}(\Phi(\cdot;\theta);\gamma_{1},\gamma_{2})$ ::=:= $\dfrac{\gamma_{1}}{N_{b_{1}}}\sum_{x\in X_{b_{1}}}\left(\Phi(x;\theta)-g(x)\right)^{2}+\dfrac{\gamma_{2}}{N_{b_{2}}}\sum_{x\in X_{b_{2}}}\left(\Phi(x;\theta)-h(x)\right)^{2},$
~$\widetilde{\mathcal{R}}(\Phi(\cdot;\theta))$ ::=:= $\dfrac{1}{N_{0}}\sum_{x\in X}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}.$
Here, $X$ contains $N_{0}$ i.i.d. samples from $\mathcal{D}_{0}$, $X_{b_{1}}$ contains $N_{b_{1}}$ i.i.d. samples from $\mathcal{D}_{b_{1}}$, and $X_{b_{2}}$ contains $N_{b_{2}}$ i.i.d. samples from $\mathcal{D}_{b_{2}}$. The sets $X,X_{b_{1}},$ and $X_{b_{2}}$ are resampled in each iteration. In the Deep Learning community, the i.i.d. samples $X\subset\Omega$, $X_{b_{1}}\subset\partial\Omega$ and $X_{b_{2}}\subset\Gamma$ are often referred to as mini-batches, and $N_{0}$, $N_{b_{1}}$ and $N_{b_{2}}$ denote the batch-size for each mini-batch.
A possible choice for the probability distributions $\mathcal{D}$, $\mathcal{D}_{b_{1}}$ and $\mathcal{D}_{b_{2}}$ are the uniform distributions over $\Omega$, $\partial\Omega$, and $\Gamma$ respectively. However, as we observe in the numerical experiments in section 6.2, other distributions can be more efficient, especially in high-dimensional domains.
Algorithm 1 SGD for Lax-Friedrichs finite-difference loss
$\gamma_{1},\gamma_{2}>0$, $\{\eta_{k}\}_{k=0}^{K}\in(0,\infty)^{K}$, $N_{0},N_{b_{1}},N_{b_{2}}\in\mathbb{N}$, $\mathcal{D}_{0}\in\mathcal{P}(\Omega)$, $\mathcal{D}_{b_{1}}\in\mathcal{P}(\partial\Omega)$, and $\mathcal{D}_{b_{2}}\in\mathcal{P}(\Gamma)$.
Input: $\theta_{0}\in\Theta$ and $\alpha,\delta>0$.
procedure SGDLxF($\theta_{0},\alpha,\delta$)
     $k\leftarrow 0$
     while $k<K$ do
         $X\leftarrow$ $N_{0}$ i.i.d. samples from $\mathcal{D}_{0}$
         $X_{b_{1}}\leftarrow$ $N_{b_{1}}$ i.i.d. samples from $\mathcal{D}_{b_{1}}$
         $X_{b_{2}}\leftarrow$ $N_{b_{2}}$ i.i.d. samples from $\mathcal{D}_{b_{2}}$
         $\theta^{k+1}:=\theta^{k}-\eta_{k}\left(\nabla_{\theta}\widetilde{\mathcal{L}}(\Phi(\cdot;\theta^{k});\gamma_{1},\gamma_{2})+\nabla_{\theta}\widetilde{\mathcal{R}}(\Phi(\cdot;\theta^{k}))\right)$
         $k\leftarrow k+1$
     end whilereturn $\theta^{K}$
end procedure
If the NN is initialised to be close to a global minimiser, then the gradient descent iterations will converge to a global minimiser. By restricting the parameters to a compact set, we can estimate the Lipschitz constant $L$. We can choose $\alpha:=\alpha_{0}>0$ and $\delta:=\delta_{0}>0$ accordingly to meet the condition set in (4.8) for the Lax-Friedrichs numerical Hamiltonian $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))$ from (4.1).
In view of Theorem 2, we can ensure that any critical point of
[…]
with Lipschitz constant $L>0$ satisfies
$\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0\qquad\text{forall}\quad x\in\Omega.$ (5.4)
Hence, the gradient descent iterations will converge to a function that solves (or approximately solves) the above equation. Once the NN is trained, we can reduce the hyper-parameters $\alpha,\delta$ by choosing
$0<\alpha:=\alpha_{1}\lesssim\alpha_{0}\quad\text{and}\quad 0<\delta:=\delta_{1}\lesssim\delta_{0}$
and continue with the gradient descent iterations applied to the new functional $\mathcal{R}_{\alpha_{1},\delta_{1}}(u)$. Since the functional $\mathcal{R}(\cdot)$ depends continuously on the parameters $\alpha$ and $\delta$, and the solutions of (5.4) also depend continuously on these parameters, we can ensure that, by initialising the NN close to a global minimiser of $\mathcal{R}_{\alpha_{0},\delta_{0}}(u)$, the iterations will converge to a global minimiser of $\mathcal{R}_{\alpha_{1},\delta_{1}}(u)$. This procedure can be iterated with a reduction schedule for $\alpha$ and $\delta$. Whenever we reduce the parameters $\alpha$ and $\delta$, we obtain a function closer to the viscosity solution of (1.1). This procedure can be interpreted as an approximation of the vanishing viscosity method, in which the viscosity solution to the boundary value problem (1.1) is obtained as the limit $\varepsilon\to 0^{+}$ of the viscous approximation (1.4).
Algorithm 2 Optimisation for a Lax-Friedrichs finite-difference loss with a progressive refinement schedule
$M\in\mathbb{N}$, $L>0$, $\mathcal{F}$ as in (5.1) and $\{(\alpha_{m},\delta_{m})\}_{m=0}^{M}\in(0,\infty)^{2M}$.
$(L,\alpha_{0},\delta_{0})$ satisfy (4.8) and $\alpha_{m+1}\leq\alpha_{m}$ and $\delta_{m+1}\leq\delta_{m}$ for all $1\leq m\leq M$.
Initialise $\theta_{0}\in\Theta$ in a way that $x\mapsto\Phi(x;\theta)$ has Lipschitz constant $L$.
$m\leftarrow 0$
while $m<M$ do
     $\theta_{m+1}:=\operatorname{SGDLxF}(\theta_{m},\alpha_{m},\delta_{m})$ as defined in Algorithm 1
     $m\leftarrow m+1$
end whilereturn $\theta_{M}$
Relating $\delta$, the total number of collocations $N$ and the dimension $d$
In this paper, we argued that using a finite difference scheme allows us to regulate the network for the viscosity solution. The effectiveness of the regularization directly depends on the probability of the finite difference stencil sampling over the region where the characteristics should collide. This region corresponds to the location where the viscosity solution is not differentiable and is typically lower dimensional. On the one hand, for a fixed number of collocation points randomly placed in the domain, larger $\delta$ in $\hat{H}$ increases the probability of the finite difference stencil regulating over that region. On the other hand, finite difference solutions with larger values of $\delta$ have worse approximation errors due to numerical diffusion.
In the following, we derive a lower bound on $\delta$ for a given total number of collocation points, $N$ used in training and the dimensionality of the problem.
We recall the following estimates derived in [21]. Assuming that $\mu$ has a density $\rho(x)$. Then, with a probability greater than $p_{0}$, the diameter of the Voronoi cell $V(x)$ for $x\in D_{N}$ can be bounded as follows:
$\text{diam}(V(x))\leq 3\sqrt{d}\left(\frac{C_{1}+C_{2}\sqrt{2+N\ln(\frac{2}{1-p_{0}})}}{\rho(x)N}\right)^{1/d},$ (5.5)
where $C_{1},C_{2}$ are two positive constants independent of $d$ and $N$. The stencil of a finite difference scheme applied to $x$ will reach different Voronoi cells if $\delta>V(x)$. Therefore, the range of influence of the finite difference residual will cover $\Omega$ if
$\delta>\max_{x\in D_{N}}\text{diam}(V(x)).$ (5.6)
The benefits of resampling
Resampling of collocation points during the training process is essential to control the generalization error. It is also a simple way to avoid overfitting.
As we mentioned earlier, collocating a ReLU-activated network with a finite-difference equation directly regulates the network in the neighbourhood of each collocation point. Thus, the resampling of collocation points effectively increases the total number of collocation points used in regulating the network while having much fewer gradient evaluations in each training step. The resulting approximation will have a more uniformly distributed error.
Finally, as we showed above, having a larger total number of collocation points allows us to use smaller $\delta$ in the finite differences to improve the accuracy of the approximations.
5.2 Computational experiments
We test the efficacy of our algorithm on various canonical HJ equations posed in different dimensions. We start with the Eikonal equation posed on simple domains (Section 5.2.1). Then, we consider a time-evolution HJ equation of the form (4.22) (Section 5.2.2). Finally, we consider HJ equations arising in control problems and differential games associated with curvature-constrained dynamics. In particular, we consider two problems associated with Reeds-Shepp’s car model: the shortest path to a given target (Section 5.2.3) and the pursuit-evasion game (Section 5.2.4). In Section 6, we analyse, through numerical tests, various aspects of our proposed algorithm, such as data efficiency and the choice of the distribution for the collocation points in the PDE domain.
In all the experiments, we use a fully connected multi-layer perceptron (MLP) as NN architecture. For a chosen number of hidden layers $\ell\in\mathbb{N}$ (NN depth) and a vector $[d_{1},\ldots,d_{\ell}]\in\mathbb{N}^{\ell}$ representing the number of neurons in each hidden layer, the network function can be written as
$\Phi(x;\theta)=W_{\ell}y_{\ell}+b_{\ell},\quad\text{where}\quad\begin{cases}y_{i}=\sigma\left(W_{i-1}y_{i-1}+b_{i-1}\right),&\text{for}\ i=1,\ldots,\ell\\ y_{0}=x.\end{cases}$ (5.7)
where $\theta:=\{(W_{i},b_{i})\}_{i=0}^{\ell}$ is the parameter of the NN. The matrices $W_{0}\in\mathbb{R}^{d_{1}\times d}$, $\{W_{i}\}_{i=1}^{\ell-1}\in\prod_{i=1}^{\ell-1}\mathbb{R}^{d_{i+1}\times d_{i}}$ and $W_{\ell}\in\mathbb{R}^{d_{\ell}\times 1}$ are typically called weights, and the vectors $\{b_{i}\}_{i=0}^{\ell-1}\in\prod_{i=0}^{\ell-1}\mathbb{R}^{d_{i+1}}$ and the scalar $b_{\ell}\in\mathbb{R}$ are called biases. The function $\sigma(\cdot)\in C(\mathbb{R};\mathbb{R}^{+})$ defined as $\sigma(s):=\max\{0,s\}$ is the activation function, which acts component-wise on vectors.
The NN architecture will be indicated in the experiment description for each case. For instance, in the case of a fully connected MLP with three hidden layers, the architecture is denoted by $[d_{1},d_{2},d_{3}]$, where $d_{1},d_{2}$ and $d_{3}$ represent the number of neurons in each hidden layer. Investigating the optimal choice of the NN architecture is beyond the scope of this paper. Therefore, in all the experiments, we chose relatively simple architectures in which the number of neurons in the hidden layers is kept constant. For the problems in Sections 5.2.3 and 5.2.4, we consider a slightly different architecture because the problem’s domain requires the solution to be $2\pi$-periodic with respect to some of the variables. This is achieved by considering a linear combination of trigonometric functions, for which the coefficients are given by MLPs as defined above.
We consider various examples for which the viscosity solution is available in closed form. Namely, we consider the Eikonal equation in simple domains such as a cube, a ball and an annulus, in different dimensions. In this case, the viscosity solution corresponds to the distance function to the boundary and can be easily computed for the three aforementioned domains. This function is used as a ground truth to evaluate the accuracy of the solution provided by the trained NN.
We use two metrics to evaluate the accuracy: the Mean Square Error (MSE) and the $L^{\infty}$-error. These are defined respectively as
[…]
where $u\in C(\overline{\Omega})$ is the viscosity solution used as ground truth.
Obviously, we cannot compute exactly the integral and the maximum in (5.8). We approximated it by the Monte Carlo method. For some $N\in\mathbb{N}$, let us define
$\widetilde{MSE}(\Phi(\cdot;\theta)):=\dfrac{1}{N}\sum_{i=1}^{N}|\Phi(x_{i};\theta)-u(x_{i})|^{2}$ (5.9)
$\widetilde{E}_{\infty}(\Phi(\cdot;\theta)):=\max_{i=0,\ldots,N}|\Phi(x_{i};\theta)-u(x_{i})|,$ (5.10)
where $\{x_{1},\ldots,x_{N}\}\subset\Omega$ is an i.i.d. uniform sampling in $\Omega$, and $x_{0}=(0,0,\ldots)\in\mathbb{R}^{d}$ is the origin. Note that to approximate the $L^{\infty}$-error we add the origin to the sampling. Of course, this does not apply when $\Omega$ is an annulus. When $\Omega$ is a cube or a ball, we observe that the highest error of the numerical solution is achieved near the origin, which is the furthest point from the boundary. Therefore, especially when the dimension is high, adding this point to the sampling improves the approximation of the $L^{\infty}$-error. In all the experiments below, we took the sample size for the Monte Carlo approximation as $N=10^{6}$.
In the following sections, the run times for the experiments are recorded from running our PyTorch implementation on a desktop Dell with Intel Core i7-14700 CPU at 2.10GHz×$\times 20$ with $15.8$ GiB of RAM. The code is available at the following link:
5.2.1 Eikonal equation
To illustrate the algorithm presented in Section 5.1, we carry out some numerical experiments for the Eikonal equation in various $d$-dimensional domains:
$\begin{cases}|\nabla u(x)|^{2}=1,&x\in\Omega,\\ u(x)=0,&x\in\partial\Omega.\end{cases}$ (5.11)
Let us start with a simple example by considering the domain $\Omega:=(-3,3)^{d}$. The solution, in this case, corresponds to the distance function to the boundary of the $d$-dimensional cube and can be written as
$u(x)=3-\|x\|_{\infty},\qquad\forall x\in(-3,3)^{d}.$ (5.12)
We shall use the explicit solution as a ground truth to evaluate the accuracy of the numerical approximations.
The two-dimensional case ($d=2$).
We use Algorithm 2, with $M=4$, to train a fully connected Neural Network with one hidden layer with $20$ neurons. The parameters $\alpha$ and $\delta$ in each round/iteration are assigned according to the following schedules:
$(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(2.5,2,1.5,1,0.5)\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2},\delta_{3},\delta_{4})=(0.75,0.5,0.3,0.1,0.05).$ (5.13)
Observe that, after each round, both parameters $\alpha$ and $\delta$ are decreased. This is designed to progressively reduce the effect of the numerical viscosity to obtain higher fidelity in the approximate solution in the final round. In each round, we applied Algorithm 1 with batch size $N_{0}=60$ for the collocation points sampled uniformly in $\Omega$, and $N_{b_{1}}=20$ for the boundary points, sampled uniformly in $\partial\Omega$. We did not consider any supervised data in the interior of the domain, i.e. $N_{b_{2}}=0$. Concerning the weighting parameter for the boundary loss, we used $\gamma_{1}=1$. The number of SGD iterations in each of the five rounds is $K=1000$.
We ran 10 numerical experiments with i.i.d. initializations and computed the average and the standard deviation of the Mean Square error (MSE) (5.9) and the $L^{\infty}$-error (5.10) after each round in the training schedule. The results are reported in Figure 3 and Table 1. The solid lines in Figure 3 represent the average error over the 10 experiments, and the shadowed areas represent the interval within one standard deviation. The run time for each experiment was 6 seconds on average. In Figure 2, we can see, for one of the ten experiments, the numerical approximation provided by the NN at initialisation and after each of the five rounds in Algorithm 2.
We see in Table 1 that the standard deviation of the MSE and the $L^{\infty}$-error decreases for smaller values of $\alpha$ and $\delta$. However, if we look at the ratio between the standard deviation and the average error, we see that it increases throughout the training rounds. Indeed, after the first round this ratio is approximately $0.04$ and $0.03$ for the MSE and $L^{\infty}$-error respectively, whereas after the fifth round, these ratios are approximately $0.5$ and $0.35$. This can also be observed in the graphical representation of the numerical experiment in Figure 3, which show a larger variance of the error for later rounds in the training algorithm. We attribute this phenomenon to the fact that for larger $\delta$, each collocation point influences a larger portion of the domain, which leads to an enhanced performance of the training algorithm in terms of data efficiency (see Section 6.1 for a more detailed analysis). The flexibility in using different values of $\delta$ to control the domain of influence for a finite set of collocation points may be another argument supporting the use of a finite-difference-based loss function over the PDE residual, which would correspond to the limit $\delta\to 0^{+}$.
Figure 2: The numerical approximation of the Eikonal equation in $\Omega=(-3,3)^{2}$ constructed by Algorithm 1. The top-left plot corresponds to the NN at initialisation. Then, from left to right and from top to bottom, each plot corresponds to the approximation after every iteration. The values of $\alpha$ and $\delta$ are taken as in (5.13).
Figure 3: (Eikonal equation in a square in $\mathbb{R}^{2}$) Evolution of the Mean Square Error and the $L^{\infty}$-Error after each round of Algorithm 2, for the Eikonal equation in $(-3,3)^{2}$. The parameters $\alpha$ and $\delta$ in each of the five training rounds are taken as in (5.13). The experiment is repeated 10 independent times. The solid line represents the average error, and the shadowed area represents the interval within one standard deviation.
Table 1: (Eikonal equation in a square in $\mathbb{R}^{2}$) Summary of the numerical experiment for the Eikonal equation in $\Omega=(-3,3)^{2}$. In each experiment, we applied 4 iterations of Algorithm 2 with decreasing values of $\alpha$ and $\delta$ as in (5.13). The experiment is repeated 10 times using different random seeds. The error values in the table correspond to the mean and the standard deviation after each training round. See Figure 3 for a graphical representation of the numerical results.
Round 1 2 3 4 5
$\delta$ 0.75 0.5 0.3 0.1 0.05
$\alpha$ 2.5 2 1.5 1 0.5
MSE (avg) 6.45e-02 1.42e-02 1.55e-03 4.11e-05 1.45e-05
MSE (std) 2.68e-03 8.88e-04 1.21e-04 2.09e-05 7.02e-06
$L^{\infty}$-error (avg) 1.01 6.12e-01 2.87e-01 6.07e-02 1.30e-02
$L^{\infty}$-error (std) 2.97e-02 1.96e-02 1.19e-02 6.93e-03 4.49e-03
Higher dimensional cases.
Next, we solve (5.11) in $(-3,3)^{d},$ for $d=5$ and $d=8$. In the experiments, we ran Algorithm 2 with $M=4$ and parameters $\alpha$ and $\delta$ following the schedule
$(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})=(2.5,2,1,0.5)\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2},\delta_{3})=(0.75,0.3,0.1,0.05).$ (5.14)
As we increase the dimension, we need to increase the complexity of the NN and the number of collocation points. Rather than increasing the batch sizes $N_{0}$ and $N_{b_{1}}$ in (5.3), we increase the number of iterations per round. We compared the performance of simple fully connected MLP with 2 or 3 hidden layers.
See Table 2 for a summary of the numerical experiments (each row corresponds to a different experiment). Each experiment is repeated 10 times with different independent random seeds. The error values reported in the table represent the average error and the standard deviation over the 10 independent runs of each experiment. In all the cases, we used a uniform distribution to sample the points in $\Omega$ and $\partial\Omega$ during the iterations of Algorithm 1. The weighting parameter for the boundary loss is taken, again, as $\gamma_{1}=1$. In Figure 4, we see the evolution of the MSE and the $L^{\infty}$-error throughout the application of Algorithm 2, for experiments 4,5 and 6, which correspond to the Eikonal equation in an $8$-dimensional cube. The solid lines represent the average error after each training round, and the shadowed area represents the interval within one standard deviation.
As mentioned in section 5.1, the reason behind decreasing the values of $\alpha$ and $\delta$ throughout the iterations of Algorithm 2, is to improve the accuracy of the NN. However, this does not always produce a more accurate approximation. See in Figure 4 that the variance of the approximation error increases over the iterations of Algorithm 2. We think that this is due to $\delta$ being too small relative to the number of collocation points and the network’s expressive power. See a more detailed discussion in subsection 6.1. This issue can be alleviated by increasing the number of collocation points. Since the collocation points are resampled at every SGD iteration, one can increase the total number of collocation points by increasing the number of SGD iterations.
Table 2: (Eikonal equation in a hypercube in $\mathbb{R}^{d}$) Summary of the numerical experiments for the Eikonal equation in $\Omega=(-3,3)^{d}$ with $d=5,8$. In each experiment, we applied 4 iterations of Algorithm 2 with decreasing values of $\alpha$ and $\delta$ as in (5.14). Each experiment is repeated 10 times using different random seeds. The error values in the table correspond to the mean and the standard deviation. See Figure 4 for the evolution of the MSE and the $L^{\infty}$-error after each iteration. In all the experiments, the batch size in the implementation of SGD in (5.3) is taken as $N_{0}=200$ and $N_{b_{1}}=80$.
Exp. Dimension Architecture Iterations (×$\times 10^{3}$) MSE $L^{\infty}$-error Runtime (s)
1 5 [30, 30] $\{2,2,2,2\}$ 0.013 ± 0.002 0.606 ± 0.079 62
2 5 [30, 30, 30] $\{2,2,2,2\}$ 0.01 ± 0.007 0.527 ± 0.119 85
3 5 [30, 30, 30] $\{2,2.5,3,5\}$ 0.006 ± 0.004 0.435 ± 0.137 133
4 8 [40, 40] $\{3,3,3,3\}$ 0.047 ± 0.006 1.129 ± 0.18 154
5 8 [40, 40, 40] $\{3,3,3,3\}$ 0.031 ± 0.011 0.87 ± 0.131 209
6 8 [40, 40, 40] $\{3,4,5,7\}$ 0.012 ± 0.006 0.751 ± 0.245 331
Figure 4: (Eikonal equation in a hypercube in $\mathbb{R}^{8}$) Evolution of the MSE (left) and the $L^{\infty}$-error (right) throughout the iterations of Algorithm 2, for the Experiments 4, 5 and 6 reported in Table 2. These correspond to the Eikonal equation in an 8-dimensional cube. Each experiment is repeated 10 times with different random seeds. The solid lines represent the average error over the 10 experiments, and the shadowed area represents the interval within one standard deviation.
In Table 3, we can see the results of some numerical experiments for the solution of the Eikonal equation in an annulus of dimension $d=5$ and $10$, with inner radius $2$ and outer radius $6$, i.e.
$\Omega:=\{x\in\mathbb{R}^{d}\ :\quad 2<|x|<6\}.$
We applied $4$ rounds of Algorithm 2 with decreasing values of $\alpha$ and $\delta$, following the training schedule in (5.14). See Figure 5 for the evolution of the MSE and the $L^{\infty}$-error of the solution after each iteration. See Figure 6 for a representation of the numerical solution in a central cross-section of the domain.
[…]
Exp. Dim. Architecture Iterations (×$\times 10^{3}$) MSE $L^{\infty}$-error Runtime (s)
1 5 [40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.019 ± 0.007 0.611 ± 0.092 62
2 5 [40, 40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.009 ± 0.003 0.434 ± 0.078 90
3 5 [40, 40, 40] $\{1.5,2,2.5,3\}$ 0.012 ± 0.013 0.396 ± 0.054 135
4 10 [60, 60] $\{2.5,2.5,2.5,2.5\}$ 0.026 ± 0.003 0.797 ± 0.043 80
5 10 [60, 60, 60] $\{2.5,2.5,2.5,2.5\}$ 0.015 ± 0.004 0.608 ± 0.059 103
6 10 [60, 60, 60] $\{2.5,3,3.5,5\}$ 0.014 ± 0.002 0.552 ± 0.04 137
Figure 5: (Eikonal equation in an annulus in $\mathbb{R}^{10}$) Evolution of the MSE (left) and the $L^{\infty}$-error (right) throughout the iterations of Algorithm 2, for the experiments 4,5 and 6 reported in Table 3. These correspond to the Eikonal equation in a 10-dimensional annulus. Each experiment is repeated 10 times with different random seeds. The solid lines represent the average error over the 10 experiments, and the shadowed area represents the interval within one standard deviation.
Figure 6: Cross-section of the numerical solution provided by the trained NN in one of the ten independent runs of Experiment 6 in Table 3. The NN approximates the viscosity solution of the Eikonal equation in a 10-dimensional annulus.
5.2.2 Time-dependent Hamilton-Jacobi equation
As mentioned in Section 4.4, our method can be applied to time-evolution Hamilton-Jacobi equations of the form (4.22). Let us consider the initial value problem
$\begin{cases}\vspace{5pt}\partial_{t}u+\dfrac{|\nabla u|^{2}}{2}+\dfrac{|x|^{2}}{2}=0,&(t,x)\in(0,T)\times\mathbb{R}^{d},\\ u_{0}(x)=\dfrac{1}{2}\left(\langle x,\,Ax\rangle-1\right),&x\in\mathbb{R}^{d},\end{cases}$ (5.15)
[…]
By the Riccati Theory, see e.g. [14], the solution to (5.15) is given by
$u(t,x)=\dfrac{1}{2}\left(\langle x,\,E(t)x\rangle-1\right),$ (5.16)
where $E(\cdot):[0,T)\to\mathbb{R}^{d\times d}$ is the quadratic form, solution to the Riccati differential equation
$\begin{cases}\dot{E}(t)=-E(t)^{2}-I,&t\in(0,T)\\ E(0)=A,\end{cases}$ (5.17)
where $I$ denotes the identity matrix in $\mathbb{R}^{d}$. The closed-form solution (5.16) for the PDE problem (5.15) will be used as ground truth to evaluate the accuracy of our method.
We consider the bounded domain $\Omega:=(-3,3)^{d}$, and minimize the functional
[…]
where $\mathcal{R}(u)$ is defined as in (4.24), and $\gamma_{0}=1$. Note that we do not use supervised data for the entire boundary of $(0,1/2)\times\Omega$. We only prescribe the initial condition on $\{0\}\times(-3,3)^{d}$. However, due to the convexity of Hamiltonian and the initial condition $u_{0}(\cdot)$, the characteristic curves starting outside $(-3,3)^{d}$ never enter in $(-3,3)^{d}$. This implies that the viscosity solution in $(-3,3)^{d}\times(0,T)$ is uniquely determined by the initial condition in $(-3,3)^{d}$, i.e. the initial value problem
$\begin{cases}\partial_{t}u+H(x,\nabla u)=0&(x,t)\in(-3,3)^{d}\times(0,1/2)\\ u(x,0)=u_{0}(x)&x\in(-3,3)^{d}\end{cases}$ (5.18)
has a unique solution. Then, by choosing a finite-difference functional $\mathcal{R}(u)$ such that any critical point approximates a viscosity solution, one can deduce that any critical point of $\mathcal{J}(u)$ approximates the unique viscosity solution to the initial value problem (5.18).
In Table 4, we report the results of various numerical experiments for the problem with $d=2$ and $d=5$. In all cases, we consider the training schedule
$(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})$ $=(2.5,2,1.5,1)$
$(\delta_{x,0},\delta_{x,1},\delta_{x,2},\delta_{x,3})$ $=(0.5,0.3,0.2,0.1)$ (5.19)
$(\delta_{t,0},\delta_{t,1},\delta_{t,2},\delta_{t,3})$ $=(0.05,0.03,0.02,0.01).$
Each experiment reported in Table 4 has been run 10 times independently, and the errors represent the mean and the standard deviation over the 10 experiments. Figure 7 shows the evolution of the 00-level set of the approximated solution for the two-dimensional case, after each training round (from left to right). These errors have been computed by taking the solution (5.16) as the ground truth. The solution in (5.16) has been computed by solving the Riccati equation (5.17) by means of RK45 method implemented in the Python package Scipy.
Table 4: (Time-evolution Hamilton-Jacobi equation) Summary of the experiments for the initial value problem problem (5.18), with $d=2$ and $d=5$. Each experiment is repeated 10 times with different random seeds. The MSE and the $L^{\infty}$-error is computed with respect to the solution given by (5.16), obtained by numerically solving the Riccati equation (5.17).
Exp. Dim. Architecture Iterations (×$\times 10^{3}$) MSE $L^{\infty}$-error Runtime (s)
1 2 [50, 50] {2, 2, 2, 2} 0.008 ± 0.002 0.672 ± 0.102 60
2 2 [50, 50, 50] {2, 2, 2, 2} 0.005 ± 0.002 0.596 ± 0.088 62
3 2 [50, 50, 50] {2, 3, 5, 8} 0.002 ± 0.001 0.463 ± 0.093 137
4 5 [80, 80] {8, 8, 8, 8} 0.021 ± 0.006 1.579 ± 0.257 294
5 5 [80, 80, 80] {8, 8, 8, 8} 0.012 ± 0.003 1.422 ± 0.323 402
6 5 [80, 80, 80] {8, 10, 15, 20} 0.009 ± 0.002 1.13 ± 0.261 643
Figure 7: (Time-evolution Hamilton-Jacobi equation) For the problem (5.15), with $d=2$, the plots represent the 00-level set of the numerical solution at times $t\in\{0,0.1,0.2,0.3,0.5\}$, provided by an NN corresponding to Exp. 3 in Table 4. From left to right, the plots represent the NN solution (in red) after training with $\alpha,\delta_{x}$ and $\delta_{t}$ as in the schedule (5.19). The green dashed line represents the solution obtained by numerically solving the Riccati Differential equation (5.17). In the plots, the zero-level sets of the approximate solutions, $\{x\in\mathbb{R}^{d}\,:\,\Phi(x,t;\theta)=0\}$, are depicted for $t=0,0.1,0.2,0.3,0.4$ and $0.5$. The label for the curves reveals the corresponding values of $t$.
5.2.3 Shortest path to a target for a Reeds-Shepp’s car
Here, we consider an optimal control problem for a dynamical system with curvature constraints. In particular, we consider Reeds and Shepp’s car model [26], which can move backward and forward with a limited turning radius. The pose of the car is represented by $(x,y,\omega)\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$, where $(x,y)$ represents the spatial position of the centre of mass of the car and $\omega$ represents its orientation angle. Here, $\mathbb{T}_{[0,2\pi)}$ denotes the one-dimensional flat torus of length $2\pi$. The optimal control problem involves steering the car to the origin $(x,y)=0$ from the given initial position $(x_{0},y_{0},\omega_{0})$ in the shortest possible time.
Denoting by $t\mapsto(x(t),y(t),\omega(t))\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$ the trajectory of the car, the dynamics are given by the following ODE:
$\begin{cases}\dot{x}(t)=\sigma a(t)\cos\omega(t)\\ \dot{y}(t)=\sigma a(t)\sin\omega(t)\\ \dot{\omega}(t)=\dfrac{b(t)}{\rho}\\ x(0)=x_{0},\ y(0)=y_{0},\ \omega(0)=\omega_{0},\end{cases}$ (5.20)
where $\sigma>0$ is the maximum speed of the car, $\rho>0$ is the inverse of the angular velocity of the car (the turning radius is therefore $\sigma\rho$) and $(x_{0},y_{0},\omega_{0})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}$ is the initial pose of the car. The time dependent functions $a(\cdot),b(\cdot):[0,\infty)\to[-1,1]$ are the controls of the car.
Given the dynamics of the car, the Hamiltonian associated with the shortest path problem is given by
$H(x,y,\omega,\nabla u)=\sigma|\partial_{x}u\cos\omega+\partial_{y}u\sin\omega|+\dfrac{1}{\rho}|\partial_{\omega}u|-1,$
where $\nabla u=(\partial_{x}u,\partial_{y}u,\partial_{\omega}u)$ represents the gradient of $u:\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}\to\mathbb{R}$ with respect to the pose of the car. The target is the ball of radius $r=0.2$ centred at the origin. Since our framework applies to bounded domains, we set $R=5$, and define the domain
$\Omega:=\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)},$
where $\mathbb{A}_{r,R}$ denotes the two-dimensional annulus with inner radius $r$ and outer radius $R$
$\mathbb{A}_{r,R}:=\{(x,y)\in\mathbb{R}^{2}\,:\ r<\sqrt{x^{2}+y^{2}}<R\}.$ (5.21)
As for the boundary conditions, we set 00 on the inner boundary of the annulus and $R$ on the outer boundary. Hence, the resulting boundary value problem reads as
$\begin{cases}H(x,y,\omega,\nabla u)=0&(x,y,\omega)\in\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=0&(x,y,\omega)\in\{\sqrt{x^{2}+y^{2}}=r\}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=R&(x,y,\omega)\in\{\sqrt{x^{2}+y^{2}}=R\}\times\mathbb{T}_{[0,2\pi)}.\end{cases}$ (5.22)
We used our approach to approximate the viscosity solution of the above boundary value problem using a NN. Since the solution $u(x,y,\omega)$ is $2\pi$-periodic with respect to $\omega$, we considered a parameterised family of functions of the form
$\Phi(x,y,\omega;\theta)=\sum_{n=0}^{10}\phi(x,y;\theta_{n}^{(1)})\cos(n\omega)+\sum_{m=1}^{10}\phi(x,y;\theta_{m}^{(2)})\sin(m\omega),$
where the function $\phi(\cdot,\cdot;\theta):\mathbb{R}^{2}\to\mathbb{R}$ is a three-layer fully connected NN with 60 neurons in each hidden layer. We trained the NN for three rounds of Algorithm 2, using the following values for the parameters $\alpha$ and $\delta$:
$(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})=(2.5,\,2.5,\,2.5)\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2},\delta_{3})=(0.75,0.6,0.3).$
In each application of Algorithm 1, we run $3000$ iterations of stochastic gradient descent, using $N_{b_{1}}=100$ boundary points for the boundary data, and in each of the three rounds, we took $N=[800,1000,1200]$ as number of collocation points to approximate the finite-difference loss. The time it took to train the NN was $90$ seconds on a CPU.
Once trained, the NN can be used, as an approximation of the value function $u(x,y,\omega)$, to obtain a feedback control for the car. It is well-known that the optimal feedback control can be obtained in terms of the value function as
$\left(a^{\ast}(x,y,\omega),b^{\ast}(x,y,\omega)\right)\in\arg\min_{(a,b)\in[-1,1]^{2}}\left(\sigma a\partial_{x}u\cos\omega+\sigma a\partial_{y}u\sin\omega+\dfrac{1}{\rho}b\partial_{\omega}u-1\right),$
which yields
$a^{\ast}(x,y,\omega):=-\operatorname{sgn}\left(\partial_{x}u\cos\omega+\partial_{y}u\sin\omega\right),\quad\text{and}\quad b^{\ast}(x,y,\omega):=-\operatorname{sgn}\left(\partial_{w}u\right).$
In practice, we use the NN $\Phi(x,y,\omega;\theta)$ as an approximation of the value function $u(x,y,\omega;\theta)$, and the partial derivatives are replaced by finite-difference approximations as
$\partial_{x}\Phi(x,y,\omega;\theta)=\dfrac{\Phi(x+\delta,y,\omega;\theta)-\Phi(x,y,\omega;\theta)}{\delta},\quad\partial_{y}\Phi(x,y,\omega;\theta)=\dfrac{\Phi(x,y+\delta,\omega;\theta)-\Phi(x,y,\omega;\theta)}{\delta}$
$\partial_{\omega}\Phi(x,y,\omega;\theta)=\dfrac{\Phi(x,y,\omega+\delta;\theta)-\Phi(x,y,\omega;\theta)}{\delta}.$
where $\delta>0$ is the parameter used in the last step of the training schedule. Figure 8 shows some trajectories obtained by the trained Neural Network.
Figure 8: Five sample trajectories for the Reeds-Shepp’s car, obtained using Neural Network approximating the viscosity solution of (5.22). The underlying optimal control problem is the shortest path problem to the origin (represented by the blue dot), subject to the curvature-constrained dynamics in (5.20). The red star represents the initial position of the car, and the blue and green circles represent the turning circumference from the initial position and orientation of the car. The bottom right figure is the fifth trajectory, zoomed-in around the initial position. The car manoeuvres to find a shorter path to the target.
5.2.4 Pursuit-evasion game with Reeds-Shepp’s cars
As an example of applying our method to a Hamilton-Jacobi equation arising from differential game theory, we consider the pursuit-evasion game for two players that move according to Reeds and Shepp’s car model (5.20). Given the initial poses for the Evader and the Pursuer, represented by
$(x_{e},y_{e},\omega_{e})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)}\qquad\text{and}\qquad(x_{p},y_{p},\omega_{p})\in\mathbb{R}^{2}\times\mathbb{T}_{[0,2\pi)},$
we consider the game in which the Pursuer’s goal is to intercept the Evader in the shortest possible time, and the Evader’s goal is to delay the interception for as long as possible.
Since the position of the game only depends on the relative position between the players, we introduce the variables $X=x_{e}-x_{p}$ and $Y=y_{e}-y_{p}$. The vector $(X,Y)$ represents the vector joining the Evader with the Pursuer. The position of the game is therefore represented by $(X,Y,\omega_{e},\omega_{p})\in\mathbb{R}^{2}\times\mathbb{T}_{[1,2\pi)}^{2}$. Given the dynamics of the two cars in (5.20), we can write down the equation for the game as
$\begin{cases}\dot{X}(t)=\sigma_{e}a_{e}(t)\cos\omega_{e}(t)-\sigma_{p}a_{p}(t)\cos\omega_{p}(t)\\ \dot{Y}(t)=\sigma_{e}a_{e}(t)\sin\omega_{e}(t)-\sigma_{p}a_{p}(t)\sin\omega_{p}(t)\\ \dot{\omega}_{e}(t)=\dfrac{b_{e}(t)}{\rho_{e}},\qquad\dot{\omega}_{p}(t)=\dfrac{b_{p}(t)}{\rho_{p}}\\ X(0)=x_{e}-x_{p},\ Y(0)=y_{e}-y_{p},\ \omega_{e}(0)=\omega_{e},\ \omega_{p}(0)=\omega_{p},\end{cases}$ (5.23)
where $\sigma_{e}$ and $\sigma_{p}$ represent the maximum velocities for the Evader and the Pursuer, respectively, and $\rho_{e}$ and $\rho_{p}$ represent the inverse of their angular velocity (the turning radius for the evader and the pursuer is $\sigma_{e}\rho_{e}$ and $\sigma_{p}\rho_{p}$ respectively). The functions $a_{e}(\cdot),a_{p}(\cdot):[0,\infty)\to[-1,1]$ are the controls for the velocities of the players, and $b_{e}(\cdot),b_{p}(\cdot):[0,\infty)\to[-1,1]$ are the controls for the direction.
In this case, we define the domain of the game as
$\Omega:=\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}^{2},$
with $r=0.2$, $R=4$, where $\mathbb{A}_{r,R}$ is the two-dimensional annulus defined in (5.21), and $\mathbb{T}_{[0,2\pi)}^{2}$ is the two-dimensional flat torus with side-length $2\pi$. Recall that the position of the game is represented by the vector joining the players $(X,Y)=(x_{e}-x_{p},\,y_{e}-y_{p})\in\mathbb{R}^{2}$ and the orientation angle of the players $(\omega_{e},\omega_{p})\in\mathbb{T}_{[0,2\pi)}^{2}$. The target for the Pursuer is the inner boundary of the annulus $\mathbb{A}_{r,R}$, where we set zero boundary condition. The target for the Evader is the outer boundary, where we set the boundary condition equal to $10$.
Since the Pursuer minimises the output of the game and the Evader maximises it, we can derive the Hamiltonian associated with the game, which reads as
$H(X,Y,\omega_{e},\omega_{p},\nabla u):=\sigma_{p}\left|\partial_{X}u\cos\omega_{p}+\partial_{Y}u\sin\omega_{p}\right|+\dfrac{1}{\rho_{p}}\left|\partial_{\omega_{p}}u\right|-\sigma_{e}\left|\partial_{X}u\cos\omega_{e}+\partial_{Y}u\sin\omega_{e}\right|-\dfrac{1}{\rho_{e}}\left|\partial_{\omega_{e}}u\right|,$
where, for a function $u:\Omega\to\mathbb{R}$ defined on the game domain, $\nabla u=(\partial_{X}u,\partial_{Y}u,\partial_{\omega_{e}}u,\partial_{\omega_{p}}u)$ denotes the gradient of $u$ with respect to the position of the game. The Hamilton-Jacobi-Isaacs equation associated with the game reads as follows:
$\begin{cases}H(X,Y,\omega_{e},\omega_{p},\nabla u)=0&(X,Y,\omega_{e},\omega_{p})\in\mathbb{A}_{r,R}\times\mathbb{T}_{[0,2\pi)}^{2}\\ u(x,y,\omega)=0&(x,y,\omega)\in\{\sqrt{X^{2}+Y^{2}}=r\}\times\mathbb{T}_{[0,2\pi)}\\ u(x,y,\omega)=R&(x,y,\omega)\in\{\sqrt{X^{2}+Y^{2}}=R\}\times\mathbb{T}_{[0,2\pi)}.\end{cases}$ (5.24)
The viscosity solution of this boundary value problem characterises the value of the game and can be used to characterise the optimal feedback strategies for the players as
$a_{p}(X,Y,\omega_{e},\omega_{p})=-\operatorname{sgn}\left(\partial_{X}u\cos\omega_{p}+\partial_{Y}u\sin\omega_{p}\right),\qquad b_{p}(X,Y,\omega_{e},\omega_{p})=-\operatorname{sgn}\left(\partial_{\omega_{p}}u\right)$ (5.25)
for the Pursuer, and
$a_{e}(X,Y,\omega_{e},\omega_{p})=\operatorname{sgn}\left(\partial_{X}u\cos\omega_{e}+\partial_{Y}u\sin\omega_{e}\right),\qquad b_{e}(X,Y,\omega_{e},\omega_{p})=\operatorname{sgn}\left(\partial_{\omega_{e}}u\right)$ (5.26)
for the Evader.
For different values of the players’ velocities $(\sigma_{e},\rho_{e})$ and $(\sigma_{p},\rho_{p})$, we used our approach to approximate the viscosity solution of (5.24) using a NN. Similarly to the previous example, since the solution $u(X,Y,\omega_{e},\omega_{p})$ is $2\pi$-periodic with respect to $\omega_{e}$ and $\omega_{p}$, we considered a parameterised family of functions of the form
$\Phi(X,Y,\omega_{e},\omega_{p};\theta)$ $=$ $\sum_{n=0}^{4}\phi(X,Y;\theta_{n}^{(1)})\cos(n\omega_{e})+\sum_{m=1}^{4}\phi(X,Y;\theta_{m}^{(2)})\sin(m\omega_{e})$
$+\sum_{n=0}^{4}\phi(X,Y;\theta_{n}^{(3)})\cos(n\omega_{p})+\sum_{m=1}^{4}\phi(X,Y;\theta_{m}^{(4)})\sin(m\omega_{p})$
where the function $\phi(\cdot,\cdot;\theta):\mathbb{R}^{2}\to\mathbb{R}$ is a three-layer fully connected NN with 80 neurons in each hidden layer. We trained the NN for three rounds of Algorithm 2, using the following values for the parameters $\alpha$ and $\delta$:
$(\alpha_{0},\alpha_{1},\alpha_{2})=(2.5,2,1.5)\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2})=(0.7,0.5,0.3).$
In each application of Algorithm 1, we run $3000$ iterations of stochastic gradient descent, using $N_{b_{1}}=100$ points for the boundary data, and $N=800$ collocation points to approximate the least-squares finite-difference loss. The running time to train each NN was 120 seconds on a CPU.
In Figure 9, each sample trajectory of the game is computed using the feedback control according to (5.25) and (5.26), with the trained NN approximating the value function $u$. As in the previous example, the partial derivatives are approximated by finite-differences. In the first example, the Pursuer moves faster than the Evader ($\sigma_{p}>\sigma_{e}$), but the Evader has a higher angular velocity ($1/\rho_{e}>1/\rho_{p}$). The plot at the right represents the evolution of the distance between the players throughout the game. We see an oscillatory behaviour due to the turns that the Evader performs for not being intercepted. In some cases we see that, near the maxima of the curve, the distance between the players increases abruptly before it starts decreasing. This is due to the fact that the Pursuer has the possibility to back up in order to change the direction faster (similarly to the last example in Figure 8). In the second example, we see a trajectory in which both players have the same angular velocity, but the pursuer moves faster. Hence, the turning radius of the Evader is smaller ($\sigma_{e}\rho_{e}<\sigma_{p}\rho_{p}$). In this case, the game enters in a loop. Finally, in the third example, we see a trajectory in which the pursuer moves faster, and both players have the same turning radius ($\sigma_{e}\rho_{e}=\sigma_{p}\rho_{p}$). In this case, the Pursuer has a clear advantage over the Evader, and interception occurs quickly. See https://github.com/carlosesteveyague/HamiltonJacobi_LeastSquares_LxF_NNs for other animated representations of trajectories related to this Pursuit-Evasion game.
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=1.2$ Refer to caption Refer to caption
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=1$ Refer to caption Refer to caption
$\sigma_{e}=0.8$ $\rho_{e}=1$ $\sigma_{p}=1$ $\rho_{p}=0.8$ Refer to caption Refer to caption
Players’ velocities Game trajectory Time versus distance between players
Figure 9: Three trajectories of the Pursuit Evasion game with Reeds-Shepp’s cars discussed in Section 5.2.4. We consider different velocity parameters for the players (left) and different initial positions shown by the stars.
6 Qualitative analysis of the algorithm
In this section, we analyse various aspects of our proposed algorithm using judiciously designed numerical experiments. In particular, we want to show how costly it is to train a NN to approximate the global minimizer of a functional of the form
[…]
where $\widehat{H}_{\alpha}(x,p^{+},p^{-})$ is the Lax-Friedrichs numerical Hamiltonian defined in (4.1), $\Gamma\subset\Omega$ is a subset of the domain where the viscosity solution $u^{\ast}$ is assumed to be known, and $\rho,\mu$ and $\nu$ are probability distributions on $\Omega$, $\partial\Omega$ and $\Gamma$ respectively. In all the experiments in this section, we consider the weighting parameters $\gamma_{1}=\gamma_{2}=1$.
  1. i.
    In Section 6.1, the experiments will show that, to prevent deterioration of generalization errors, the number of collocation points should increase as $\delta$ decreases.
  2. ii.
    Also in Section 6.1, we will see that resampling the collocation points at each SGD iteration drastically increases the performance of the trained NNs.
  3. iii.
    In Section 6.2, we study the choice of the distribution for the collocation points $\rho$ in (6.1). We will see that, especially in high-dimensional scenarios, a suitable choice of $\rho$ may greatly impact the accuracy of the obtained numerical solution.
  4. iv.
    In Section 6.3, we will show that adding supervised data in the interior of $\Omega$ allows for smaller values of $\delta$ and $\alpha$ while keeping the guarantee of convergence to the viscosity solution. This was anticipated in section 4.3.
For simplicity, we will always consider the Eikonal equation in simple domains.
6.1 Data efficiency and sampling strategy
As argued in Section 3.1 and Section 5.1, minimising the functional $\mathcal{J}(u)$ with smaller values of $\delta$ may demand more collocation points in the domain $\Omega$ due to a smaller range of influence per point. While we cannot prove it rigorously, we illustrate this phenomenon through simple numerical experiments. In our tests, we consider three different scenarios in terms of training data (collocation points + boundary points), and compare the generalisation power of the NN after beging trained with decreasing values of $\delta$ and a fixed parameter $\alpha$.
For each value of $\delta$, the NN is trained until the training error is close to zero. When the training data is insufficient, the finite-difference residual $\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}$ is not necessarily small at points away from the training data.
We consider the Eikonal equation (5.11) in a $d$-dimensional ball of radius $3$, i.e. $\Omega=B(0,3)\in\mathbb{R}^{d}$, first with $d=2$ and then with $d=5$. The viscosity solution to this boundary value problem is given by
$u(x)=3-\|x\|,\qquad\text{for}\ x\in\Omega,$ (6.2)
which will be used as the ground truth solution to evaluate the accuracy of the numerical solution.
To study the data efficiency during the training process, we trained three-layer fully connected MLPs (with $20$ neurons in each hidden layer for $d=2$, and with $30$ neurons for $d=5$) by applying Algorithm 2 to minimise the loss functional $\mathcal{J}(\cdot)$ in (6.1), with $\Gamma=\emptyset$ (i.e. we consider supervised data only on the boundary). The loss functional $\mathcal{J}(\cdot)$ is approximated, using Monte Carlo approximation, as
$\mathcal{J}\left(\Phi(\cdot;\theta)\right)\approx\tilde{\mathcal{J}}\left(\Phi(\cdot;\theta)\right):=\dfrac{1}{N_{0}}\sum_{x\in X}\left[\widehat{H}_{\alpha}(D_{\delta}^{+}\Phi(x;\theta),D_{\delta}^{-}\Phi(x;\theta))\right]^{2}+\dfrac{1}{N_{b}}\sum_{x\in X_{b}}\left(\Phi(x;\theta)\right)^{2}$
where the collocation points $X=\{x_{j}\}_{j=1}^{N_{0}}$ are uniformly sampled from $\Omega$ and the boundary points $X_{b}=\{x_{j}\}_{j=1}^{N_{b}}$ are uniformly sampled from $\partial\Omega$. As for the training schedule, we considered $5$ training steps ($M=4$ in Algorithm 2) in which the parameters $\alpha$ and $\delta$ in the numerical Hamiltonian are set as follows:
$\alpha_{i}=2\quad\forall i=0,\ldots,4\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2},\delta_{3},\delta_{4})=(0.7,0.5,0.2,0.1,0.05).$ (6.3)
In the first two scenarios, the collocation points $X$ and $X_{b}$ are sampled once at the beginning (we shall refer to them as training data) and then re-used throughout the training of the NN. In particular, for $d=2$, we considered $N_{0}=80$ and $N_{b}=40$ in the first case, and $N_{0}=160$ and $N_{b}=80$ in the second case. For $d=5$, we considered $N_{0}=200$ and $N_{b}=40$ in the first case, and $N_{0}=400$ and $N_{b}=80$ in the second case. In both cases, we trained the NN for $500$ epochs using SGD with a batch size of $20$ for $d=2$ and $50$ for $d=5$. The training data is randomly shuffled and partitioned into batches in each epoch. Then, a gradient step is performed using the data in each batch. One epoch refers to when all the training data has been used once.
In the third case, the training data $X$ and $X_{b}$ are resampled at every SGD iteration, as described in Algorithm 1. In particular, for $d=2$, we carried $J=2500$ SGD iterations [^2] Note that $500$ epochs of SGD with a dataset of size $120$ splitted in batches of $20$, as in the first case, corresponds to $3000$ iterations. with $N_{0}=20$ and $N_{b}=8$. For $d=5$, we carried $J=2500$ SGD iterations with $N_{0}=300$ and $N_{b}=50$. To put the three cases into the same setting, the latter case can be seen as a single training epoch with a dataset of size $(20+8)\times 2500=70000$ for $d=2$, and $(300+50)\times 2500=875000$ for $d=5$.
We test the trained NNs on a uniform grid in $\Omega$ with $\delta=0.01$ for $d=2$ and $\delta=0.3$ for $d=5$. We shall refer to the grid nodes as the test data, and denote it by $X_{t}\subset\Omega$. It contains around $N_{t}\approx 2.8\times 10^{5}$ points for $d=2$ and around $N_{t}\approx 5.3\times 10^{5}$ for $d=5$.
Figure 10 shows the heat maps of the finite-difference residual $x\mapsto\left[\widehat{H}_{\alpha}(D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}$ after training the NN with different values of $\delta$ for the two dimensional case. The top and middle rows of figures correspond to the two first scenarios described above, where the collocation points (represented by red dots) are sampled at the beginning and then re-used throughout the training. The areas not covered by the training data appear brighter since the finite-difference residual is larger. This effect is mitigated for larger values of $\delta$, showing improved data efficiency in this case. In the bottom row, we show the heat maps from using the resampling strategy described above. The training data is not represented in this case due to visualisation purposes. We see that the finite-difference residual is smaller, also for small values of $\delta$.
(a) $\delta=0.7$
(b) $\delta=0.5$
(c) $\delta=0.2$
(d) $\delta=0.1$
(e) $\delta=0.05$
Figure 10: Heat map of the finite-difference residual $x\mapsto\left[\widehat{H}_{\alpha}(D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}$. We show one example for each training scenario described in section 6.1: $N_{0}=80$ collocation points (top), $N_{0}=160$ collocation points (middle) and using the resampling strategy (bottom). The collocation points for the first two cases are represented by red dots. The training data for the resampling strategy is not represented as it covers most of the domain.
Figures 11 and 12 show a summary of ten independent experiments using each of the three training data scenarios described above, for the case $d=2$. Figure 11 shows the evolution of the finite-difference residual, evaluated on the test data. The left plot represents the average over the test data, and the right plot shows the maximum value:
^$\widehat{\mathcal{R}}(u)=\dfrac{1}{N_{t}}\sum_{x\in X_{t}}\left[\widehat{H}(D_{\delta}^{+}u(x),D_{\delta}^{-}(x))\right]^{2}\qquad\text{and}\qquad\widehat{\mathcal{R}}_{\infty}(u)=\max_{x\in X_{t}}\left[\widehat{H}(D_{\delta}^{+}u(x),D_{\delta}^{-}(x))\right]^{2}.$ (6.4)
In Figure 12 we see the evolution of the mean square error (5.9) and the $L^{\infty}$-error (5.10), with respect to the ground truth solution (6.2), as the parameter $\delta$ is decreased following the training schedule in (6.3). The solid line shows the averaged value over the 10 experiments, and the shadowed area shows the interval within the standard deviation. Figures 13 and 14 show the same summary of ten independent experiments using each of the three training data scenarios described above for the case $d=5$.
In Figures 12 and 14, we observe that the error with respect to the ground truth solution decreases as $\delta$ is reduced. On the contrary, the finite-difference loss evaluated on test data does not decrease. It actually increases if we look at the maximum over the test data (right plots in Figures 11 and 13). Although these two statements might seem to be in contradiction, we stress that, when $\delta$ is smaller, the improved accuracy in terms of error with respect to the ground truth solution is due to less numerical diffusion in the discretised Hamiltonian. In conclusion, reducing $\delta$ has two opposite effects on the error with respect to the ground truth solution: on one hand, the reduced numerical diffusion enhances the accuracy, and on the other hand, the smaller range of influence per collocation point makes training less data efficient. We also observe that the resampling strategy gives better results, and this improvement is more noticeable for smaller values of $\delta$. In Figure 14, we see that when the dimension is higher, the improvement of the resampling strategy is more substantial. We also note that, when using resampling, the standard deviation over the ten experiments is smaller, indicating that this strategy makes the method more robust.
Figure 11: (Residual versus $\delta$ for the eikonal equation in a disc) The plots show the evolution of the finite-difference residual, evaluated on the test data $X_{t}$, as the parameter $\delta$ is reduced according to (6.3), for the problem (5.11) with $\Omega=B(0,3)$ and $d=2$. We show the average over the domain (left) ^$\widehat{\mathcal{R}}(u)$ and its maximum (right) ^$\widehat{\mathcal{R}}_{\infty}(u)$, as defined in (6.4). The solid line represents the mean value computed over ten independent experiments, and the shadowed area represents the interval within one standard deviation.
Figure 12: (Error versus $\delta$ for the eikonal equation in a disc) Evolution of the error with respect to the ground truth solution (6.2) with $d=2$, evaluated on the test data $X_{t}$, as the parameter $\delta$ is reduced according to (6.3). We show the mean square error (left) and the $L^{\infty}$-error (right), as defined in (5.9) and (5.10) respectively. The solid line represents the mean value computed over ten independent experiments, and the shadowed area represents the interval within one standard deviation.
Figure 13: (Residual versus $\delta$ for the eikonal equation in a ball in $\mathbb{R}^{5}$) Evolution of the finite-difference residual, evaluated on the test data $X_{t}$, as the parameter $\delta$ is decreased according to (6.3), for the problem (5.11) with $\Omega=B(0,3)$ and $d=5$. We show the average over the domain (left) ^$\widehat{\mathcal{R}}(u)$ and its maximum (right) ^$\widehat{\mathcal{R}}_{\infty}(u)$, as defined in (6.4). The solid line represents the mean value computed over ten independent experiments, and the shadowed area represents the interval within one standard deviation.
Figure 14: (Error versus $\delta$ for the eikonal equation in a ball in $\mathbb{R}^{5}$) Evolution of the error with respect to the ground truth solution (6.2) with $d=5$, evaluated on the test data $X_{t}$, as the parameter $\delta$ is decreased according to (6.3). We show the mean square error (left) and the $L^{\infty}$-error (right), as defined in (5.9) and (5.10) respectively. The solid line represents the mean value computed over ten independent experiments, and the shadowed area represents the interval within one standard deviation.
6.2 Data distribution in high-dimensional scenarios
When the dimension is high, the sampling distribution used during the iterations of Algorithm 1 dramatically impacts the numerical solution. Let the domain $\Omega$ be a $d$-dimensional ball. Sample points of the uniform distribution over $\Omega$ with respect to the Lebesgue measure, will be closer to the boundary with higher probability. This is simply because the outer rings of the ball occupy a larger volume (of the order $\sim r^{d-1}$). This “concentration” causes the accuracy near the centre of the ball to be relatively poor compared to the accuracy of the solution near the boundary. This issue can be alleviated by using a different distribution for the collocation points. In the case when $\Omega$ is a $d$-dimensional ball of radius $R$, one possibility is to use a radially uniform sampling, i.e.
$X=rv,\quad\text{where}\ r\sim\operatorname{Unif}(0,R)\ \text{and}\ v\sim\operatorname{Unif}(\mathbb{S}^{d-1}),$ (6.5)
where $\mathbb{S}^{d-1}$ denotes the $(d-1)-$dimensional unit sphere.
In Table 5, we see the summary of the numerical experiments for the Eikonal equation in a $10$-dimensional ball of radius $6$. Again, for each experiment we carried out $4$ rounds of Algorithm 2, with parameters $\alpha$ and $\delta$ following the schedule
$(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})=(2.5,2,1,0)\quad\text{and}\quad(\delta_{0},\delta_{1},\delta_{2},\delta_{3})=(0.7,0.3,0.1,0.01).$ (6.6)
Each experiment is repeated 10 times with different random seeds. The error values shown in the table represent the average and one standard deviation. The evolution of the MSE and the $L^{\infty}$-error throughout the iterations of Algorithm 2 is shown in Figure 15. We observe that the accuracy of the numerical solution in terms of MSE is similar when using a uniform distribution or a radially uniform distribution to sample the collocation points in $\Omega$. See Table 5 and Figure 15 (left). However, in terms of $L^{\infty}$-error, the accuracy is remarkably better when using the radially uniform sampling defined in (6.5). See Table 5 and Figure 15 (right). This is because, when using uniform sampling, most of the points are close to the boundary, and the accuracy near the centre of the ball is poorer. See Figure 16 for a representation of the numerical solution using different sampling distributions for the Eikonal equation in the $10$-dimensional ball $\Omega=B_{10}(0,6)$. In the plot on the left, we used a uniform sampling for the collocation points during the optimisation iterations. In contrast, in the plot at the right, we used the radially uniform sampling described in (6.5). In both cases, the plots represent the NN evaluated on a uniform grid of the two-dimensional cross-section $B_{2}(0,6)\times\{0\}^{8}$.
Table 5: Summary of the numerical experiments for the Eikonal equation in $\Omega=B_{10}(0,6)$, i.e. the $10$-dimensional ball of radius 6. In each experiment, we applied 4 iterations of Algorithm 2 with decreasing values of $\alpha$ and $\delta$, following the schedule (6.6). See Figure 15 for the evolution of the MSE and the $L^{\infty}$-error after each iteration. In all the experiments, the batch size in the implementation of SGD in (5.3) is taken as $N_{0}=200$ and $N_{b_{1}}=80$. We see that the use of a radially uniform distribution (6.5) for the collocation points (experiments 4,5 and 6) improves the accuracy in terms of $L^{\infty}$-error.
Exp. Dim. Distr. Architecture Iterations (×$\times 10^{3}$) MSE $L^{\infty}$-error Runtime (s)
1 10 Unif. [40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.02 ± 0.003 2.355 ± 0.202 55
2 10 Unif. [40, 40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.015 ± 0.004 2.773 ± 0.077 74
3 10 Unif. [40, 40, 40] $\{1.5,2,2.5,3\}$ 0.017 ± 0.008 2.646 ± 0.148 115
4 10 Rad. [40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.026 ± 0.004 0.768 ± 0.053 55
5 10 Rad. [40, 40, 40] $\{1.5,1.5,1.5,1.5\}$ 0.031 ± 0.011 0.795 ± 0.064 77
6 10 Rad. [40, 40, 40] $\{1.5,2,2.5,3\}$ 0.017 ± 0.003 0.619 ± 0.034 121
Figure 15: (Error versus iteration index for the eikonal equation in a ball in $\mathbb{R}^{10}$)Evolution of the MSE (left) and the $L^{\infty}$-error (right) throughout the iterations of Algorithm 2, for the experiments 3 and 6 reported in Table 5. These correspond to the Eikonal equation in a 10-dimensional ball.
Figure 16: Eikonal equation in a ball in $\mathbb{R}^{10}$) The numerical solution for the Eikonal equation in $\Omega=B_{10}(0,6)\subset\mathbb{R}^{10}$, evaluated on a uniform grid of the two-dimensional cross-section $B_{2}(0,6)\times\{0\}^{8}$. In the figure at the left, a uniform distribution over $\Omega$ has been used in Algorithm 1, whereas in the figure at the right, we used the radially uniform sampling defined in (6.5).
6.3 Supervised data in domain interior
Theorem 1 provides a sufficient condition for the convergence of Algorithm 2 to the solution of the difference equation. However, when (4.10) is not satisfied, we cannot a priori guarantee the convergence of the method to the searched solution.
To better understand the behaviour of the proposed method under this uncertain regime, we carry out experiments to estimate the success rate of convergence to the right minimizer for the Eikonal Equation (see Section 5.2.1 for the description of the problem and settings) on different domains. For these tests, we fix $\alpha=2$ and consider decreasing values of $\delta$, namely
$(\delta_{0},\delta_{1},\delta_{2})=(0.7,0.2,0.01).$
The value of $\delta$ is kept constant in each experiment.
The method is monotone for $\alpha=2$, so to check convergence to the correct minimizer for the Eikonal equation, we only need to consider the sign of the solution. Therefore, we consider solutions correct if they have the right sign after training for a set number of epochs or if they can switch to the correct sign after an additional training session. The success rate (for a fixed value of the parameters $\alpha$ and $\delta$) is then obtained as the ratio between the number of correct solutions and the total number of experiments. Table 6 shows that the success rate decreases as $\delta$ as we expected.
Table 6: Success rate of Convergence to the right minimizer for different domains and values of $\delta$
Domain $\alpha$ $\delta$ Success rate ($\Gamma=\emptyset$) Success rate (localized $\Gamma$) Success rate (uniform $\Gamma$)
Square $2$ $0.7$ $1.00$ $1.00$ $1.00$
$0.2$ $0.83\pm 0.08$ $1.00-0.03$ $1.00-0.03$
$0.01$ $0.48\pm 0.09$ $0.96\pm 0.03$ $1.00-0.03$
Ball $2$ $0.7$ $1.00$ $1.00$ $1.00$
$0.2$ $0.91\pm 0.05$ $1.00-0.03$ $1.00-0.03$
$0.01$ $0.53\pm 0.09$ $0.96\pm 0.03$ $1.00-0.03$
Annulus $2$ $0.7$ $1.00$ $1.00$ $1.00$
$0.2$ $0.43\pm 0.09$ $0.96\pm 0.03$ $1.00-0.03$
$0.01$ $0.02\pm 0.02$ $0.31\pm 0.09$ $0.75\pm 0.08$
Now we investigate the effect of having additonal supervised data, $\Gamma$ in the interior of $\Omega$. We assume that
$\text{dist}(\Gamma,\partial\Omega)>C$ (6.7)
for some $C>0$ sufficiently large. The additional supervised data change the spectrum of the discrete Laplacian and could potentially convexify the least square fucntional; see (4.10).
Let us then suppose (6.7) to hold for a suitable constant $C$ depending on $\delta$. To verify the beneficial effects of adding supervised data, we run experiments first setting $\Gamma$ to be composed by $10$ randomly selected points sampled in a small internal region of $\Omega$ and then by $10$ randomly chosen points on the all domain $\Omega$.
We first run experiments where $\Gamma$ contains $10$ randomly selected points sampled from a small subdomain of $\Omega$. The success rates are reported in the fifth column of Table 6. We see that the added knowledge on $\Gamma$ enable a remarkable improvement in the success rate. Nevertheless, it is worth underlining that, as before, the shape of the domain seems to affect the results so that the positive influence of $\Gamma$ is mitigated, for instance, in the annulus case.
Next, run experiments where $\Gamma$ contains $10$ points sampled unformly from the interior of $\Omega$. The success rates are reported in the last column of Table 6). We can appreciate an improvement in the success rate even on the annulus, confirming the dependency of the choice of $\Gamma$ on the shape of the domain $\Omega$.
These experiments suggest that supervised data points in the domain interior can improve the overall data efficiency and generalization power of the trained NNs. In practice, one may use other pathwise Hamilton-Jacobi solvers, such as those based on Hopf-Lax formula [11] (when applicable), to general supversied data points.
7 Proofs of the main results
7.1 Notation and preliminaries
In this section, we only consider the $d$-dimensional cube $\Omega:=(0,1)^{d}$. For any $N\in\mathbb{N}$, fix $\delta=1/N$, and consider the uniform Cartesian grids of the domains $\overline{\Omega}$ and $\Omega$, given by
$\overline{\Omega}_{\delta}:=\delta\mathbb{Z}^{d}\cap\overline{\Omega}=\left\{x_{\beta}:=\delta\beta\,:\ \beta\in\{0,1,\ldots,N\}^{d}\right\}$ (7.1)
$\Omega_{\delta}:=\delta\mathbb{Z}^{d}\cap\Omega=\left\{x_{\beta}:=\delta\beta\,:\ \beta\in\{1,\ldots,N-1\}^{d}\right\}$ (7.2)
From now on, we will use the multi-index notation $\beta:=(\beta_{1},\ldots,\beta_{d})\in\{0,1,\ldots,N\}^{d}$. Furthermore, we denote by $\{e_{i}\}_{i=1}^{d}$ the canonical basis of $\mathbb{R}^{d}$, so that $x_{\beta+e_{i}}=x_{\beta}+\delta e_{i}$ and $x_{\beta-e_{i}}=x_{\beta}-\delta e_{i}$. We also introduce the index domains
$\overline{\mathfrak{B}}:=\{0,\ldots,N\}^{d}\quad\text{and}\quad\mathfrak{B}:=\{1,\ldots,N-1\}^{d}.$ (7.3)
Then, we can write $\overline{\Omega}_{\delta}=\{x_{\beta}\,:\ \beta\in\overline{\mathfrak{B}}\}$ and $\Omega_{\delta}=\{x_{\beta}\,:\ \beta\in\mathfrak{B}\}$.
For a given a function $u\in C(\overline{\Omega})$, let us denote by $U$ the grid function associated to $u$ on the grid $\overline{\Omega}_{\delta}$, i.e.
$U:=\left\{U_{\beta}=u(x_{\beta})\,:\ \beta\in\overline{\mathfrak{B}}\right\}.$
Using the above notation, we can write the functional ^$\widehat{\mathcal{R}}(u)$ as
^$\widehat{\mathcal{R}}(u)$ $=$ $\delta^{d}\sum_{x\in\Omega_{\delta}}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2}$ (7.4)
$=$ $\delta^{d}\sum_{x\in\Omega_{\delta}}\left[H\left(x,\dfrac{D_{\delta}^{+}u(x)+D_{\delta}^{-}u(x)}{2}\right)-\alpha\sum_{i=1}^{d}\left(\dfrac{D_{\delta}^{+}u(x)-D_{\delta}^{-}u(x)}{2}\right)_{i}\right]^{2}$
$=$ $\delta^{d}\sum_{\beta\in\mathfrak{B}}\left[H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}\right]^{2}=:F(U),$
where, for any $\beta\in\mathfrak{B}$, $D_{\delta}U_{\beta}$ denotes the centred finite-difference approximation of the gradient, i.e.
$D_{\delta}U_{\beta}=\left(\dfrac{U_{\beta+e_{1}}-U_{\beta-e_{1}}}{2\delta},\ldots,\dfrac{U_{\beta+e_{d}}-U_{\beta-e_{d}}}{2\delta}\right)\in\mathbb{R}^{d},$ (7.5)
and, for any $i\in\{1,\ldots,d\},$ $D_{\delta,i}^{2}U_{\beta}$ is given by
$D_{\delta,i}^{2}U_{\beta}=\dfrac{U_{\beta+e_{i}}+U_{\beta-e_{i}}-2U_{\beta}}{2\delta}.$ (7.6)
Note that the notation $D_{\delta,i}^{2}$ does not correspond to the application of $D_{\delta,i}$ twice.
In view of the assumption (1.3), the functional $U\mapsto F(U)$ defined in (7.4) is differentiable in the space of grid functions on $\overline{\Omega}_{\delta}$. The main step in the proof of Theorem 1 is to verify that if $\nabla F(U)=0$, then
$H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}=0,\qquad\forall\beta\in\mathfrak{B}.$
In the next lemma, we compute the partial derivatives of $F(U)$ with respect to $U_{\beta}$, for all $\beta\in\mathfrak{B}$.
Lemma 1.
Let $H$ satisfy (1.3), $\alpha>0$ and, for any $N\in\mathbb{N}$ and $\delta=1/N$, consider the grid $\overline{\Omega}_{\delta}$ as in (7.1), and the index domains $\overline{\mathfrak{B}}$ and $\mathfrak{B}$ as in (7.3). Let $F(U)$ be defined as in (7.4) for any grid function $U$ on $\overline{\Omega}_{\delta}$. Then we have
$\partial_{U_{\beta}}F(U)=-\delta^{d-1}\left(\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right)+\alpha\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right)\right),\qquad\forall\beta\in\mathfrak{B},$
where
$W_{\beta}:=\begin{cases}H\left(x_{\beta},D_{\delta}U_{\beta}\right)-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta},&\forall\beta\in\mathfrak{B},\\ 0&\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B},\end{cases}$ (7.7)
and for all $i\in\{1,\ldots,d\}$,
$V_{\beta}^{(i)}:=\begin{cases}\partial_{p_{i}}H\left(x_{\beta},D_{\delta}U_{\beta}\right),&\forall\beta\in\mathfrak{B}\\ 0&\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}.\end{cases}$ (7.8)
Proof.
Note that, in view of the definition of $F(U)$ in (7.4) and the definition of $W=\{W_{\beta}\}_{\beta\in\overline{\mathfrak{B}}}$ in (7.7), we have
$F(U)=\sum_{\beta\in\overline{\mathfrak{B}}}W_{\beta}^{2}.$
For any $\beta\in\mathfrak{B}$, the term $U_{\beta}$ appears only in the terms associated to the index $\beta$ and the indexes $\beta\pm e_{i}$ for all $i\in\{1,\ldots,d\}$. Moreover, we note that, in view of (7.5), the term $H(x_{\beta},D_{\delta}U_{\beta})$ is independent of $U_{\beta}$, and for any $i$, the terms $H(x_{\beta\pm e_{i}},D_{\delta}U_{\beta\pm e_{i}})$ depend on $U_{\beta}$ only through the $i$-th component of the vector $D_{\delta}U_{\beta\pm e_{i}}$.
Hence, we can write $\partial_{U_{\beta}}F(U)$ as
$\partial_{U_{\beta}}F(U)$ $=$ $-2\delta^{d}\alpha W_{\beta}\sum_{i=1}^{d}\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta}\right)$
$+2\delta^{d}\sum_{i=1}^{d}W_{\beta+e_{i}}\Big{(}\underbrace{\partial_{p_{i}}H\left(x_{\beta+e_{i}},D_{\delta}U_{\beta+e_{i}}\right)}_{V_{\beta+e_{i}}^{(i)}}\partial_{U_{\beta}}\left(D_{\delta}U_{\beta+e_{i}}\right)_{i}-\alpha\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta+e_{i}}\right)\Big{)}$
$+2\delta^{d}\sum_{i=1}^{d}W_{\beta-e_{i}}\Big{(}\underbrace{\partial_{p_{i}}H\left(x_{\beta-e_{i}},D_{\delta}U_{\beta-e_{i}}\right)}_{V_{\beta-e_{i}}^{(i)}}\partial_{U_{\beta}}\left(D_{\delta}U_{\beta-e_{i}}\right)_{i}-\alpha\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta-e_{i}}\right)\Big{)}.$
Using (7.6), we can compute
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+e_{i}}+U_{\beta-e_{i}}-2U_{\beta}}{2\delta}\right)=-\dfrac{1}{\delta}.$
Next, for any $i\in\{1,\ldots,d\}$ such that $\beta+e_{i}\in\mathfrak{B}$ (note that the terms such that $\beta+e_{i}\not\in\mathfrak{B}$ do not matter since $W_{\beta+e_{i}}$ would be zero), we can use (7.5) and (7.6) to compute
$\partial_{U_{\beta}}\left(D_{\delta}U_{\beta+e_{i}}\right)_{i}=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+2e_{i}}-U_{\beta}}{2\delta}\right)=-\dfrac{1}{2\delta},$
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta+e_{i}}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta+2e_{i}}+U_{\beta}-2U_{\beta+e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta}.$
Similarly, for any $i\in\{1,\ldots,d\}$ such that $\beta-e_{i}\in\mathfrak{B}$, we can compute
$\partial_{U_{\beta}}\left(D_{\delta}U_{\beta-e_{i}}\right)_{i}=\partial_{U_{\beta}}\left(\dfrac{U_{\beta}-U_{\beta-2e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta},$
$\partial_{U_{\beta}}\left(D_{\delta,i}^{2}U_{\beta-e_{i}}\right)=\partial_{U_{\beta}}\left(\dfrac{U_{\beta}+U_{\beta-2e_{i}}-2U_{\beta-e_{i}}}{2\delta}\right)=\dfrac{1}{2\delta}.$
By plugging all these values in (7.1), we obtain
$\partial_{U_{\beta}}F(U)$ $=$ $\delta^{d-1}\sum_{i=1}^{d}2\alpha W_{\beta}+\delta^{d-1}\sum_{i=1}^{d}\left(-V_{\beta+e_{1}}^{(i)}-\alpha\right)W_{\beta+e_{i}}+\delta^{d-1}\sum_{i=1}^{d}\left(V_{\beta-e_{i}}^{(i)}-\alpha\right)W_{\beta-e_{i}},$
and the conclusion follows by regrouping the terms in the above sum. ∎
7.2 Proof of Theorem 1 and Corollary 1
We are now in a position to carry out the proof of Theorem 1, which consists of two steps:
  1. i.
    First we prove that if $u\in C(\overline{\Omega})$ is a critical point of $\widehat{\mathcal{R}}(\cdot)$, then the associated grid function $U$ on $\overline{\Omega}_{\delta}$ satisfies $\nabla F(U)=0$, where $F(U)$ is given by (7.4).
  2. ii.
    Then, we use Lemma 1 to prove that if $u$ is $L$-Lipschitz, and $\alpha$ and $\delta$ satisfy (4.8), then $\nabla F(U)=0$ implies that $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0$ for all $x\in\Omega_{\delta}$.
Proof.
Step 1: Let $u\in C(\overline{\Omega})$. If $u$ is a critical point of $\widehat{\mathcal{R}}(\cdot)$, then it must hold that
$\lim_{\varepsilon\to 0}\dfrac{\widehat{\mathcal{R}}(u+\varepsilon\phi)-\widehat{\mathcal{R}}(u)}{\varepsilon}=0,\qquad\text{for any function $\phi\in C(\overline{\Omega})$.}$ (7.10)
Now, for any grid function $\Phi$ on $\overline{\Omega}_{\delta}$, let $\phi\in C(\overline{\Omega})$ be a function such that $\phi(x_{\beta})=\Phi_{\beta}$ for all $\beta\in\overline{\mathfrak{B}}$. In other words, consider a continuous function $\phi$ such that the associated grid function is $\Phi$. In view of (7.10), and using the function $F(\cdot)$ as defined in (7.4), we deduce that
$0=\lim_{\varepsilon\to 0}\dfrac{\widehat{\mathcal{R}}(u+\varepsilon\phi)-\widehat{\mathcal{R}}(u)}{\varepsilon}=\lim_{\varepsilon\to 0}\dfrac{F(U+\varepsilon\Phi)-F(U)}{\varepsilon}=\nabla F(U)\cdot\Phi.$
Since this holds for any grid function $\Phi$, we deduce that $\nabla F(U)=0$.
Step 2: In view of Lemma 1, the equation $\nabla F(U)=0$ can be written as
$-\delta^{d-1}(A_{N}(V)+\alpha\Delta_{N})W=0,$
where $W$ and $V=(V^{(1)},\ldots,V^{(d)})$ are the grid functions on $\Omega_{\delta}$ defined in the statement of Lemma 1, and $A_{N}(V)$ and $\Delta_{N}$ are linear operators on the space of grid functions on $\Omega_{\delta}$, defined below, in Lemmas 2 and 3 respectively. The conclusion of the proof follows after proving that the linear operator $A_{N}(V)+\alpha\Delta_{N}$ is invertible, so the optimality condition $\nabla F(U)=0$ implies that $W=0$, and hence,
$\widehat{H}_{\alpha}(x_{\beta},D_{\delta}^{+}u(x_{\beta}),D_{\delta}^{-}u(x_{\beta}))=H(x_{\beta},D_{\delta}U_{\beta})-\alpha\sum_{i=1}^{d}D_{\delta,i}^{2}U_{\beta}=W_{\beta}=0,\quad\forall\beta\in\mathfrak{B}.$
Recall that, by the Definition (7.2), $\Omega_{\delta}=\{x_{\beta}\,:\ \beta\in\mathfrak{B}\}$. To prove that the operator $A_{N}(V)+\alpha\Delta_{N}$ is invertible, it is enough to prove that
$\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)=\min_{W}\dfrac{\|\left(A_{N}(V)+\alpha\Delta_{N}\right)W\|_{2}}{\|W\|_{2}}>0.$
Note that $\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)$ is a lower bound for the modulus of the eigenvalues of $A_{N}(V)+\alpha\Delta_{N}$. Using Lemmas 2 and 3 below, we have that
$\underline{\sigma}\left(A_{N}(V)+\alpha\Delta_{N}\right)$$\geq$ $\min_{W}\dfrac{\alpha\|\Delta_{N}W\|_{2}-\|A_{N}(V)W\|_{2}}{\|W\|_{2}}$
$\geq$ $\min_{W}\dfrac{\alpha\|\Delta_{N}W\|_{2}}{\|W\|_{2}}-\max_{W}\dfrac{\|A_{N}(V)W\|_{2}}{\|W\|_{2}}$
$\geq$ $4\alpha d\sin^{2}\left(\dfrac{\pi}{2}\delta\right)-2dC_{H}(L).$
Hence, we deduce that $\underline{\sigma}(A_{N}(V)+\alpha\Delta_{N})>0$ provided $2\alpha\sin^{2}\left(\frac{\pi}{2}\delta\right)>C_{H}(L)$, where $C_{H}(L)$ is defined in the statement of Lemma 2. ∎
Next, we prove the two following lemmas, which are used, at the end of the proof of Theorem 1, to bound the spectrum of the operators $A_{N}(V)$ and $\Delta_{N}$. Although these results might be known to the expert reader, we include them for completeness.
Lemma 2.
Let $H$ satisfy (1.3), consider the domain $\Omega=(0,1)^{d}$ and, for any $N\in\mathbb{N}$ and $\delta=1/N$, let $\overline{\Omega}_{\delta}$ be the grid defined in (7.1). For any function $u\in C(\overline{\Omega})$ with Lipschitz constant $L>0$, let $U$ be the grid function associated to $u$ on the grid $\overline{\Omega}_{\delta}$, and set $V=(V^{(1)},\ldots,V^{(d)})$, where for each $i\in\{1,\ldots,d\}$, $V^{(i)}$ is the grid function on $\overline{\Omega}_{\delta}$ defined by (7.8).
Then, the linear operator defined, for any grid function $W$ on $\overline{\Omega}_{\delta}$, by
$A_{N}(V)W=\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right),$
satisfies
$\overline{\sigma}\left(A_{N}(V)\right):=\max_{W}\dfrac{\|A_{N}(V)W\|_{2}}{\|W\|_{2}}\leq 2dC_{H}(L),$
where $C_{H}(L)=\displaystyle\max_{\begin{subarray}{c}\|p\|\leq L\\ x\in\overline{\Omega}\end{subarray}}\|\nabla_{p}H(x,p)\|$.
Proof.
Using the definition of $V^{(i)}$ in (7.8) and the fact that $u$ has Lipschitz constant $L$, it holds that
$V_{\beta}^{(i)}\leq C_{H}(L),\quad\forall\beta\in\mathfrak{B}\ \text{and}\ i\in\{1,\ldots,d\}.$
Then, in view of the definition of $A_{N}(V)$, it follows that
$\overline{\sigma}(A_{N}(V))\leq C_{H}(L)\overline{\sigma}(\overline{A}),$ (7.11)
where $\overline{A}$ is the linear operator, on the space of grid functions on $\Omega_{\delta}$, given by
$\overline{A}=\sum_{i=1}^{d}A^{(i)},$
with
$A^{(1)}$ $=$ $B\otimes I\otimes\cdots\otimes I$
$A^{(2)}$ $=$ $I\otimes B\otimes\cdots\otimes I$
$\vdots$
$A^{(d)}$ $=$ $I\otimes I\otimes\cdots\otimes B,$
where $I$ is the $(N-1)\times(N-1)$ identity matrix, and $B\in\mathbb{R}^{(N-1)\times(N-1)}$ is given by
$B=\begin{bmatrix}0&1&\\ 1&0&1\\ &1&0&1\\ &&\ddots&\ddots&\ddots\\ &&&1&0&1\\ &&&&1&0\end{bmatrix}.$
Finally, by the Gershgorin Theorem, the largest eigenvalue of $B$ is smaller or equal to $2$, so we conclude that
$\overline{\sigma}(\overline{A})\leq\sum_{i=1}^{d}\overline{\sigma}\left(A^{(i)}\right)\leq 2d,$
and the conclusion of the lemma follows from (7.11). ∎
Lemma 3.
Let $N\in\mathbb{N}$, $\delta=1/N$, and consider the grid $\Omega_{\delta}$ as in (7.1) and the index domain $\mathfrak{B}$ as in (7.3). The linear operator $\Delta_{N}$, defined for any grid function $W$ on $\Omega_{\delta}$ as
$\Delta_{N}W=\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right),$
with $W_{\beta\pm e_{i}}=0$ whenever $\beta\pm e_{i}\not\in\mathfrak{B}$, satisfies
$\underline{\sigma}(\Delta_{N}):=\min_{W}\dfrac{\|\Delta_{N}W\|_{2}}{\|W\|_{2}}=4d\sin^{2}\left(\dfrac{\pi}{2}\delta\right).$
Proof.
The operator $\Delta_{N}$ is the Dirichlet Laplacian associated with the $d$-dimensional grid $\Omega_{\delta}$ and is a symmetric negative definite operator. Hence, $\underline{\sigma}(\Delta_{N})$ coincides with the smallest eigenvalue of $-\Delta_{N}$. We can write it as
$-\Delta_{N}=-\sum_{i=1}^{d}\Delta_{N}^{(i)},$
with
$\Delta_{N}^{(1)}$ $=$ $L_{N}\otimes I\otimes\cdots\otimes I$
$\Delta_{N}^{(2)}$ $=$ $I\otimes L_{N}\otimes\cdots\otimes I$
$\vdots$
$\Delta_{N}^{(d)}$ $=$ $I\otimes I\otimes\cdots\otimes L_{N},$
and where
$L_{N}=\begin{bmatrix}-2&1&\\ 1&-2&1\\ &1&-2&1\\ &&\ddots&\ddots&\ddots\\ &&&1&-2&1\\ &&&&1&-2\end{bmatrix}\in\mathbb{R}^{(N-1)\times(N-1)}.$
Any eigenvector of $-\Delta_{N}$ can be written as $\Phi=\bigotimes_{i=1}^{d}\phi^{(i)},$ where $\phi^{(i)}\in\mathbb{R}^{N-1}$ is an eigenvector of $-L_{N}$ with associated eigenvalue $\Lambda=\sum_{i=1}^{d}\lambda^{(i)}$, where each $\lambda^{(i)}$ is the eigenvalue associated to $\phi^{(i)}$. Hence, the smallest eigenvalue of $-\Delta_{N}$ is $d\lambda_{1}$, where $\lambda_{1}$ is the smallest eigenvalue of $-L_{N}$. Since the latter is well-known to be $\lambda_{1}=4\sin^{2}\left(\frac{\pi}{2N}\right)$, we conclude that
$\underline{\sigma}(\Delta_{N})=d\lambda_{1}=4d\sin^{2}\left(\dfrac{\pi}{2N}\right)=4d\sin^{2}\left(\dfrac{\pi}{2}\delta\right).$
Next, we give the proof of Corollary 1. The proof is almost identical to that of Theorem 1, with the only difference being that one has to consider the functional’s partial derivatives with respect to the boundary points.
Proof of Corollary 1.
First, for any function $u\in C(\overline{\Omega})$, let us write the functional $\widehat{\mathcal{J}}(u)$ as
$\widehat{\mathcal{J}}(u)=F(U)+\gamma G(U),$
where $F(U)$ is defined as in (7.4), and $G(U)$ is given by
$G(U):=\delta^{d-1}\sum_{\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}}\left(U_{\beta}-g(x_{\beta})\right)^{2}.$
We use the multi-index notation introduced in (7.3).
Using precisely the same argument as in step 1 in the proof of Theorem 1, we have that if $u$ is a critical point of $\widehat{\mathcal{J}}(u)$, then it must hold that $\nabla F(U)+\gamma\nabla G(U)=0$.
Since $G(U)$ does not depend on $U_{\beta}$ for any $\beta\in\mathfrak{B}$, we have that $u$ being a critical point of $\widehat{\mathcal{J}}(u)$ implies that
$\partial_{U_{\beta}}F(U)=-\delta^{d-1}\left(\sum_{i=1}^{d}\left(V_{\beta+e_{i}}^{(i)}W_{\beta+e_{i}}-V_{\beta-e_{i}}^{(i)}W_{\beta-e_{i}}\right)+\alpha\sum_{i=1}^{d}\left(W_{\beta+e_{i}}+W_{\beta-e_{i}}-2W_{\beta}\right)\right)=0,\qquad\forall\beta\in\mathfrak{B},$
where $W_{\beta}$ and $V_{\beta}^{(i)}$ are given as in (7.7) and (7.8). Now, using the step 2 in the proof of Theorem 1, we obtain that, if (4.8) holds, then $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$.
Since $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$, we have that
$\partial_{U_{\beta}}F(U)+\gamma\partial_{U_{\beta}}G(U)=\gamma\partial_{U_{\beta}}G(U)=2\gamma\delta^{d-1}\left(U_{\beta}-g(x_{\beta})\right),\qquad\forall\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}.$
Hence, $\partial_{U_{\beta}}F(U)+\gamma\partial_{U_{\beta}}G(U)=0$ implies that $U_{\beta}=g(x_{\beta})$ for all $\beta\in\overline{\mathfrak{B}}\setminus\mathfrak{B}$. This, combined with $W_{\beta}=0$ for all $\beta\in\mathfrak{B}$ concludes the proof. ∎
7.3 Proof of Theorem 2
The main idea of the proof is to re-write the functional $\mathcal{R}(\cdot)$ in such a way that we can apply the conclusion of Theorem 1 for the discrete functional $\widehat{\mathcal{R}}(\cdot)$.
Proof.
Let $N\in\mathbb{N}$ and fix $\delta=1/(N-1)$. The interval $[0,1)$ can be partitioned in $N-1$ disjoint intervals as
$[0,1)=\left[0,\dfrac{1}{N-1}\right)\sqcup\left[\dfrac{1}{N-1},\dfrac{2}{N-1}\right)\sqcup\ldots\sqcup\left[\dfrac{N-2}{N-1},1\right)=\bigsqcup_{\beta\in\{1,\ldots,N-1\}}\left(\delta\beta+[-\delta,0)\right).$
Hence, for any dimension $d$, we can write $[0,1)^{d}$ similarly as a disjoint union of $(N-1)^{d}$ disjoint $d$-dimensional cubes as
$[0,1)^{d}=\bigsqcup_{\beta\in\mathfrak{B}}\left(\delta\beta+[-\delta,0)^{d}\right),\quad\text{where $\mathfrak{B}:=\{1,\ldots,N-1\}^{d}$.}$
Now, recalling that $\Omega=(0,1)^{d}$, we can use this partition to write $\mathcal{R}(u)$ defined in (4.11) as
[…]
$=$ $\int_{(-\delta,0)^{d}}\sum_{\beta\in\mathfrak{B}}\left[\widehat{H}_{\alpha}(z-\delta\beta,D_{\delta}^{+}u(z-\delta\beta),D_{\delta}^{-}u(z-\delta\beta))\right]^{2}dz$
[…]
where
^$\widehat{\mathcal{R}}(u;z):=\sum_{x\in\Omega_{\delta}(z)}\left[\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))\right]^{2},$
with $\Omega_{\delta}(z):=\{x_{\beta}=z+\delta\beta\,:\ \beta\in\mathfrak{B}\}$ for all $z\in(0,1)^{d}$.
Note that $\Omega_{\delta}(z)$ is simply a shift of the grid $\Omega_{\delta}$ defined in (7.2). Hence, by virtue of Theorem 1, if $u\in C(\overline{\Omega})$ is a critical point of ^$\widehat{\mathcal{R}}(u;z)$, then $\widehat{H}_{\alpha}(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x))=0$ for all $x\in\Omega_{\delta}(z)$.
For any functions $u,v\in C(\overline{\Omega})$, and any $z\in(0,1)^{d}$, let $U(z)$ and $V(z)$ be the grid functions of $u$ and $v$ associated to the grid $\Omega_{\delta}(z)$, i.e.
$U(z):=\{u(x)\,:x\in\Omega_{\delta}(z)\},\quad\text{and}\quad V(z):=\{v(x)\,:x\in\Omega_{\delta}(z)\}.$
We also define, for any $z\in(0,\delta)^{d}$, the functional on the space of grid functions on $\Omega_{\delta}(z)$ given by
$F_{z}(U(z)):=\sum_{x\in\Omega_{\delta}(z)}\left[\widehat{H}_{\alpha}\left(x,D_{\delta}^{+}u(x),D_{\delta}^{-}u(x)\right)\right]^{2}=\widehat{\mathcal{R}}(u;z).$
If $u\in C(\overline{\Omega})$ is a critical point of $\mathcal{R}(\cdot)$, then for any function $v\in C(\overline{\Omega})$ it must hold that
$\lim_{\varepsilon\to 0}\dfrac{\mathcal{R}(u+\varepsilon v)-\mathcal{R}(u)}{\varepsilon}=0,$
or equivalently, we can use (7.12) to write
[…]
Since this holds for any function $v\in C(\overline{\Omega})$, we deduce that $\nabla F_{z}\left(U(z)\right)$ is identically 00 for all $z\in(-\delta,0)^{d}$. Whence, $u$ is a critical point of ^$\widehat{\mathcal{R}}(u;z)$ for all $z$, and the conclusion follows. ∎
8 Summary and perspectives
In this paper, we propose to solve the boundary value problems for a class of Hamilton-Jacobi equations using Deep Learning. Artificial neural networks are trained by typical gradient descent-based algorithms, optimising for the least square problem defined by the residual of a consistent and monotone finite difference scheme for the equation. We analyze the critical points of the least square problem in the space of continuous functions and reveal conditions that guarantee the critical points are the unique viscosity solution of the problem. These conditions may involve the parameters of the chosen finite difference scheme or those coming from additional “supervised” data points. We demonstrated numerically that the proposed approach effectively solves Hamilton-Jacobi equations in higher dimensions under a favourable setup.
An important point is that artificial neural networks are used instead of grid functions. Thus, one may take advantage of neural networks’ approximation power for the smooth parts of a solution and the scaling property of Monte-Carlo sampling involved in the training procedure. Thus, under favourable conditions, it is reasonable to expect that the proposed approach would suffer less the so-called “curse of dimensionality” than the classical algorithms.
But during the training iterations, one must ensure that the (high dimensional) computational domain is sufficiently “covered” — enough to effectively follow the characteristics from the boundaries to everywhere in the domain and “monitor” how the monotone scheme resolves the collision of the characteristics. This is crucial for computing accurate approximation of the viscosity solutions.
It is widely acknowledged that the choice of collocation points can determine the success of a PINN approach. This is also true for the proposed approach, given a limited training budget. In higher dimensions, due to the “concentration of measure” errors, the errors “deeper” in the interior of $\Omega$ are significantly “discounted.” However, reducing the errors in the interior may be essential to constructing approximate solutions with the proper global structure (such as convexity). Therefore, uniform sampling (relative to the Lebesque measure) in training or assigning additional “supervised” data may be sub-optimal. We plan to develop adaptive algorithms for the proposed approach in the near future.
We point out that it is possible to use high-resolution schemes. If a gradient descent-based optimization scheme relying automatic differentiation is employed, the numerical scheme selected must be a differentiable function with respect to the variables defined by the scheme’s stencil. This formerly precludes the use of higher order essentially non-oscillatory (ENO) schemes [24]. However, the weighted essentially oscillatory (WENO) schemes [17] can be considered. A potential advantage may come from the wider stencils of such a scheme that have a large domain of influence in the optimization model.
Finally, we remark that the idea of using a convergent numerical method to form least square principles is not limited to only finite difference methods. It is possible to develop such an algorithm using the finite element methodology. In that case, we envision that proper mesh refinement strategies may be involved during training to enhance error control.
Acknoledgement
Esteve-Yagüe is partly supported by the EPSRC grant EP/V029428/1, and by the Ramón y Cajal 2022 grant RYC2022-035966-I. Tsai is partially supported by NSF grant DMS-2110895 and Army Research Office Grant W911NF2320240.